### v1

# AnomalyDINO: Boosting Patch-based Few-shot Anomaly Detection with DINOv2

Abstract
Recent advances in multimodal foundation models have set new standards in few-shot anomaly detection. This paper explores whether high-quality visual features alone are sufficient to rival existing state-of-the-art vision-language models. We affirm this by adapting DINOv2 for one-shot and few-shot anomaly detection, with a focus on industrial applications. We show that this approach does not only rival existing techniques but can even outmatch them in many settings. Our proposed vision-only approach, AnomalyDINO, is based on patch similarities and enables both image-level anomaly prediction and pixel-level anomaly segmentation. The approach is methodologically simple and training-free and, thus, does not require any additional data for fine-tuning or meta-learning. Despite its simplicity, AnomalyDINO achieves state-of-the-art results in one- and few-shot anomaly detection (e.g., pushing the one-shot performance on MVTec-AD from an AUROC of $93.1\%$ to $96.6\%$). The reduced overhead, coupled with its outstanding few-shot performance, makes AnomalyDINO a strong candidate for fast deployment, for example, in industrial contexts.
1 Introduction
Anomaly detection (AD) in machine learning attempts to identify instances that deviate substantially from the nominal data distribution $p_{\mathrm{norm}}(x)$. Anomalies, therefore, raise suspicion of being ‘generated by a different mechanism’ (Hawkins, 1980)—often indicating critical, rare, or unforeseen events. The ability to reliably distinguish anomalies from normal samples is highly valuable across various domains, including security (Siddiqui et al., 2019), healthcare (Fernando et al., 2021; Tian et al., 2023), and industrial inspection. In this work, we focus on the latter, where fully automated systems necessitate the ability to detect defective or missing parts to prevent malfunctions in downstream products, raise alerts for potential hazards, or analyze these to optimize production lines. See the right-hand side of Figure 1 for anomalous samples in this context (with predicted anomaly maps from AnomalyDINO).
Figure 1: Anomaly detection with AnomalyDINO based on a single immaculate reference sample (here category ‘Screw’ from MVTec-AD). Whenever suitable we select the relevant patches from the (potentially augmented) reference sample via masking and collect the nominal patch representations in $\mathcal{M}$. For both, masking and feature extraction, we utilize DINOv2. At test time, the distances of patch representations to those in $\mathcal{M}$ give rise to an anomaly map and the corresponding anomaly score after aggregation with $q$. Further examples for other categories are depicted on the right (and in Figures 3 and 4 in Appendix A).
AD for industrial images has gained tremendous interest over the last couple of years. The close-to-optimal results on benchmark data make it seem that the problem of anomaly detection is essentially solved. For instance, Mousakhan et al. (2023) report $99.8\%$ and $98.9\%$ AUROC on the popular benchmarks MVTec-AD (Bergmann et al., 2021) and VisA (Zou et al., 2022), respectively. The most popular AD techniques use the training data to train an anomaly classifier (e.g., Ruff et al., 2020), or a generative model coupled with reconstruction-based (e.g., Wolleb et al., 2022; Livernoche et al., 2023; Mousakhan et al., 2023) or likelihood-based (e.g., Rudolph et al., 2022; Daehyun Kim, 2023) anomaly scoring. However, these approaches operate within the full-shot setting, meaning they rely on access to a sufficiently large amount of training data. Given the challenges associated with dataset acquisition, the attractivity of a fast and easy-to-deploy methodology, and the requirement to rapidly adapt to covariate shifts in the nominal data distribution (Li et al., 2024a), there is an increasing interest in few-shot and zero-shot anomaly detection.
Few-shot techniques, however, heavily rely on meaningful features, or as Reiss et al. (2022) frame it: ‘anomaly detection requires better representations’. Such better representations are now available with the increasing availability and capabilities of foundation models, i.e., large-scale models trained on massive datasets in unsupervised/self-supervised fashion (e.g., Radford et al., 2021; Caron et al., 2021; Oquab et al., 2024). The performance of few-shot anomaly detection techniques has already been boosted by the use of foundation models, mostly by multimodal approaches (e.g., Jeong et al., 2023; Cao et al., 2023; Zhou et al., 2024; Kwak et al., 2024)
Here, we propose to focus on a vision-only approach, in contrast to such multimodal techniques. This perspective is motivated by the observation that few-shot anomaly detection is feasible for human annotators based on visual features only, and does not require additional textual description of the given object or the expected types of anomalies (which are typically not known a priori).
Our approach, termed AnomalyDINO, is memory-bank-based and leverages DINOv2 (Oquab et al., 2024) as a backbone (but any architecture with strong patch-level feature extraction is suitable for the proposed approach). We carefully design a suitable preprocessing pipeline for the one-shot scenario, using the segmentation abilities of DINOv2 (which alleviates the additional overhead of another model). At test time, anomalous samples are detected based on the high distances between their patch representations and the closest counterparts in the nominal memory bank.
Due to its simplicity, AnomalyDINO can be deployed in industrial contexts very easily—in strong contrast to more complex approaches such as those proposed by Chen et al. (2023) or Kwak et al. (2024). Yet, the proposed method achieves new state-of-the-art performance on anomaly detection in the few-shot regime on MVTec-AD (Bergmann et al., 2021) and outperforms all but one competing method on VisA (Zou et al., 2022).
The structure of the paper is as follows: Section 2 reviews relevant prior studies and clarifies the distinctions between the settings addressed by zero- and few-shot, and batched zero-shot techniques. Section 3 introduces our proposed method, AnomalyDINO. An extension of this method to the batched zero-shot scenario is detailed in Appendix D. Section 4 presents the experimental outcomes. Additional results and an ablation study are provided in Appendices A and C, respectively. We address identified failure cases of AnomalyDINO in Appendix B.
Contributions
  • We propose AnomalyDINO, a simple and training-free yet highly effective patch-based technique for visual anomaly detection. Our method builds upon the high-quality feature representation extracted by DINOv2.
  • An extensive analysis demonstrates the efficiency and effectiveness of the proposed approach, outperforming other multimodal few-shot techniques, thereby reducing the gap between few- and full-shot anomaly detection. Specifically, AnomalyDINO achieves state-of-the-art results for few-shot anomaly detection on MVTec-AD, e.g., pushing the one-shot detection from an AUROC of 93.1% to 96.6%. Moreover, our results on VisA are not only competitive with other few-shot methods but also establish new state-of-the-art performance among all training-free few-shot anomaly detectors.
2 Related Work
Foundation Models for Vision
Multimodal foundation models have emerged as powerful tools for a wide range of tasks (see e.g., Bommasani et al., 2022; Li et al., 2023; Chen et al., 2020; He et al., 2022; Radford et al., 2021; Liu et al., 2023; OpenAI, 2024). Most relevant to visual AD are multimodal approaches based on CLIP (Radford et al., 2021) or recent LLMs (e.g., OpenAI, 2024), but also vision-only approaches like DINO (Caron et al., 2021; Oquab et al., 2024). CLIP (Radford et al., 2021) learns visual concepts from natural language descriptions by training on a dataset of images paired with textual annotations. The model uses a contrastive learning objective that aligns the embeddings from image and text encoders, optimizing the similarity between corresponding image-text pairs. This common feature space for vision and language can be utilized for several downstream tasks, such as zero-shot image classification by assessing similarities to a set of class-specific prompts. DINO (Caron et al., 2021; Oquab et al., 2024) leverages a self-supervised student-teacher framework based on vision transformers (Dosovitskiy et al., 2021). It employs a multi-view strategy to train models to predict softened teacher outputs, thereby learning robust and high-quality features for downstream tasks. DINOv2 (Oquab et al., 2024) combines ideas from DINO with patch-level reconstruction techniques (Zhou et al., 2022) and scales to larger architectures and datasets. The features extracted by DINO are well-suited for anomaly detection as they incorporate both local and global information, are robust to multiple views and crops, and benefit from large-scale pre-training. GroundingDINO (Liu et al., 2023), builds upon the DINO framework and focuses on improving the alignment of textual and visual information, enhancing the model’s performance in tasks requiring detailed object localization and multimodal understanding.
Anomaly Detection
Given a predefined notion of normality, the anomaly detection task is to detect test samples that deviate from this concept (Yang et al., 2021; Ruff et al., 2021). In this work, we focus on low-level sensory anomalies of industrial image data, i.e., we do not target the detection of semantic anomalies but of low-level features such as scratches of images of industrial products (see e.g., Figure 1). Several works tackled this task by either training an anomaly classifier (e.g., Ruff et al., 2020) or a generative model, which allows for reconstruction-based or likelihood-based AD (e.g., Wolleb et al., 2022; Livernoche et al., 2023; Zhang et al., 2023; Daehyun Kim, 2023; Mousakhan et al., 2023).
However, this work focuses on few-shot AD, which tackles anomaly detection under strong restrictions on the number of available training samples. Various recent methods recognized the effectiveness of pre-trained vision models for few-shot AD. The underlying idea in these approaches is to leverage the feature representations originating from pre-trained ResNets (He et al., 2015), Wide ResNets (Zagoruyko and Komodakis, 2016), and Vision Transformers (ViT, Dosovitskiy et al., 2020). These models can extract representations on a patch and pixel level, which can be compared against a memory bank, which aggregates patch and pixel features provided by the training set. The comparison with the memory bank is typically conducted using a nearest neighbor approach (Bergman et al., 2020; Cohen and Hoshen, 2020; Defard et al., 2020; Roth et al., 2022; Xie et al., 2023; Li et al., 2024b). Loosely speaking, patches and pixels corresponding to anomalies are expected to have a high distance to their nearest neighbor in the memory bank. Another line of work builds upon the success of pre-trained language-vision models in zero-shot classification. The underlying idea consists of two steps. First, these approaches define sets of prompts describing nominal samples and anomalies. Second, the corresponding textual embeddings are compared against the image embeddings (Jeong et al., 2023; Chen et al., 2023; Kwak et al., 2024; Zhou et al., 2024). Images whose visual embedding is close to the textual embedding of a prompt associated with an anomaly are classified as anomalous. However, these methods either require significant prompt engineering (e.g., Chen et al., 2023, use a total of $35\times 7$ different prompts for describing normal samples) or fine-tuning of the prompt(-embeddings). Lastly, another type of few-shot anomaly detection builds upon the success of multimodal chatbots. These methods require more elaborate prompting and techniques for interpreting textual outputs (Xu et al., 2024). Since these methods do not require a memory bank, they are capable of performing zero-shot anomaly detection.
Categorization of Few-/Zero-Shot Anomaly Detectors
Previous works consider different AD setups, which complicates their evaluation and comparison. To remedy this, we provide a taxonomy of recent few- and zero-shot AD based on the particular ‘shot’-setting, the training requirements, and the modes covered by the underlying models. We categorize three ‘shot’-settings: zero-shot, few-shot, and batched zero-shot. Zero- and few-shot settings are characterized by the number of nominal training samples a method can process, before making predictions on the test samples. In batched zero-shot, inference is not performed sample-wise but based on a whole batch of test samples, usually the full test set. For instance, the method proposed by Li et al. (2024b) benefits from the fact that a significant majority of pixels correspond to normal pixels, which motivates the strategy of matching patches across a batch of images. Another work that considers this setting is that of Li et al. (2024a), who deploy a parameter-free anomaly detector based on the effect of batch normalization. We split the training requirements into the categories ‘Training-Free’, ‘Fine-Tuning’, and ‘Meta-Training’. ‘Training-Free’ approaches do not require any training, while ‘Fine-Tuning’ methods use the few accessible samples to modify the underlying model. In contrast, ‘Meta-Training’ is associated with training the model on a dataset related to the test data. For example, Li et al. (2024a) train their model on MVTec-AD containing all classes except the class they test against. Zhou et al. (2024) and Chen et al. (2023) train their model on VisA when evaluating the test performance on MVTec-AD and vice versa. Finally, we differentiate the leveraged models, which are either vision models (such as pre-trained ViT) or language-vision models (such as CLIP). We provide a detailed summary in Table 1.
Table 1: Taxonomy of recent few- and zero-shot anomaly detection methods. The $\dagger$ indicates approaches that were introduced as full-shot detectors but then considered as few-shot detectors in later works, see e.g., Roth et al. (2022).
Method Setting Training Type Modes
DN2 (Bergman et al., 2020) Few-Shot Training-Free Vision
SPADE (Cohen and Hoshen, 2020) Few-Shot Training-Free Vision
PaDiM (Defard et al., 2020) Few-Shot Fine-Tuning Vision
PatchCore (Roth et al., 2022) Few-Shot Training-Free Vision
PatchCore–opt (Santos et al., 2023) Few-Shot Training-Free Vision
MuSc (Li et al., 2024b) Batched Zero-Shot Training-Free Vision
GraphCore (Xie et al., 2023) Few-Shot Training (GNN) Vision
WinCLIP (Jeong et al., 2023) Zero-/Few-Shot Training-Free Vision + Language
APRIL-GAN (Chen et al., 2023) Zero-/Few-Shot Meta-Training Vision + Language
ADP (Kwak et al., 2024) Few-Shot Fine-Tuning Vision + Language
AnomalyCLIP (Zhou et al., 2024) Zero-Shot Meta-Training Vision + Language
GPT4-V (Xu et al., 2024) Zero-Shot Training-Free Vision + Language
ACR (Li et al., 2024a) Batched Zero-Shot Meta-Training /
AnomalyDINO (Ours) Few-Shot Training-Free Vision
3 Patch-Matching for Visual Anomaly Detection
This section introduces AnomalyDINO, which leverages DINOv2 to extract meaningful patch-level features. The strong feature extraction capabilities of the utilized backbone allow us to fall back on a simple patch-based matching strategy to perform few-shot and training-free AD for image data. We follow a memory-bank approach, i.e., we first gather relevant patch representations of nominal samples in a memory bank $\mathcal{M}$. Then, for each test patch, we compute its distance to the nearest nominal patch in $\mathcal{M}$. A suitable aggregation of the patch-based distances gives the anomaly score on image-level.
3.1 Anomaly Detection with Nearest Patch Distances
Let us consider the image space $\mathcal{X}$. We assume to have access to a suitable feature extractor $f:\mathcal{X}\rightarrow\mathcal{F}^{n}$ that maps each image $\mathbf{x}$ to a tuple of patch-features $f(\mathbf{x})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n})$, where $\mathcal{F}$ denotes the feature space. Note that $n$ depends on the image resolution and the patch size. Given $k\geq 1$ nominal reference sample(s) $X_{\text{ref}}:=\bigl{\{}\mathbf{x}^{(i)}\mid\,i\in\{1,\dots,k\}\bigr{\}}$ ($k$-shot setting), we collect the nominal patch features and store them in a memory bank
$\mathcal{M}:=\bigcup_{\mathbf{x}^{(i)}\in X_{\operatorname{ref}}}\bigl{\{}\mathbf{p}_{j}^{(i)}\;|\;f(\mathbf{x}^{(i)})=(\mathbf{p}_{1}^{(i)},\dots,\mathbf{p}_{n}^{(i)}),\,j\in\{1,\dots,n\}\bigr{\}}\enspace.$ (1)
To score a test sample $\mathbf{x}_{\mathrm{test}}$, we collect the extracted patch representations $f(\mathbf{x}_{\mathrm{test}})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n})$ and then check how well they comply with our memory bank $\mathcal{M}$. To do so, we leverage a nearest-neighbor approach to find the distance to the closest reference patch for a given test patch $\mathbf{p}\in\mathcal{F}$
$d_{\operatorname{NN}}(\mathbf{p};\mathcal{M}):=\min_{\mathbf{p}_{\operatorname{ref}}\in\mathcal{M}}d(\mathbf{p},\,\mathbf{p}_{\operatorname{ref}})$ (2)
for some distance metric $d$. In our experiments, we set $d$ as the cosine distance, that is,
$d(\mathbf{x},\mathbf{y}):=1-\frac{\langle\mathbf{x},\;\mathbf{y}\rangle}{\|\mathbf{x}\|\,\|\mathbf{y}\|}\enspace.$ (3)
The image-level score $s(\mathbf{x}_{\mathrm{test}})$ is given by aggregating the patch distances via a suitable statistic $q$
$s(\mathbf{x}_{\mathrm{test}}):=q\big{(}\{d_{\operatorname{NN}}(\mathbf{p}_{1};\mathcal{M}),\dots,d_{\operatorname{NN}}(\mathbf{p}_{n};\mathcal{M})\}\big{)}\enspace.$ (4)
Throughout this paper, we define $q$ as the average distance of the 1% most anomalous patches, i.e, $q(\mathcal{D}):=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ with $H_{0.01}(\mathcal{D})$ containing the 1% highest values in $\mathcal{D}$. The statistic $q$ can be understood as an empirical estimate of the tail value at risk for the 99% quantile (e.g., McNeil, 1999) and turns out suitable for a wide range of settings because it balances two desirable properties: First, we want $s(\mathbf{x}_{\mathrm{test}})$ to depend on the highest patch distances, as they may provide the strongest anomaly signal. Similarly, we want a certain degree of robustness against a singular high patch distance (in particular in few-shot scenarios where $\mathcal{M}$ is sparsely populated). However, one could also replace $q$ with a different statistic to cater to special cases. If anomalies are expected to cover larger parts of the image, for example, a high percentile of the patch distances might be a suitable choice. If, on the contrary, anomalies may occur only locally such that only very few patches/or a single patch might be affected, the score $q$ should be sensitive to the highest patch distances. Frequently, the maximum pixel-wise anomaly score is considered. Such an anomaly score on pixel-level is usually obtained by upsampling to full image resolution and applying some smoothing operation.
Following Roth et al. (2022), we utilize bilinear upsampling and Gaussian smoothing (with $\sigma=4.0$) to turn the patch distances into pixel-level anomaly scores for localization of potential defects. Examples of the resulting anomaly maps are visualized in Figure 1 and Appendix A (Figures 3 and 4).
We extend AnomalyDINO also to the batched zero-shot scenario, see Appendix D (Figure 13).
3.2 Enriching the Memory Bank & Filtering Relevant Patches
Access to only a single/few normal sample(s) poses (at least) two challenges. First, all relevant information about the concept of normality must be extracted from the single/few available nominal sample(s). However, it is expected that the variability at test time is (considerably) higher than in $X_{\mathrm{ref}}$. Secondly, one must determine a suitable preprocessing for reference and test samples, solely based on $X_{\mathrm{ref}}$. To address low variations of nominal patch features, we apply simple augmentations in the form of rotations, following the insights of Xie et al. (2023). In addition, to ensure that $\mathcal{M}$ is of high quality, we consider masking the object of interest as background noise could poison $\mathcal{M}$. Note, that a suitable preprocessing, i.e., whether to apply masking or rotations, should depend on the object of interest (see, e.g., Figures 7 and 10, and Section C.1 for further discussion).
Figure 2: Masking test on MVTec-AD. The dashed red regions mark areas that belong to the object of interest but are falsely predicted to be background. See App. C.1 for the outcomes per category.
Masking
To minimize the overhead of the proposed method, we utilize DINOv2 also for masking, i.e., delineating the primary object(s) in an image from its background. This is achieved by thresholding the first PCA component of the patch features (Oquab et al., 2024). We observe empirically that this procedure sometimes produces erroneous masks. Frequently, such failure cases occur for close-up shots, where the objects of interest account for $\gtrsim 50\%$ of patches, see Figure 2. To resolve that issue we test whether the PCA-based mask correctly captures the object in the first reference sample and apply masking accordingly. In general, we do not mask textures (e.g., wood or tile in MVTec-AD).
Rotation
Rotating the reference sample may improve the detection performance by enriching the memory bank, but can also be detrimental in cases where rotations of (parts of the) object of interest can be considered anomalies themselves. We consider two different settings.
In the ‘agnostic’ case, which we focus on in the main paper, we always augment the reference sample with rotations. We also consider the case where we know about the samples’ potential rotations (‘informed’). The ‘informed’ case is sensible as the data collection process can usually be controlled in an industrial/medical setting (or test images can be aligned), and may lead to significantly lower inference time. See Section C.1 for the effect of masking and rotation, and a comparison of ‘informed’ vs. ‘agnostic’.
4 Experiments
AnomalyDINO – Defaults
Using DINOv2111Code and model weights of DINOv2 (Oquab et al., 2024) are available under Apache 2.0 license. as backbone allows us to choose from the distillation sizes, which range from small (ViT-S, $21\times 10^{6}$ parameters) to giant (ViT-G, $1.1\times 10^{9}$ parameters). To prioritize low latency, we take the smallest model (S) as our default. As discussed above, the ‘agnostic’ preprocessing is our default.
Datasets
For the experiments, we consider MVTec-AD [^2] The MVTec-AD dataset (Bergmann et al., 2021) is available under the CC BY-NC-SA 4.0 license. and VisA, [^3] The VisA dataset (Zou et al., 2022) is released under the CC BY 4.0 license. two datasets with high-resolution images for industrial anomaly detection. MVTec-AD consists of fifteen categories, both objects and textures, with a small amount of clean training samples (between 60 and 391), and up to eight anomaly types per category. VisA has twelve categories with up to 1006 normal samples per category and diverse types of anomalies condensed into one anomaly class ‘bad’.
Evaluation Metrics
We assess the image-level detection and the pixel-level segmentation performance with three metrics each (similar to, e.g., Jeong et al., 2023; Chen et al., 2023; Li et al., 2024b). For evaluating the detection performance, we measure the area under the receiver-operator curve (AUROC), the F1-score at the optimal threshold (F1-max), and the average precision (AP) using the respective image-level anomaly scores. We quantify the anomaly segmentation performance using the AUROC, F1-max, and the per-region overlap (PRO, Bergmann et al., 2021) of the segmentation using the pixel-wise anomaly scores. Note in this context that due to a high imbalance of nominal and anomalous pixels—for MVTec-AD we have 97.26% nominal pixel, for VisA even 99.45% (Li et al., 2024b)—it is not recommended to assess performance solely on segmentation AUROC (Rafiei et al., 2023). We repeat each experiment three times and report mean and standard deviation. [^4] The randomness stems from the choice of reference samples $X_{\mathrm{ref}}$. For straightforward reproducibility, we simply set $X_{\mathrm{ref}}$ to the first $k$, second $k$ and third $k$ train samples for the three different runs, respectively.
[…]
Setting Method Detection Segmentation
AUROC F1-max AP AUROC F1-max PRO
1-shot SPADE 82.9$\pm$2.6 91.1$\pm$1.0 91.7$\pm$1.2 92.0$\pm$0.3 44.5$\pm$1.0 85.7$\pm$0.7
PatchCore 86.3$\pm$3.3 92.0$\pm$1.5 93.8$\pm$1.7 93.3$\pm$0.6 53.0$\pm$1.7 82.3$\pm$1.3
[…]
AnomalyDINO 96.6$\pm$0.5 96.0$\pm$0.2 98.2$\pm$0.3 96.3$\pm$0.1 58.2$\pm$0.8 91.9$\pm$0.1
2-shot SPADE 81.0$\pm$2.0 90.3$\pm$0.8 90.6$\pm$0.8 91.2$\pm$0.4 42.4$\pm$1.0 83.9$\pm$0.7
PatchCore 83.4$\pm$3.0 90.5$\pm$1.5 92.2$\pm$1.5 92.0$\pm$1.0 50.4$\pm$2.1 79.7$\pm$2.0
GraphCore 91.9$\pm$/ / / 96.9$\pm$/ / /
[…]
4-shot SPADE 84.8$\pm$2.5 91.5$\pm$0.9 92.5$\pm$1.2 92.7$\pm$0.3 46.2$\pm$1.3 87.0$\pm$0.5
PatchCore 88.8$\pm$2.6 92.6$\pm$1.6 94.5$\pm$1.5 94.3$\pm$0.5 55.0$\pm$1.9 84.3$\pm$1.6
GraphCore 92.9$\pm$/ / / 97.4$\pm$/ / /
[…]
8-shot GraphCore 95.9$\pm$/ / / 97.8$\pm$/ / /
WinCLIP+ 94.6$\pm$0.1 / / / / /
[…]
16-shot WinCLIP+ 94.8$\pm$0.1 / / / / /
[…]
AnomalyDINO 98.4$\pm$0.1 97.8$\pm$0.2 99.3$\pm$0.0 97.1$\pm$0.1 60.2$\pm$0.1 93.0$\pm$0.1
Baselines
We compare AnomalyDINO with a range of modern zero- and few-shot AD models, e.g., SPADE (Cohen and Hoshen, 2020), PaDiM (Defard et al., 2020), PatchCode (Roth et al., 2022), WinCLIP+ (Jeong et al., 2023), and APRIL-GAN (Chen et al., 2023). It is important to note that ACR (Li et al., 2024a) and MuSc (Li et al., 2024b) consider the batched zero-shot setting (see Table 1), thus, covering a different setting than AnomalyDINO. We adapt AnomalyDINO to this setting in Appendix D. Moreover, APRIL-GAN and AnomalyCLIP, require training on a related dataset, which is in contrast to our proposed training-free method. In Tables 2 and 3, results reported by Jeong et al. (2023) are indicated by , the result of the WinCLIP re-implementation by Kwak et al. (2024) (where J. Jeong is also coauthor) by . All other results are taken from the original publication. For ADP (Kwak et al., 2024) we report the (usually slightly better) variant ADP (with access to the class label). For GraphCore no standard deviations are reported (Xie et al., 2023).
4.1 Few-Shot Anomaly Detection and Anomaly Segmentation
Results
We summarize the results for few-shot anomaly detection on MVTec-AD and VisA in Table 2 and Table 3, respectively. Regarding MVTec-AD, our method achieves state-of-the-art $k$-shot detection performance across all $k\in\{1,2,4,8,16\}$ in every reported metric, outperforming even approaches that require training on an additional data set (e.g., ADP and APRIL-GAN).
[…]
Setting Method Detection Segmentation
AUROC F1-max AP AUROC F1-max PRO
1-shot SPADE 79.5$\pm$4.0 80.7$\pm$1.9 82.0$\pm$3.3 95.6$\pm$0.4 35.5$\pm$2.2 84.1$\pm$1.6
PaDiM 62.8$\pm$5.4 75.3$\pm$1.2 68.3$\pm$4.0 89.9$\pm$0.8 17.4$\pm$1.7 64.3$\pm$2.4
PatchCore 79.9$\pm$2.9 81.7$\pm$1.6 82.8$\pm$2.3 95.4$\pm$0.6 38.0$\pm$1.9 80.5$\pm$2.5
WinCLIP+ 83.8$\pm$4.0 83.1$\pm$1.7 85.1$\pm$4.0 96.4$\pm$0.4 41.3$\pm$2.3 85.1$\pm$2.1
APRIL-GAN 91.2$\pm$0.8 86.9$\pm$0.6 93.3$\pm$0.8 96.0$\pm$0.0 38.5$\pm$0.3 90.0$\pm$0.1
[…]
2-shot SPADE 80.7$\pm$5.0 81.7$\pm$2.5 82.3$\pm$4.3 96.2$\pm$0.4 40.5$\pm$3.7 85.7$\pm$1.1
PaDiM 67.4$\pm$5.1 75.7$\pm$1.8 71.6$\pm$3.8 92.0$\pm$0.7 21.1$\pm$2.4 70.1$\pm$2.6
PatchCore 81.6$\pm$4.0 82.5$\pm$1.8 84.8$\pm$3.2 96.1$\pm$0.5 41.0$\pm$3.9 82.6$\pm$2.3
[…]
ADP 86.9$\pm$0.9 / / / / /
APRIL-GAN 92.2$\pm$0.3 87.7$\pm$0.3 94.2$\pm$0.3 96.2$\pm$0.0 39.3$\pm$0.2 90.1$\pm$0.1
[…]
4-shot SPADE 81.7$\pm$3.4 82.1$\pm$2.1 83.4$\pm$2.7 96.6$\pm$0.3 43.6$\pm$3.6 87.3$\pm$0.8
PaDiM 72.8$\pm$2.9 78.0$\pm$1.2 75.6$\pm$2.2 93.2$\pm$0.5 24.6$\pm$1.8 72.6$\pm$1.9
[…]
ADP 88.4$\pm$0.4 / / / / /
APRIL-GAN 92.6$\pm$0.4 88.4$\pm$0.5 94.5$\pm$0.3 96.2$\pm$0.0 40.0$\pm$0.1 90.2$\pm$0.1
[…]
8-shot WinCLIP+ 85.0$\pm$0.0 / / / / /
ADP 89.2$\pm$0.1 / / / / /
[…]
16-shot WinCLIP+ 85.0$\pm$0.1 / / / / /
ADP 90.1$\pm$0.5 / / / / /
[…]
AnomalyDINO reduces the gap between few-shot and full-shot methods in anomaly detection significantly, and it demonstrates the strongest performance in anomaly localization. Furthermore, we see a clear improvement in the method’s results as the quantity of samples increases.
In terms of anomaly detection in the VisA benchmark (Table 3), APRIL-GAN demonstrates the strongest performance for $k\in\{1,2,4\}$. Nonetheless, AnomalyDINO consistently achieves second-best results, comparable results with APRIL-GAN in the 8-shot setting, and sets new state-of-the-art for $k=16$. This can be attributed to AnomalyDINO’s ability to benefit more from a richer memory bank $\mathcal{M}$ than APRIL-GAN. We hypothesize that meta-learning exerts a greater influence on APRIL-GAN (i.e., training on MVTec-AD, when testing on VisA, and vice-versa) compared to learning from the nominal features of the given reference samples.Note also that, AnomalyDINO outperforms all other training-free approaches.
Table 4: Detection results for other settings. All results are AUROC values in (%).
Setting Method MVTec-AD VisA
0-shot WinCLIP 91.8 78.1
AnomalyCLIP 91.5 82.1
APRIL-GAN 86.1 78.0
Batched 0-shot ACR 85.8 /
MuSc 97.8 92.8
AnomalyDINO 93.3 89.6
Comparing the segmentation performance on VisA, Table 3 reveals a mixed picture: WinCLIP+ shows strong localization performance in terms of pixel-level AUROC, and APRIL-GAN achieves the highest scores in PRO. AnomalyDINO consistently achieves results in per-region-overlap that are second-best compared to all other, and the best among the training-free methods. In addition, AnomalyDINO shows the best segmentation performance when measured in F1-max, a metric that accounts for the strong imbalance of nominal and anomalous pixels.
We have further adapted AnomalyDINO to the batched zero-shot setting (see Appendix D). This adaptation is straightforward and relatively simple, especially compared to MuSc. Notably, we did not employ additional techniques such as ‘Re-scoring with Constrained Image-level Neighborhood’ or ‘Local Neighborhood Aggregation with Multiple Degrees’ (Li et al., 2024b). The results obtained, as detailed in Table 4, are therefore quite satisfactory.
We conclude that AnomalyDINO—despite its simplicity—rivals other methods in all settings and even comes out ahead in most of the settings. Within the class of training-free models, it is the clear winner essentially across the board. These results do not only demonstrate the virtues of AnomalyDINO itself but, more generally, highlight the merits of strong visual features as compared to highly engineered architectures. We provide further qualitative results in Figures 3 and 4, and discuss the limitations (such as detecting semantic anomalies) and specific failure cases of our approach in Appendix B.
4.2 Ablation Study
We conduct additional experiments to assess the effect of specific design choices in our pipeline. The full results are provided in Appendix C. Here, we briefly describe the main insights.
Preprocessing
Comparing the two different preprocessing settings, ‘agnostic’ (our default) and ‘informed’ we observe that the agnostic preprocessing slightly outperforms the informed counterpart, see Figures 8 and 9. Depending on the product category suitable preprocessing steps lead to substantial improvements. In some cases augmenting the reference sample with rotations also negatively impacts the detection performance. For the full discussion, see Section C.1.
Inference time
Data augmentation increases the computational complexity of the proposed pipeline, as well as the memory bank size, and thus, the runtime of our nearest neighborhood approach. [^5] In practice, our implementation leverages GPU-accelerated neighborhood search (Johnson et al., 2019), leading to only slightly increased inference time. With increasing size of $\mathcal{M}$ reduction techniques such as coreset subsampling (Roth et al., 2022) are advisable. The effects of masking and augmentations on the inference time, presented in Figure 11, are not negligible but can be reduced by approximately 30% in the ‘informed’ scenario (on MVTec-AD).
Aggregation statistic
We compare the default scoring method, the empirical tail value at risk, to other suitable aggregations in Section C.2. The mean of the 1% highest distances from test patches to $\mathcal{M}$ improves over the standard choice (maximum of the upsampled and smoothed patch distances).
Architecture size
In addition, we evaluated our pipeline with DINOv2 at different scales (ViT-S, ViT-B, ViT-L). The full results are given in Section C.3, Figure 12. We observe no considerable differences, in particular, also larger backbones give state-of-the-art results on MVTec-AD (all $k$) and VisA ($k=16$). Interestingly, larger architectures do not necessarily translate to higher performance. As the dimensionality of the feature space $\mathcal{F}$ increases with larger architectures, the distances between nominal patches also increases. The smallest model demonstrates the best performance on MVTec-AD, which primarily consists of single objects and less complex scenes. In contrast, the larger architecture size L performs better on VisA, which often involves multiple and more complex objects. This suggests the importance of achieving the ‘just right’ level of abstraction for the specific task, particularly in the few-shot regime.
5 Conclusion
This paper proposes a vision-only approach for one- and few-shot anomaly detection. Our method is based on similarities between patch representations extracted by DINOv2 (Oquab et al., 2024). We carefully design means to populate the nominal memory bank with diverse and relevant features utilizing zero-shot segmentation and simple data augmentation. Industrial settings require fast deployment, easy debugging and error correction, and rapid adaptation for covariance shifts in the normal data distribution. Our pipeline caters to these requirements through its simplicity and computational efficiency. The proposed method, AnomalyDINO, achieves state-of-the-art results on MVTec-AD and rivals all competing methods on VisA while outperforming all other training-free approaches. Its simplicity and strong performance render AnomalyDINO an excellent candidate for practitioners for industrial anomaly detection and an effective baseline for assessing few-shot and even full-shot anomaly detection in ongoing research.
Follow-up research directions
The specific pipeline exemplified in the paper focuses on simplicity and high throughput. However, the individual parts of our pipeline can easily be exchanged for more sophisticated alternatives. For example, our simple masking approach could be replaced by more specialized and adaptive masking techniques (which may also be relevant to other methods building on features extracted by DINOv2), and the simple upsampling and smoothing approach could be substituted by more sophisticated methods (e.g., Fu et al., 2024) to obtain the anomaly map from patch features. It would be interesting to see if that leads to further improvements in anomaly detection and localization, thereby further reducing the gap between few- and full-shot anomaly detectors. We also plan to improve the batched zero-shot performance of AnomalyDINO.
Broader Impacts
Advancing few-shot visual anomaly detection methods can offer various benefits by enhancing manufacturing quality control through the rapid identification of defects with minimal nominal examples, which improves efficiency, reduces waste, and improves overall safety in the product lifecycle. Similar positive benefits can be expected outside the industrial domain, e.g., for healthcare diagnostics or environmental monitoring. It is, however, essential to recognize the shortcomings of automated anomaly detection systems. We believe that simpler methods can be adapted more quickly, monitored more effectively, and are therefore more reliable. In this context, awareness of the risks of overreliance must be heightened (see Appendix B for identified failure cases for the proposed method). In addition, strong visual anomaly detectors could also lead to potentially malicious or unintended uses. To address these concerns, including potential privacy infringements and possible socioeconomic impacts of automation, strategies such as establishing robust data governance, and implementing strict privacy protections are essential. Additionally, investing in workforce development can help manage the socioeconomic effects of automation and leverage the full potential of strong visual anomaly detection.
Appendix A Detailed experimental results
In this section, we present the results of the default setting (AnomalyDINO-S and ‘agnostic’ preprocessing) in detail. The full results per category for MVTec-AD and VisA are given in Tables 5 and 6. Additional visualizations of anomaly maps are presented in Figures 3 and 4 for MVTec-AD and VisA, respectively.
query GT 1-shot 8-shot query GT 1-shot 8-shot
Figure 3: Examples – MVTec-AD. Depicted are, from left to right, a test sample per category (query), the ground truth anomaly annotation (GT), and the predicted anomaly map from AnomalyDINO in the 1- and 8-shot settings. The color coding is normalized by the maximum score on the ‘good’ test samples (per category).
Table 5: Detailed results on MVTec-AD. Reported are results on anomaly detection (image-level AUROC) and segmentation (PRO) of AnomalyDINO with default preprocessing (mean and standard deviation over three independent runs for each metric, all results in %).
Shots 1-shot 2-shot 4-shot 8-shot 16-shot
AUROC PRO AUROC PRO AUROC PRO AUROC PRO AUROC PRO
Bottle 99.5$\pm$0.4 96.1$\pm$0.3 99.7$\pm$0.1 96.4$\pm$0.1 99.9$\pm$0.1 96.6$\pm$0.2 100.0$\pm$0.0 96.6$\pm$0.1 100.0$\pm$0.0 96.5$\pm$0.2
Cable 91.6$\pm$1.2 88.2$\pm$0.7 92.1$\pm$0.7 88.5$\pm$0.7 93.8$\pm$1.4 89.5$\pm$0.5 94.7$\pm$0.3 89.6$\pm$0.3 94.4$\pm$1.1 89.5$\pm$0.7
Capsule 89.8$\pm$5.9 96.4$\pm$0.4 88.9$\pm$7.1 96.7$\pm$0.5 95.4$\pm$0.9 97.3$\pm$0.1 94.7$\pm$1.3 97.4$\pm$0.1 95.5$\pm$0.3 97.5$\pm$0.1
Carpet 100.0$\pm$0.0 97.9$\pm$0.1 100.0$\pm$0.0 97.9$\pm$0.0 100.0$\pm$0.0 97.7$\pm$0.1 100.0$\pm$0.0 97.7$\pm$0.1 100.0$\pm$0.0 97.6$\pm$0.2
Grid 99.2$\pm$0.1 96.1$\pm$0.0 99.2$\pm$0.1 96.3$\pm$0.1 99.2$\pm$0.1 96.2$\pm$0.2 99.4$\pm$0.2 96.2$\pm$0.1 99.8$\pm$0.3 95.9$\pm$0.3
Hazelnut 98.1$\pm$1.1 97.5$\pm$0.4 99.7$\pm$0.4 97.7$\pm$0.4 99.8$\pm$0.1 97.8$\pm$0.1 100.0$\pm$0.0 97.8$\pm$0.2 100.0$\pm$0.0 97.7$\pm$0.1
Leather 100.0$\pm$0.0 97.9$\pm$0.2 100.0$\pm$0.0 97.5$\pm$0.3 100.0$\pm$0.0 97.1$\pm$0.3 100.0$\pm$0.0 97.0$\pm$0.1 100.0$\pm$0.0 96.7$\pm$0.1
Metal nut 99.9$\pm$0.2 93.8$\pm$0.1 100.0$\pm$0.1 94.4$\pm$0.3 100.0$\pm$0.0 95.1$\pm$0.1 100.0$\pm$0.0 95.2$\pm$0.3 100.0$\pm$0.0 95.5$\pm$0.1
Pill 93.1$\pm$1.5 96.9$\pm$0.2 94.9$\pm$2.2 97.1$\pm$0.2 96.3$\pm$0.1 97.3$\pm$0.0 97.1$\pm$0.4 97.4$\pm$0.1 97.6$\pm$0.3 97.5$\pm$0.1
Screw 90.1$\pm$1.0 92.5$\pm$0.2 92.3$\pm$0.2 93.5$\pm$0.0 91.3$\pm$2.4 93.6$\pm$0.7 92.5$\pm$0.8 94.6$\pm$0.4 93.5$\pm$1.4 95.3$\pm$0.2
Tile 100.0$\pm$0.0 88.6$\pm$0.2 100.0$\pm$0.0 88.2$\pm$0.2 100.0$\pm$0.0 87.9$\pm$0.2 100.0$\pm$0.0 87.5$\pm$0.2 100.0$\pm$0.0 86.9$\pm$0.6
Toothbrush 99.4$\pm$0.5 94.5$\pm$0.1 99.9$\pm$0.1 94.9$\pm$0.3 99.9$\pm$0.1 95.0$\pm$0.3 99.9$\pm$0.1 95.5$\pm$0.5 99.6$\pm$0.3 95.9$\pm$0.6
Transistor 90.9$\pm$2.1 64.4$\pm$0.5 89.2$\pm$5.9 64.8$\pm$3.1 91.9$\pm$3.5 66.5$\pm$1.4 96.4$\pm$1.8 70.3$\pm$1.6 98.3$\pm$0.5 72.7$\pm$1.0
Wood 98.8$\pm$0.4 93.9$\pm$0.3 98.7$\pm$0.6 93.8$\pm$0.2 98.6$\pm$0.6 93.8$\pm$0.1 98.6$\pm$0.2 93.8$\pm$0.2 98.2$\pm$0.4 93.5$\pm$0.5
Zipper 98.9$\pm$0.1 83.2$\pm$0.4 99.0$\pm$0.1 84.4$\pm$0.7 99.3$\pm$0.2 85.6$\pm$0.2 99.2$\pm$0.2 85.5$\pm$0.6 99.3$\pm$0.3 85.9$\pm$0.7
Mean 96.6$\pm$0.5 91.9$\pm$0.1 96.9$\pm$0.8 92.1$\pm$0.2 97.7$\pm$0.1 92.5$\pm$0.1 98.2$\pm$0.1 92.8$\pm$0.0 98.4$\pm$0.1 93.0$\pm$0.1
query GT 1-shot 8-shot query GT 1-shot 8-shot
Figure 4: Examples – VisA. Depicted are, from left to right, a test sample per category (query), the ground truth anomaly annotation (GT), and the predicted anomaly map from AnomalyDINO in the 1- and 8-shot settings. The color bar is normalized by the maximum score on the ‘good’ test samples (per category). For the category ‘Capsules’ (left column, second from top) we changed the color map for better visibility. Best viewed at a higher zoom level as some anomalies are quite small.
Table 6: Detailed results on VisA. Reported are results on anomaly detection (image-level AUROC) and segmentation (PRO) of AnomalyDINO with default preprocessing (mean and standard deviation over three independent runs for each metric, all results in %).
Shots 1-shot 2-shot 4-shot 8-shot 16-shot
AUROC PRO AUROC PRO AUROC PRO AUROC PRO AUROC PRO
Candle 83.8$\pm$0.7 77.6$\pm$0.6 86.2$\pm$3.8 78.4$\pm$0.1 88.0$\pm$3.0 79.3$\pm$0.3 89.9$\pm$1.0 79.7$\pm$0.2 92.0$\pm$0.8 80.7$\pm$0.3
Capsules 98.4$\pm$0.3 80.4$\pm$0.5 99.1$\pm$0.2 81.0$\pm$0.4 99.1$\pm$0.2 81.5$\pm$0.2 99.1$\pm$0.5 81.9$\pm$0.1 98.8$\pm$0.4 82.4$\pm$0.2
Cashew 85.3$\pm$5.6 95.0$\pm$1.5 89.3$\pm$5.6 95.7$\pm$1.1 93.8$\pm$0.5 96.2$\pm$0.8 95.1$\pm$0.7 95.9$\pm$0.3 95.5$\pm$0.1 96.2$\pm$0.3
Chewinggum 98.2$\pm$0.3 90.6$\pm$0.8 98.7$\pm$0.3 91.8$\pm$0.8 98.8$\pm$0.2 92.2$\pm$0.5 98.9$\pm$0.1 92.5$\pm$0.3 99.0$\pm$0.0 92.6$\pm$0.0
Fryum 95.1$\pm$0.3 92.6$\pm$0.3 96.6$\pm$0.5 93.2$\pm$0.1 97.2$\pm$0.5 93.6$\pm$0.1 97.9$\pm$0.6 93.8$\pm$0.1 97.8$\pm$0.6 94.0$\pm$0.1
Macaroni1 83.4$\pm$3.0 86.7$\pm$1.0 83.3$\pm$1.7 87.6$\pm$0.2 85.8$\pm$0.9 88.7$\pm$0.2 86.9$\pm$0.8 89.6$\pm$0.2 88.2$\pm$0.7 90.0$\pm$0.2
Macaroni2 61.4$\pm$3.8 66.3$\pm$0.7 67.1$\pm$5.3 67.5$\pm$0.3 70.0$\pm$1.8 68.5$\pm$0.4 74.0$\pm$1.2 70.3$\pm$0.6 76.5$\pm$1.5 71.9$\pm$0.5
PCB1 89.0$\pm$1.2 90.6$\pm$0.3 88.9$\pm$1.5 90.6$\pm$0.6 92.3$\pm$2.3 91.2$\pm$0.4 92.9$\pm$0.6 91.8$\pm$0.2 95.2$\pm$0.8 92.5$\pm$0.3
PCB2 87.8$\pm$2.0 86.5$\pm$1.1 89.4$\pm$2.5 87.5$\pm$0.2 91.8$\pm$0.7 88.4$\pm$0.1 92.8$\pm$0.6 89.1$\pm$0.2 93.0$\pm$0.4 89.5$\pm$0.3
PCB3 84.1$\pm$2.4 85.6$\pm$0.8 88.5$\pm$1.3 87.1$\pm$0.6 91.8$\pm$0.3 87.8$\pm$0.4 93.8$\pm$0.7 89.4$\pm$0.6 94.5$\pm$0.2 90.0$\pm$0.2
PCB4 76.7$\pm$11.5 79.8$\pm$5.0 86.5$\pm$7.2 83.3$\pm$3.7 94.9$\pm$3.0 86.1$\pm$0.8 97.3$\pm$0.6 88.8$\pm$1.1 99.0$\pm$0.3 90.6$\pm$0.9
Pipe fryum 91.8$\pm$2.6 97.4$\pm$0.3 93.5$\pm$3.0 97.5$\pm$0.3 94.8$\pm$2.7 97.2$\pm$0.1 96.5$\pm$1.6 97.3$\pm$0.1 98.0$\pm$0.3 97.3$\pm$0.1
Mean 86.2$\pm$1.4 85.8$\pm$0.6 88.9$\pm$1.9 86.8$\pm$0.4 91.5$\pm$0.7 87.6$\pm$0.2 92.9$\pm$0.1 88.3$\pm$0.2 94.0$\pm$0.0 89.0$\pm$0.2
The results presented in Tables 5 and 6 show that a) anomalies in some categories are more difficult to than others, b) detection and segmentation performance correlate strongly (measured in image-level AUROC and PRO), and c), that the performance of AnomalyDINO increases across the board with more available reference samples. In particular, more complex objects like ‘PCB3’ and ‘PCB4’ in VisA or ‘Transistor’ in MVTec-AD seem to benefit the most from more available reference samples. In addition, we see that also the variance in the reported metrics decreases with the number of nominal samples.
In this context, we observe a peculiarity of the few-shot regime (which does not occur for sufficiently populated $\mathcal{M}$). The results crucially depend on the chosen reference image(s). The high variances, predominantly in the 1- and 2-shot setting, for category ‘Capsule’ for MVTec-AD, or ‘Cashew’ or ‘PCB4’ for VisA demonstrate this. We discuss the potential failure cases of choosing a sub-optimal reference sample in the following section (Section B.2).
Appendix B Limitations and failure cases
The proposed method relies on similarities to patch representations captured in $\mathcal{M}$. Therefore, we can only expect the model to detect those anomalies caused by regions in the test images that are particularly different from patches of the given reference sample(s). Our experiments reveal that this may lead to some specific failure cases.
B.1 Semantic Anomalies
The first relates to the distinction between low-level sensory anomalies and high-level semantic anomalies. Semantic anomalies might be present because a logical rule or a specific semantic constraint is violated. In contrast, AnomalyDINO is developed with low-level sensory anomalies in mind.
(a) Reference sample
(b) Anomaly type ‘Bent Wire’
(c) Anomaly type ‘Cable Swap’
Figure 5: Example of a semantic anomaly in category ‘Cable’ (MVTec-AD). Depicted are a nominal reference sample, an example of a sensory anomaly (Figure 5(b)) and a semantic anomaly (Figure 5(c)) with anomaly maps predicted by AnomalyDINO (1-shot).
Consider, for instance, the anomalies shown in Figure 5. While the cable with anomaly type ‘Bent Wire’ (Figure 5(b)) contains patches that cannot be well matched to patches of the reference image, this does not apply to the anomaly ‘Cable Swap’ (Figure 5(c)). The latter shows a semantic anomaly with two blue wires, while a nominal image of a ‘Cable’ should depict all three different cable types. As a result, all test patches of Figure 5(b) can be matched well with reference patches in $\mathcal{M}$, and thus, AnomalyDINO does not detect this anomaly type. Evaluating the detection performance for this specific failure case, ‘Cable Swap’, our proposed method essentially performs on chance level, giving a detection AUROC of 50.2% ($\pm$ 4.9%).
B.2 The importance of informative reference samples
The second failure case occurs if the reference sample(s) does not resemble all concepts of normality, therefore $\mathcal{M}$ does not capture all variations of the nominal distribution $p_{\mathrm{norm}}$. As an illustration, consider the nominal samples of ‘Capsule’ in MVTec-AD, which may be rotated in such a way that the imprinted text is hidden (see Figure 6(b)). Therefore, parts of the text of a nominal test sample may be falsely recognized as anomalies. Due to the strong dependency on a suitable reference sample, we observe higher AD variances for some products in the one-shot setting, as shown in Tables 5 and 6. We like to remark, that this is only relevant to the few-shot setting, and with an increasing number of reference samples (and higher diversity of nominal patches in $\mathcal{M}$), the variance decreases notably.
(a) Reference sample (left) and estimated anomaly map of ‘good’ test sample (right).
(b) Reference sample (left) and estimated anomaly map of ‘good’ test sample (right).
Figure 6: Example of an uninformative reference sample in category ‘Capsule’ (MVTec-AD). Some reference samples do not resemble the full concept of normality (here, the sample in Figure 6(b) does not show the text on the capsule such that any nominal sample with text will get a high anomaly score). Anomaly maps predicted by AnomalyDINO based on the depicted reference sample.
(a) Test sample (left), the same test sample with ground-truth anomaly annotation (center), and a magnified view of a region with strong background artifacts (right).
(b) Anomaly map predicted by AnomalyDINO (1-shot) without (left) and with masking (right).
Figure 7: Visualization of the effect of masking in the presence of high background noise for the category ‘Capsules’ in VisA. As in Figure 4 we depict the anomaly map for ‘Capsules’ using a different colormap (red instead of yellow) for better visibility. Best viewed on a higher zoom level.
Appendix C Ablation Study
C.1 Preprocessing
As discussed in Section 3, we consider two potential preprocessing steps in our pipeline: masking and rotations. We mask out irrelevant background patches, whenever the zero-shot segmentation of DINOv2 captures the object correctly (see Figure 2). Discarding background patches helps to mitigate the problem of potential background noise, thereby reducing the number of false positives. We find that some samples of anomalous samples also show artifacts in the background. An example of the category ‘Capsules’ from VisA is depicted in Figure 7. Without masking, the method would correctly predict the depicted sample to show an anomaly, but for the wrong reason (high anomaly scores for the background noise). We conclude that masking is essential to faithfully detect anomalies whenever higher variations in the background are expected. A quantitative analysis is provided in the next paragraphs.
In addition to masking (which decreases the size of $\mathcal{M}$), we consider augmenting the reference sample with rotations (which increases the size of $\mathcal{M}$). Here we distinguish the ‘agnostic’ scenario, where we do not know a priori about potential rotations, and augment by default. Whenever we know about potential rotations of reference or test samples, we can deactivate the augmentation to reduce the test time.
The preprocessing decision per category for MVTec-AD and VisA, inferred (solely) from the first nominal reference sample in the test set (to comply with the one-shot setting), are given in Table 7.
Table 7: Default preprocessing steps for MVTec-AD and VisA. We do not mask textures, indicated by (T). In addition, we do not apply masking when the masking test failed, indicated by (MT). See Section 3 and Figure 2 for further discussion and a visualization.
MVTec-AD Mask? Rotation?
Object informed agnostic
Bottle ✗(MT)
Cable ✗(MT)
Capsule
Carpet ✗(T)
Grid ✗(T)
Hazelnut
Leather ✗(T)
Metal nut ✗(MT)
Pill
Screw
Tile ✗(T)
Toothbrush
Transistor ✗(MT)
Wood ✗(T)
Zipper ✗(MT)
VisA Mask? Rotation?
Object informed agnostic
Candle
Capsules
Cashew
Chewinggum
Fryum
Macaroni1
Macaroni2
PCB1
PCB2
PCB3
PCB4
Pipe fryum
Effect of preprocessing on detection performance
As discussed in Section 3 (and in the previous paragraph), suitable means to fill $\mathcal{M}$ influence the detection performance. The right preprocessing is in some cases more important than others. The full results are given in Figures 8 and 9.
Figure 8: Effect of preprocessing on MVTec-AD. Anomaly detection of AnomalyDINO in the 1-shot setting for different choices of the preprocessing pipeline (detection AUROC on image-level in %, mean and standard deviation over three independent runs).
Figure 9: Effect of preprocessing on VisA. Anomaly detection with AnomalyDINO in the 1-shot setting for different choices of the preprocessing pipeline (detection AUROC on image-level in %, mean and standard deviation over three independent runs).
Rotation
Increasing the diversity of nominal patch representations in $\mathcal{M}$ by rotating the reference sample can significantly improve the detection performance. Consider, e.g., the category ‘Screw’ in MVTec-AD, where the detection AUROC can be boosted from 65.6% to 90.1% in the 1-shot setting. This is intuitive, as the test samples of ‘Screw’ are taken from various angles. With sufficiently many reference samples, such data augmentation is not necessary anymore, but for the few-shot setting, we see major improvements.
Figure 10: Rotations as anomalies (‘misplaced’ transistor from MVTec-AD). The left anomaly map is estimated from a reference sample with rotations (‘agnostic’), while the right anomaly map is based on a reference sample without rotations (‘informed’).
The same holds—although to a lesser extent—for the categories ‘Hazelnut’, and ‘Metal nut’ (both MVTec-AD), as well as for ‘Macaroni2’, ‘Cashew’, ‘PCB2’ and ‘PCB4’ categories (all VisA). However, we also observe that rotations of the reference sample can actually decrease the detection performance in some categories, namely ‘PCB1’ and ‘PCB3’ (both VisA), and to a small extent also ‘Transistor’ (MVTec-AD).
We attribute this to the fact the specific anomalies for these printed circuit boards contain rotated or bent connectors (see the PCB examples in Figure 4, right column). And rotations of nominal samples may falsely reduce the distance between those patches depicting bent connectors and the nominal memory bank $\mathcal{M}$. Such a failure case based on a suboptimal preprocessing decision is depicted in Figure 10. The anomaly refers to a rotated transistor (anomaly type ‘misplaced’), and when (falsely) rotating the reference sample, such anomalies will not be detected—in contrast to the ‘informed’ preprocessing (right side of Figure 10).
Masking
Across the board, we see that masking improves the detection performance. The only exception is ‘Candle’ from VisA with a slight decrease in performance. On further inspection, we find that the estimated foreground masks of DINOv2 occasionally fail to capture all relevant patches (the decision on whether to mask or not is inferred from the first nominal sample only, 1-shot setting). We leave the development of more robust and adaptive masking techniques for future work.
Runtime analysis
The design choices in our pipeline influence the test time. To see the potential effects we measure the test time per sample on MVTec for the different combinations of masking (yes/no) and rotation (yes/no) for the full pipeline. This includes preparation of the memory bank $\mathcal{M}$ (i.e., feature extraction and the optional augmentation and masking operations), as well as the inference time for all test samples (i.e., feature extraction and optional masking). The results are given in Figure 11. The runtime was assessed utilizing a single NVIDIA A40 GPU, consistently employed throughout all experiments detailed in this paper (each experiment can be executed on a single NVIDIA A40 GPU).
The resulting run time increases by a factor of 4, when comparing the baseline (no masking, no rotation) to the scenario where masking and rotations are applied to each product. Not rotating a reference sample in cases where it is unnecessary (the difference between ‘informed’ and ‘agnostic’) lowers the test time by approximately 30%.
Figure 11: Effect on masking and augmentation on test time (per sample) in the 1-shot setting with AnomalyDINO. We measure the test time on a single NVIDIA A40 GPU and report mean and standard deviation over three independent runs on MVTec-AD.
C.2 Scoring
In Section 3 we investigate different ways of aggregating the anomaly scores $\mathcal{D}$ on patch-level (in our case, distances to the nominal memory bank $\mathcal{M}$) to an image-level anomaly score via a statistic $q$. Note that the segmentation results are therefore not affected by the choice of $q$.
A standard choice is upsampling the patch distances of lower resolution to the full resolution of the test image using bilinear interpolation, then applying Gaussian smoothing operation (we follow Roth et al., 2022, and set $\sigma=4.0$), to obtain an anomaly map $\mathcal{A}$, and set $q=\max{(\mathcal{A})}$. We also evaluate two potential alternatives of $q$, our default choice $q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ (mean of the 1% highest entries in $\mathcal{D}$) and the maximum of $\mathcal{D}$.
Table 8: Effect of aggregation statistics $q$ on the detection performance on MVTec-AD and VisA. $\mathcal{D}$ denotes the set of patch distances to the nominal memory bank, $H_{0.01}(\mathcal{D})$ the 1% highest entries thereof, and $\mathcal{A}$ denotes the anomaly map derived from $\mathcal{D}$ (anomaly scores for each pixel, after upsampling and smoothing, see Section 3). All results in %.
Scoring $q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ $q=\mathrm{max}(\mathcal{\mathcal{A}})$ $q=\mathrm{max}(\mathcal{D})$
Shots AUROC F1-max AP AUROC F1-max AP AUROC F1-max AP
MVTec-AD 1 96.6$\pm$0.5 96.0$\pm$0.2 98.2$\pm$0.3 95.8$\pm$0.6 95.4$\pm$0.3 97.7$\pm$0.4 95.1$\pm$0.5 94.8$\pm$0.1 97.4$\pm$0.4
2 96.9$\pm$0.8 96.6$\pm$0.4 98.2$\pm$0.7 96.2$\pm$1.1 96.0$\pm$0.4 97.7$\pm$0.7 95.6$\pm$1.0 95.4$\pm$0.5 97.4$\pm$0.6
4 97.7$\pm$0.1 97.1$\pm$0.3 98.5$\pm$0.3 97.2$\pm$0.2 96.7$\pm$0.3 98.1$\pm$0.4 96.7$\pm$0.2 96.1$\pm$0.1 97.9$\pm$0.4
8 98.2$\pm$0.1 97.5$\pm$0.1 99.1$\pm$0.2 97.7$\pm$0.1 97.1$\pm$0.3 98.8$\pm$0.1 97.3$\pm$0.2 96.5$\pm$0.4 98.6$\pm$0.1
16 98.4$\pm$0.1 97.8$\pm$0.2 99.3$\pm$0.0 98.0$\pm$0.1 97.4$\pm$0.2 99.1$\pm$0.1 97.7$\pm$0.2 96.9$\pm$0.3 98.9$\pm$0.1
VisA 1 86.2$\pm$1.4 83.5$\pm$1.0 87.7$\pm$1.4 85.1$\pm$1.7 82.7$\pm$1.0 87.1$\pm$1.3 83.8$\pm$1.8 82.0$\pm$1.0 85.8$\pm$1.6
2 88.9$\pm$1.9 85.3$\pm$1.4 90.4$\pm$1.4 87.7$\pm$1.9 84.3$\pm$1.4 89.7$\pm$1.2 86.6$\pm$1.9 83.4$\pm$1.2 88.5$\pm$1.2
4 91.5$\pm$0.7 87.6$\pm$0.9 92.6$\pm$0.6 90.7$\pm$0.3 86.4$\pm$0.6 92.1$\pm$0.3 89.8$\pm$0.3 85.7$\pm$0.5 91.2$\pm$0.5
8 92.9$\pm$0.1 89.0$\pm$0.3 93.8$\pm$0.1 92.0$\pm$0.1 87.8$\pm$0.3 93.2$\pm$0.3 91.2$\pm$0.1 87.0$\pm$0.2 92.4$\pm$0.4
16 94.0$\pm$0.0 90.0$\pm$0.1 94.8$\pm$0.1 93.3$\pm$0.2 88.9$\pm$0.4 94.4$\pm$0.2 92.6$\pm$0.3 88.2$\pm$0.4 93.7$\pm$0.3
The results in Table 8 show that just the maximum patch distance leads to already good results, but upsampling and smoothing the patch distances improves upon that (maximum of $\mathcal{A}$). Taking the mean of all patch distances above the 99% quantile ($q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$) improves above the standard choice. We did not optimize over the percentile and instead fixed it to 99%. Typically, the number of patches per image ranges between 200 and 1000 (depending on resolution and masking), such that 2 between 10 patches are considered. For specific products/objects and expected anomaly types, other statistics $q$ might be (more) suitable.
C.3 Architecture Size
DINOv2 is available in different distillation sizes (S, B, L, and G). This section analyzes the implications of choosing different backbone sizes for AnomalyDINO. The comparison including the best competing methods is depicted in Figure 12.
We see that different architecture sizes have some effect on the image-level AUROC. AnomalyDINO-L outperforms smaller architectures, S (our default) and B, on VisA in the few-shot setting with $k\in\{1,2,4\}$. For $k\in{8,16}$ AnomalyDINO-L is on par with AnomalyDINO-S (and APRIL-GAN). For $k=16$ all architecture sizes outperform the best competitor (APRIL-GAN).
On MVTec-AD, we see that AnomalyDINO-S performs slightly better, but the gap between AnomalyDINO-S and AnomalyDINO-B reduces with increasing shot size. In particular, all architecture sizes outperform the closest competitor (ADP).
Refer to captionRefer to caption
Figure 12: Effect of model size on detection AUROC. We report mean and standard deviation for the 1-shot setting on MVTec-AD (left) and VisA (right) over three seeds. Note, that the results for WinCLIP+ for 8 and 16 shots are those of the WinCLIP re-implementation by Kwak et al. (2024).
Appendix D Extending AnomalyDINO to the Batched Zero-Shot Setting
We extend the proposed method to the batched zero-shot setting. Recall, that in this setting all test samples $X_{\mathrm{test}}$ are provided (or at least a sufficiently large batch), but no (labeled) training or reference samples. The underlying (and necessary) assumption to meaningfully predict anomalies solely based on test samples, is that the majority of samples (or in our case, patches) at test time are from the nominal data distribution (see e.g., Li et al., 2024a).
We need to alter our method, outlined in Section 3, only slightly to adapt it to the batched setting. We score a test sample $\mathbf{x}^{(j)}\in X_{\mathrm{test}}$ in comparison to all remaining test sample $X_{\operatorname{test}}\setminus\{\mathbf{x}^{(j)}\}$, following the idea of mutual scoring (Li et al., 2024b). For each test sample $\mathbf{x}^{(j)}\in X_{\mathrm{test}}$ we therefore collect all patch representations not belonging to $\mathbf{x}^{(j)}$ in a memory bank, again utilizing DINOv2 as patch-level feature extractor $f$,
$\mathcal{M}_{j}:=\bigcup_{\mathbf{x}^{(i)}\in X_{\operatorname{test}}\setminus\{\mathbf{x}^{(j)}\}}\bigl{\{}\mathbf{p}_{m}\;|\;f(\mathbf{x}^{(i)})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n}),\,m\in\{1,\dots,n\}\bigr{\}}\enspace.$ (5)
We need to infer anomaly scores for each patch representation $\mathbf{p}_{\mathrm{test}}$ of $\mathbf{x}^{(j)}$. We could again assess the distances between $\mathbf{p}_{\mathrm{test}}$ and $\mathcal{M}_{j}$ based on the distance to the nearest neighbor, as done in Equation 2. Note, however, that $\mathcal{M}_{j}$ may now contain nominal and anomalous patches. Thus, the nearest-neighbor approach is not suitable anymore: the nearest neighbor in $\mathcal{M}_{j}$ for (the representation of) an anomalous patch might also be abnormal, and the resulting distance therefore not informative.
A simple solution, based on the assumption that the majority of patches are nominal, is to replace the nearest neighbor with a suitable aggregation statistic over the distribution of patches distances in $\mathcal{M}_{j}$
$\mathcal{D}(\mathbf{p}_{\mathrm{test}},\mathcal{M}_{j}):=\big{\{}d(\mathbf{p}_{\mathrm{test}},\mathbf{p})\mid\mathbf{p}\in\mathcal{M}_{j}\big{\}}\enspace,$ (6)
where we again use the cosine distance $d$, defined in Equation 3. Specifically, we can again make use of the tail value at risk—now for the lowest quantile as we are interested in the tail behavior of the distribution of distances to the nearest neighbors—to derive a patch-level anomaly score
$s(\mathbf{p}_{\mathrm{test}}):=\mathrm{mean}\big{(}L_{\alpha}\big{(}\mathcal{D}(\mathbf{p}_{\mathrm{test}},\mathcal{M}_{j})\big{)}\big{)}\enspace,$ (7)
where $L_{\alpha}(\mathcal{D})$ contains the values below the $\alpha$ quantile of $\mathcal{D}$. We set $\alpha=0.1\%$ as anomalous patches are rare by assumption, and $\mathcal{D}(\mathbf{p}_{\mathrm{test}},\mathcal{M}_{j})$ large enough to accurately estimate the tail of the distribution. [^6] For MVTec-AD the total number of test patches extracted by DINOv2 (ViT-S) range between 43.008 and 171.008 per category, and for VisA between 163.200 and 334.464. The image-level score $s(\mathbf{x}_{\mathrm{test}})$ is again given by aggregating the patch-level anomaly score following Equation 4. Computation of the cosine distances and the proposed aggregation statistics can be effectively implemented as matrix operations on GPU such that the batched zero-shot inference time for AnomalyDINO amounts to roughly 0.1 seconds/sample for MVTec-AD (again measured on an NVIDIA A40; feature extraction takes approximately the same amount of time).
Some resulting anomaly maps in the batched zero-shot setting are depicted in Figure 13.
Figure 13: Anomaly maps for the batched zero-shot setting on MVTec-AD. The left-most sample in each category is a ‘good’ test sample for reference, followed by four samples with anomalies.

### v2

@@
 Despite its simplicity, AnomalyDINO achieves state-of-the-art results in one- and few-shot anomaly detection (e.g., pushing the one-shot performance on MVTec-AD from an AUROC of $93.1\%$ to $96.6\%$).
-The reduced overhead, coupled with its outstanding few-shot performance, makes AnomalyDINO a strong candidate for fast deployment, for example, in industrial contexts.
+The reduced overhead, coupled with its outstanding few-shot performance, makes AnomalyDINO a strong candidate for fast deployment, e.g., in industrial contexts.
 1 Introduction
 Anomaly detection (AD) in machine learning attempts to identify instances that deviate substantially from the nominal data distribution $p_{\mathrm{norm}}(x)$.
-Anomalies, therefore, raise suspicion of being ‘generated by a different mechanism’ (Hawkins, 1980)—often indicating critical, rare, or unforeseen events.
-The ability to reliably distinguish anomalies from normal samples is highly valuable across various domains, including security (Siddiqui et al., 2019), healthcare (Fernando et al., 2021; Tian et al., 2023), and industrial inspection.
+Anomalies, therefore, raise suspicion of being ‘generated by a different mechanism’ [15]—often indicating critical, rare, or unforeseen events.
+The ability to reliably distinguish anomalies from normal samples is highly valuable across various domains, including security [38], healthcare [13, 39], and industrial inspection.
 In this work, we focus on the latter, where fully automated systems necessitate the ability to detect defective or missing parts to prevent malfunctions in downstream products, raise alerts for potential hazards, or analyze these to optimize production lines.
-See the right-hand side of Figure 1 for anomalous samples in this context (with predicted anomaly maps from AnomalyDINO).
+See the right-hand side of Figure 1 for anomalous samples in this context.
 Figure 1: Anomaly detection with AnomalyDINO based on a single immaculate reference sample (here category ‘Screw’ from MVTec-AD).
-Whenever suitable we select the relevant patches from the (potentially augmented) reference sample via masking and collect the nominal patch representations in $\mathcal{M}$.
+We collect the nominal patch representations from the (potentially augmented) reference sample(s) in the memory bank $\mathcal{M}$.
+At test time, we select the relevant patch representation via masking (if applicable).
+The distances of those to the nominal representations in $\mathcal{M}$ give rise to an anomaly map and the corresponding anomaly score $s(\mathbf{x}_{\mathrm{test}})$ using the aggregation statistic $q$.
 For both, masking and feature extraction, we utilize DINOv2.
-At test time, the distances of patch representations to those in $\mathcal{M}$ give rise to an anomaly map and the corresponding anomaly score after aggregation with $q$.
-Further examples for other categories are depicted on the right (and in Figures 3 and 4 in Appendix A).
+Further examples for other categories are depicted on the right (and in Figures 4 and 5 in Appendix A).
 AD for industrial images has gained tremendous interest over the last couple of years.
@@
 For instance, Mousakhan et al.
-(2023) report $99.8\%$ and $98.9\%$ AUROC on the popular benchmarks MVTec-AD (Bergmann et al., 2021) and VisA (Zou et al., 2022), respectively.
-The most popular AD techniques use the training data to train an anomaly classifier (e.g., Ruff et al., 2020), or a generative model coupled with reconstruction-based (e.g., Wolleb et al., 2022; Livernoche et al., 2023; Mousakhan et al., 2023) or likelihood-based (e.g., Rudolph et al., 2022; Daehyun Kim, 2023) anomaly scoring.
+[27] report $99.8\%$ and $98.9\%$ AUROC on the popular benchmarks MVTec-AD [2] and VisA [48], respectively.
+The most popular AD techniques use the training data to train an anomaly classifier [36], or a generative model coupled with reconstruction-based [40, 25, 27], or likelihood-based [34, 9] anomaly scoring.
 However, these approaches operate within the full-shot setting, meaning they rely on access to a sufficiently large amount of training data.
-Given the challenges associated with dataset acquisition, the attractivity of a fast and easy-to-deploy methodology, and the requirement to rapidly adapt to covariate shifts in the nominal data distribution (Li et al., 2024a), there is an increasing interest in few-shot and zero-shot anomaly detection.
-Few-shot techniques, however, heavily rely on meaningful features, or as Reiss et al.
-(2022) frame it: ‘anomaly detection requires better representations’.
-Such better representations are now available with the increasing availability and capabilities of foundation models, i.e., large-scale models trained on massive datasets in unsupervised/self-supervised fashion (e.g., Radford et al., 2021; Caron et al., 2021; Oquab et al., 2024).
-The performance of few-shot anomaly detection techniques has already been boosted by the use of foundation models, mostly by multimodal approaches (e.g., Jeong et al., 2023; Cao et al., 2023; Zhou et al., 2024; Kwak et al., 2024)
+Given the challenges associated with dataset acquisition, the attractivity of a fast and easy-to-deploy methodology, and the requirement to rapidly adapt to covariate shifts in the nominal data distribution [21], there is an increasing interest in few-shot and zero-shot anomaly detection.
+Few-shot techniques, however, heavily rely on meaningful features, or as [32] frame it: ‘anomaly detection requires better representations’.
+Such better representations are now available with the increasing availability and capabilities of foundation models, i.e., large-scale models trained on massive datasets in unsupervised/self-supervised fashion [30, 5, 29].
+The performance of few-shot anomaly detection techniques has already been boosted by the use of foundation models, mostly by multimodal approaches incorporating language and vision[18, 4, 47, 20].
 Here, we propose to focus on a vision-only approach, in contrast to such multimodal techniques.
 This perspective is motivated by the observation that few-shot anomaly detection is feasible for human annotators based on visual features only, and does not require additional textual description of the given object or the expected types of anomalies (which are typically not known a priori).
-Our approach, termed AnomalyDINO, is memory-bank-based and leverages DINOv2 (Oquab et al., 2024) as a backbone (but any architecture with strong patch-level feature extraction is suitable for the proposed approach).
+Our approach, termed AnomalyDINO, is memory bank-based and leverages DINOv2 [29] as a backbone (but the approach could be adapted to other architectures with strong patch-level feature extraction).
 We carefully design a suitable preprocessing pipeline for the one-shot scenario, using the segmentation abilities of DINOv2 (which alleviates the additional overhead of another model).
 At test time, anomalous samples are detected based on the high distances between their patch representations and the closest counterparts in the nominal memory bank.
-Due to its simplicity, AnomalyDINO can be deployed in industrial contexts very easily—in strong contrast to more complex approaches such as those proposed by Chen et al.
-(2023) or Kwak et al.
-(2024).
-Yet, the proposed method achieves new state-of-the-art performance on anomaly detection in the few-shot regime on MVTec-AD (Bergmann et al., 2021) and outperforms all but one competing method on VisA (Zou et al., 2022).
+Due to its simplicity, AnomalyDINO can be deployed in industrial contexts very easily—in strong contrast to more complex approaches such as those proposed by [7] or [20].
+Yet, the proposed method achieves new state-of-the-art performance on anomaly detection in the few-shot regime on MVTec-AD [2] and outperforms all but one competing method on VisA [48].
 The structure of the paper is as follows: Section 2 reviews relevant prior studies and clarifies the distinctions between the settings addressed by zero- and few-shot, and batched zero-shot techniques.
@@
 We address identified failure cases of AnomalyDINO in Appendix B.
+The code to reproduce the experiments is available at https://github.com/dammsi/AnomalyDINO.
 Contributions
@@
 Our method builds upon the high-quality feature representation extracted by DINOv2.
-  • An extensive analysis demonstrates the efficiency and effectiveness of the proposed approach, outperforming other multimodal few-shot techniques, thereby reducing the gap between few- and full-shot anomaly detection.
-Specifically, AnomalyDINO achieves state-of-the-art results for few-shot anomaly detection on MVTec-AD, e.g., pushing the one-shot detection from an AUROC of 93.1% to 96.6%.
-Moreover, our results on VisA are not only competitive with other few-shot methods but also establish new state-of-the-art performance among all training-free few-shot anomaly detectors.
+  • An extensive analysis demonstrates the efficiency and effectiveness of the proposed approach, outperforming other multimodal few-shot techniques in terms of performance and inference speed.
+Specifically, AnomalyDINO achieves state-of-the-art results for few-shot anomaly detection on MVTec-AD, e.g., pushing the one-shot detection from an AUROC of 93.1% to 96.6% (thereby halving the gap between the few- and full-shot setting).
+Moreover, our results on VisA are not only competitive with other few-shot methods but also establish a new state-of-the-art for all training-free few-shot anomaly detectors, achieving the best localization performance across all methods.
 2 Related Work
 Foundation Models for Vision
-Multimodal foundation models have emerged as powerful tools for a wide range of tasks (see e.g., Bommasani et al., 2022; Li et al., 2023; Chen et al., 2020; He et al., 2022; Radford et al., 2021; Liu et al., 2023; OpenAI, 2024).
-Most relevant to visual AD are multimodal approaches based on CLIP (Radford et al., 2021) or recent LLMs (e.g., OpenAI, 2024), but also vision-only approaches like DINO (Caron et al., 2021; Oquab et al., 2024).
-CLIP (Radford et al., 2021) learns visual concepts from natural language descriptions by training on a dataset of images paired with textual annotations.
+Multimodal foundation models have emerged as powerful tools for a wide range of tasks, see e.g., [3, 22, 6, 16, 30, 24, 28].
+Most relevant to visual AD are multimodal approaches based on CLIP [30] or recent LLMs [28], but also vision-only approaches like DINO [5, 29].
+CLIP [30] learns visual concepts from natural language descriptions by training on a dataset of images paired with textual annotations.
 The model uses a contrastive learning objective that aligns the embeddings from image and text encoders, optimizing the similarity between corresponding image-text pairs.
 This common feature space for vision and language can be utilized for several downstream tasks, such as zero-shot image classification by assessing similarities to a set of class-specific prompts.
-DINO (Caron et al., 2021; Oquab et al., 2024) leverages a self-supervised student-teacher framework based on vision transformers (Dosovitskiy et al., 2021).
+DINO [5, 29] leverages a self-supervised student-teacher framework based on vision transformers [12].
 It employs a multi-view strategy to train models to predict softened teacher outputs, thereby learning robust and high-quality features for downstream tasks.
-DINOv2 (Oquab et al., 2024) combines ideas from DINO with patch-level reconstruction techniques (Zhou et al., 2022) and scales to larger architectures and datasets.
+DINOv2 [29] combines ideas from DINO with patch-level reconstruction techniques [46] and scales to larger architectures and datasets.
 The features extracted by DINO are well-suited for anomaly detection as they incorporate both local and global information, are robust to multiple views and crops, and benefit from large-scale pre-training.
-GroundingDINO (Liu et al., 2023), builds upon the DINO framework and focuses on improving the alignment of textual and visual information, enhancing the model’s performance in tasks requiring detailed object localization and multimodal understanding.
+GroundingDINO [24], builds upon the DINO framework and focuses on improving the alignment of textual and visual information, enhancing the model’s performance in tasks requiring detailed object localization and multimodal understanding.
 Anomaly Detection
-Given a predefined notion of normality, the anomaly detection task is to detect test samples that deviate from this concept (Yang et al., 2021; Ruff et al., 2021).
+Given a predefined notion of normality, the anomaly detection task is to detect test samples that deviate from this concept [43, 35].
 In this work, we focus on low-level sensory anomalies of industrial image data, i.e., we do not target the detection of semantic anomalies but of low-level features such as scratches of images of industrial products (see e.g., Figure 1).
-Several works tackled this task by either training an anomaly classifier (e.g., Ruff et al., 2020) or a generative model, which allows for reconstruction-based or likelihood-based AD (e.g., Wolleb et al., 2022; Livernoche et al., 2023; Zhang et al., 2023; Daehyun Kim, 2023; Mousakhan et al., 2023).
+Several works tackled this task by either training an anomaly classifier [36] or a generative model, which allows for reconstruction-based or likelihood-based AD [40, 25, 45, 9, 27].
 However, this work focuses on few-shot AD, which tackles anomaly detection under strong restrictions on the number of available training samples.
 Various recent methods recognized the effectiveness of pre-trained vision models for few-shot AD.
-The underlying idea in these approaches is to leverage the feature representations originating from pre-trained ResNets (He et al., 2015), Wide ResNets (Zagoruyko and Komodakis, 2016), and Vision Transformers (ViT, Dosovitskiy et al., 2020).
+The underlying idea in these approaches is to leverage the feature representations originating from pre-trained ResNets [17], Wide ResNets [44], and Vision Transformers (ViT) [11].
 These models can extract representations on a patch and pixel level, which can be compared against a memory bank, which aggregates patch and pixel features provided by the training set.
-The comparison with the memory bank is typically conducted using a nearest neighbor approach (Bergman et al., 2020; Cohen and Hoshen, 2020; Defard et al., 2020; Roth et al., 2022; Xie et al., 2023; Li et al., 2024b).
+The comparison with the memory bank is typically conducted using a nearest neighbor approach [1, 8, 10, 33, 41, 23].
 Loosely speaking, patches and pixels corresponding to anomalies are expected to have a high distance to their nearest neighbor in the memory bank.
@@
 First, these approaches define sets of prompts describing nominal samples and anomalies.
-Second, the corresponding textual embeddings are compared against the image embeddings (Jeong et al., 2023; Chen et al., 2023; Kwak et al., 2024; Zhou et al., 2024).
+Second, the corresponding textual embeddings are compared against the image embeddings [18, 7, 20, 47].
 Images whose visual embedding is close to the textual embedding of a prompt associated with an anomaly are classified as anomalous.
-However, these methods either require significant prompt engineering (e.g., Chen et al., 2023, use a total of $35\times 7$ different prompts for describing normal samples) or fine-tuning of the prompt(-embeddings).
+However, these methods either require significant prompt engineering (e.g., [7] use a total of $35\times 7$ different prompts for describing normal samples) or fine-tuning of the prompt(-embeddings).
 Lastly, another type of few-shot anomaly detection builds upon the success of multimodal chatbots.
-These methods require more elaborate prompting and techniques for interpreting textual outputs (Xu et al., 2024).
+These methods require more elaborate prompting and techniques for interpreting textual outputs [42].
 Since these methods do not require a memory bank, they are capable of performing zero-shot anomaly detection.
+Table 1: Taxonomy of recent few- and zero-shot anomaly detection methods.
+The $\dagger$ indicates approaches that were introduced as full-shot detectors but then considered as few-shot detectors in later works, see e.g., [33].
+Method Setting Training Type Modes
+DN2 [1] Few-Shot Training-Free Vision
+SPADE [8] Few-Shot Training-Free Vision
+PaDiM [10] Few-Shot Fine-Tuning Vision
+PatchCore [33] Few-Shot Training-Free Vision
+PatchCore–opt [37] Few-Shot Training-Free Vision
+MuSc [23] Batched Zero-Shot Training-Free Vision
+GraphCore [41] Few-Shot Training (GNN) Vision
+WinCLIP [18] Zero-/Few-Shot Training-Free Vision + Language
+APRIL-GAN [7] Zero-/Few-Shot Meta-Training Vision + Language
+ADP [20] Few-Shot Fine-Tuning Vision + Language
+AnomalyCLIP [47] Zero-Shot Meta-Training Vision + Language
+GPT4-V [42] Zero-Shot Training-Free Vision + Language
+ACR [21] Batched Zero-Shot Meta-Training /
+AnomalyDINO (Ours) Few-Shot Training-Free Vision
 Categorization of Few-/Zero-Shot Anomaly Detectors
@@
 In batched zero-shot, inference is not performed sample-wise but based on a whole batch of test samples, usually the full test set.
-For instance, the method proposed by Li et al.
-(2024b) benefits from the fact that a significant majority of pixels correspond to normal pixels, which motivates the strategy of matching patches across a batch of images.
-Another work that considers this setting is that of Li et al.
-(2024a), who deploy a parameter-free anomaly detector based on the effect of batch normalization.
+For instance, the method proposed in [23] benefits from the fact that a significant majority of pixels correspond to normal pixels, which motivates the strategy of matching patches across a batch of images.
+Another work that considers this setting [21], deploys a parameter-free anomaly detector based on the effect of batch normalization.
 We split the training requirements into the categories ‘Training-Free’, ‘Fine-Tuning’, and ‘Meta-Training’.
@@
 In contrast, ‘Meta-Training’ is associated with training the model on a dataset related to the test data.
-For example, Li et al.
-(2024a) train their model on MVTec-AD containing all classes except the class they test against.
-Zhou et al.
-(2024) and Chen et al.
-(2023) train their model on VisA when evaluating the test performance on MVTec-AD and vice versa.
+For example, [21] train their model on MVTec-AD containing all classes except the class they test against.
+[47] and [7] train their model on VisA when evaluating the test performance on MVTec-AD and vice versa.
 Finally, we differentiate the leveraged models, which are either vision models (such as pre-trained ViT) or language-vision models (such as CLIP).
 We provide a detailed summary in Table 1.
-Table 1: Taxonomy of recent few- and zero-shot anomaly detection methods.
-The $\dagger$ indicates approaches that were introduced as full-shot detectors but then considered as few-shot detectors in later works, see e.g., Roth et al.
-(2022).
-Method Setting Training Type Modes
-DN2 (Bergman et al., 2020) Few-Shot Training-Free Vision
-SPADE (Cohen and Hoshen, 2020) Few-Shot Training-Free Vision
-PaDiM (Defard et al., 2020) Few-Shot Fine-Tuning Vision
-PatchCore (Roth et al., 2022) Few-Shot Training-Free Vision
-PatchCore–opt (Santos et al., 2023) Few-Shot Training-Free Vision
-MuSc (Li et al., 2024b) Batched Zero-Shot Training-Free Vision
-GraphCore (Xie et al., 2023) Few-Shot Training (GNN) Vision
-WinCLIP (Jeong et al., 2023) Zero-/Few-Shot Training-Free Vision + Language
-APRIL-GAN (Chen et al., 2023) Zero-/Few-Shot Meta-Training Vision + Language
-ADP (Kwak et al., 2024) Few-Shot Fine-Tuning Vision + Language
-AnomalyCLIP (Zhou et al., 2024) Zero-Shot Meta-Training Vision + Language
-GPT4-V (Xu et al., 2024) Zero-Shot Training-Free Vision + Language
-ACR (Li et al., 2024a) Batched Zero-Shot Meta-Training /
-AnomalyDINO (Ours) Few-Shot Training-Free Vision
-3 Patch-Matching for Visual Anomaly Detection
+3 Matching Patch Representations for Visual Anomaly Detection
 This section introduces AnomalyDINO, which leverages DINOv2 to extract meaningful patch-level features.
 The strong feature extraction capabilities of the utilized backbone allow us to fall back on a simple patch-based matching strategy to perform few-shot and training-free AD for image data.
-We follow a memory-bank approach, i.e., we first gather relevant patch representations of nominal samples in a memory bank $\mathcal{M}$.
+We first gather relevant patch representations of nominal samples in a memory bank $\mathcal{M}$.
 Then, for each test patch, we compute its distance to the nearest nominal patch in $\mathcal{M}$.
 A suitable aggregation of the patch-based distances gives the anomaly score on image-level.
-3.1 Anomaly Detection with Nearest Patch Distances
-Let us consider the image space $\mathcal{X}$.
-We assume to have access to a suitable feature extractor $f:\mathcal{X}\rightarrow\mathcal{F}^{n}$ that maps each image $\mathbf{x}$ to a tuple of patch-features $f(\mathbf{x})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n})$, where $\mathcal{F}$ denotes the feature space.
+Inspired by PatchCore [33], our method simplifies the memory bank framework by selecting DINOv2 as the backbone, thereby reducing the complexity of the feature engineering stage—specifically in deciding which representations or layers to use and how to aggregate them.
+We also tailor the memory bank concept to the few-shot learning regime by introducing an improved aggregation statistic, and a new preprocessing pipeline that incorporates zero-shot masking and augmentations.
+The proposed method is described in detail in the following subsections.
+3.1 Anomaly Detection via Patch (Dis-) Similarities
+Let us briefly review the idea of the memory bank approach.
+We assume to have access to a suitable feature extractor $f:\mathcal{X}\rightarrow\mathcal{F}^{n}$ that maps each image $\mathbf{x}\in\mathcal{X}$ to a tuple of patch-features $f(\mathbf{x})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n})$, where $\mathcal{X}$ denotes some space of images and $\mathcal{F}$ the feature space.
 Note that $n$ depends on the image resolution and the patch size.
-Given $k\geq 1$ nominal reference sample(s) $X_{\text{ref}}:=\bigl{\{}\mathbf{x}^{(i)}\mid\,i\in\{1,\dots,k\}\bigr{\}}$ ($k$-shot setting), we collect the nominal patch features and store them in a memory bank
-$\mathcal{M}:=\bigcup_{\mathbf{x}^{(i)}\in X_{\operatorname{ref}}}\bigl{\{}\mathbf{p}_{j}^{(i)}\;|\;f(\mathbf{x}^{(i)})=(\mathbf{p}_{1}^{(i)},\dots,\mathbf{p}_{n}^{(i)}),\,j\in\{1,\dots,n\}\bigr{\}}\enspace.$ (1)
-To score a test sample $\mathbf{x}_{\mathrm{test}}$, we collect the extracted patch representations $f(\mathbf{x}_{\mathrm{test}})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n})$ and then check how well they comply with our memory bank $\mathcal{M}$.
+Given $k\geq 1$ nominal reference sample(s) $X_{\text{ref}}:=\bigl{\{}\mathbf{x}^{(i)}\mid\,i\in[k]\bigr{\}}$ (with shorthand $[k]:=\{1,\dots,k\}$), we collect the nominal patch features and store them in a memory bank
+$\begin{split}\mathcal{M}:=\bigcup_{\mathbf{x}^{(i)}\in X_{\operatorname{ref}}}\bigl{\{}\mathbf{p}_{j}^{(i)}\mid f(\mathbf{x}^{(i)})=(\mathbf{p}_{1}^{(i)},\dots,\mathbf{p}_{n}^{(i)}),\,j\in[n]\bigr{\}}\enspace.\end{split}$ (1)
+To score a test sample $\mathbf{x}_{\mathrm{test}}$, we collect the extracted patch representations $f(\mathbf{x}_{\mathrm{test}})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n})$ and then check how well they comply with $\mathcal{M}$.
 To do so, we leverage a nearest-neighbor approach to find the distance to the closest reference patch for a given test patch $\mathbf{p}\in\mathcal{F}$
@@
 Throughout this paper, we define $q$ as the average distance of the 1% most anomalous patches, i.e, $q(\mathcal{D}):=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ with $H_{0.01}(\mathcal{D})$ containing the 1% highest values in $\mathcal{D}$.
-The statistic $q$ can be understood as an empirical estimate of the tail value at risk for the 99% quantile (e.g., McNeil, 1999) and turns out suitable for a wide range of settings because it balances two desirable properties: First, we want $s(\mathbf{x}_{\mathrm{test}})$ to depend on the highest patch distances, as they may provide the strongest anomaly signal.
+The statistic $q$ can be understood as an empirical estimate of the tail value at risk for the 99% quantile [26] and turns out suitable for a wide range of settings because it balances two desirable properties: First, we want $s(\mathbf{x}_{\mathrm{test}})$ to depend on the highest patch distances, as they may provide the strongest anomaly signal.
 Similarly, we want a certain degree of robustness against a singular high patch distance (in particular in few-shot scenarios where $\mathcal{M}$ is sparsely populated).
@@
 Such an anomaly score on pixel-level is usually obtained by upsampling to full image resolution and applying some smoothing operation.
-Following Roth et al.
-(2022), we utilize bilinear upsampling and Gaussian smoothing (with $\sigma=4.0$) to turn the patch distances into pixel-level anomaly scores for localization of potential defects.
-Examples of the resulting anomaly maps are visualized in Figure 1 and Appendix A (Figures 3 and 4).
+Following [33], we utilize bilinear upsampling and Gaussian smoothing ($\sigma=4.0$) to turn the patch distances into pixel-level anomaly scores for localization of potential defects.
+Examples of the resulting anomaly maps are visualized in Figure 1 and Appendix A (Figures 4 and 5).
 We extend AnomalyDINO also to the batched zero-shot scenario, see Appendix D (Figure 15).
 3.2 Enriching the Memory Bank & Filtering Relevant Patches
-Access to only a single/few normal sample(s) poses (at least) two challenges.
-First, all relevant information about the concept of normality must be extracted from the single/few available nominal sample(s).
-However, it is expected that the variability at test time is (considerably) higher than in $X_{\mathrm{ref}}$.
-Secondly, one must determine a suitable preprocessing for reference and test samples, solely based on $X_{\mathrm{ref}}$.
-To address low variations of nominal patch features, we apply simple augmentations in the form of rotations, following the insights of Xie et al.
-(2023).
-In addition, to ensure that $\mathcal{M}$ is of high quality, we consider masking the object of interest as background noise could poison $\mathcal{M}$.
-Note, that a suitable preprocessing, i.e., whether to apply masking or rotations, should depend on the object of interest (see, e.g., Figures 7 and 10, and Section C.1 for further discussion).
-Figure 2: Masking test on MVTec-AD. The dashed red regions mark areas that belong to the object of interest but are falsely predicted to be background. See App. C.1 for the outcomes per category.
+In the few-shot anomaly detection setting, the primary challenge is to effectively capture the concept of normality from the limited set of nominal samples.
+A useful strategy involves applying simple augmentations, like rotations (following the insights in [41]), to enhance the variability of nominal patch features.
+This is essential because the variability at test time is likely to be significantly greater than in the reference data, $X_{\mathrm{ref}}$.
+To avoid irrelevant areas of the test image from leading to falsely high anomaly scores, we propose masking the object of interest.
+This approach mitigates the risk of false positives, particularly in the few-shot regime where the limited reference samples may not adequately capture the natural variations in the background, as exemplified in Figure 8, Appendix C.
+It is important to note that the appropriate preprocessing technique—whether to apply masking and/or augmentations—should depend on the specific characteristics of the object of interest.
+For a more detailed discussion on the challenges and considerations in designing an effective preprocessing pipeline, see Section C.1.
 Masking
-To minimize the overhead of the proposed method, we utilize DINOv2 also for masking, i.e., delineating the primary object(s) in an image from its background.
-This is achieved by thresholding the first PCA component of the patch features (Oquab et al., 2024).
-We observe empirically that this procedure sometimes produces erroneous masks.
-Frequently, such failure cases occur for close-up shots, where the objects of interest account for $\gtrsim 50\%$ of patches, see Figure 2.
-To resolve that issue we test whether the PCA-based mask correctly captures the object in the first reference sample and apply masking accordingly.
-In general, we do not mask textures (e.g., wood or tile in MVTec-AD).
+Masking, i.e., delineating the primary object(s) in an image from its background, can help to reduce false-positive predictions.
+To minimize the overhead of the proposed pipeline, we utilize DINOv2 also for masking.
+This is achieved by thresholding the first PCA component of the patch features [29].
+We observe empirically that this sometimes produces erroneous masks.
+Frequently, such failure cases occur for close-up shots, where the objects of interest account for $\gtrsim 50\%$ of patches.
+To address this issue, we check if the PCA-based mask accurately captures the object in the first reference sample and apply the mask accordingly, which gives rise to the ‘masking test’ in Figure 2.
+This test is performed only once per object as the procedure yields very consistent outputs.
+In addition, we utilize dilation and morphological closing to eliminate small holes and gaps within the predicted masks.
+See Figures 12 and 13 for examples of this masking procedure, Table 8 for the outcomes of the masking test per object, and Figure 8 for a visualization of the benefits of masking in the presence of background noise (all in Appendix C).
+In general, we do not mask textures (e.g., ‘Wood’ or ‘Tile’ in MVTec-AD).
+Figure 2: Masking test on MVTec-AD.
+For ‘Capsule’ and ‘Hazelnut’ the masking works successfully (top row), while for ‘Cable’ and ‘Transistor’ (bottom row) some areas are incorrectly predicted as background that should belong to the object of interest (highlighted in red).
+See App.
+C.1 for the outcomes per object.
 Rotation
-Rotating the reference sample may improve the detection performance by enriching the memory bank, but can also be detrimental in cases where rotations of (parts of the) object of interest can be considered anomalies themselves.
+Rotating the reference sample may improve the detection performance by better resembling the variations within the concept of normality captured in $\mathcal{M}$.
+Consider, e.g., the ‘Screw’ in MVTec-AD as depicted in Figure 1 for which rotation-invariant features are desirable.
+On the other hand, rotation can also be detrimental in cases where rotations of (parts of) the object of interest can be considered anomalies themselves (see Figure 11 for an example).
 We consider two different settings.
@@
 We also consider the case where we know about the samples’ potential rotations (‘informed’).
-The ‘informed’ case is sensible as the data collection process can usually be controlled in an industrial/medical setting (or test images can be aligned), and may lead to significantly lower inference time.
+The ‘informed’ case is sensible as the data collection process can usually be controlled in an industrial/medical setting (or test images can be aligned), and may lead to lower inference time.
 See Section C.1 for the effect of masking and rotation, and a comparison of ‘informed’ vs.
@@
 AnomalyDINO – Defaults
-Using DINOv2111Code and model weights of DINOv2 (Oquab et al., 2024) are available under Apache 2.0 license.
-as backbone allows us to choose from the distillation sizes, which range from small (ViT-S, $21\times 10^{6}$ parameters) to giant (ViT-G, $1.1\times 10^{9}$ parameters).
-To prioritize low latency, we take the smallest model (S) as our default.
-As discussed above, the ‘agnostic’ preprocessing is our default.
+Using DINOv2111Code/weights of DINOv2 [29] are available under Apache 2.0 license.
+as backbone allows us to choose from different distillation sizes, which range from small (ViT-S, $21\times 10^{6}$ parameters) to giant (ViT-G, $1.1\times 10^{9}$ parameters).
+To prioritize low latency, we take the smallest model as our default (and denote our default pipeline AnomalyDINO-S, accordingly) and evaluate two input resolutions, 448 and 672 pixels (smaller edge).
+As discussed above, we utilize the ‘agnostic’ preprocessing by default (see Appendix C).
 Datasets
-For the experiments, we consider MVTec-AD [^2] The MVTec-AD dataset (Bergmann et al., 2021) is available under the CC BY-NC-SA 4.0 license.
-and VisA, [^3] The VisA dataset (Zou et al., 2022) is released under the CC BY 4.0 license.
+For the experiments, we consider MVTec-AD [^2] The MVTec-AD dataset [2] is available under CC BY-NC-SA 4.0.
+and VisA, [^3] The VisA dataset [48] is released under the CC BY 4.0 license.
 two datasets with high-resolution images for industrial anomaly detection.
-MVTec-AD consists of fifteen categories, both objects and textures, with a small amount of clean training samples (between 60 and 391), and up to eight anomaly types per category.
-VisA has twelve categories with up to 1006 normal samples per category and diverse types of anomalies condensed into one anomaly class ‘bad’.
+MVTec-AD consists of fifteen categories, depicting single objects or textures, and up to eight anomaly types per category.
+VisA has twelve categories and comes with diverse types of anomalies condensed into one anomaly class ‘bad’.
+Some categories in VisA can be considered more challenging as multiple objects and complex structures are present.
 Evaluation Metrics
-We assess the image-level detection and the pixel-level segmentation performance with three metrics each (similar to, e.g., Jeong et al., 2023; Chen et al., 2023; Li et al., 2024b).
+We assess the image-level detection and the pixel-level segmentation performance with three metrics each, following [18, 7, 23].
 For evaluating the detection performance, we measure the area under the receiver-operator curve (AUROC), the F1-score at the optimal threshold (F1-max), and the average precision (AP) using the respective image-level anomaly scores.
-We quantify the anomaly segmentation performance using the AUROC, F1-max, and the per-region overlap (PRO, Bergmann et al., 2021) of the segmentation using the pixel-wise anomaly scores.
-Note in this context that due to a high imbalance of nominal and anomalous pixels—for MVTec-AD we have 97.26% nominal pixel, for VisA even 99.45% (Li et al., 2024b)—it is not recommended to assess performance solely on segmentation AUROC (Rafiei et al., 2023).
+We quantify the anomaly segmentation performance using the AUROC, F1-max, and the per-region overlap (PRO, [2]) of the segmentation using the pixel-wise anomaly scores.
+Note that due to a high imbalance of nominal and anomalous pixels—for MVTec-AD we have 97.26% nominal pixel, for VisA even 99.45% [23]—it is not recommended to assess performance solely on segmentation AUROC [31].
 We repeat each experiment three times and report mean and standard deviation.
@@
 For straightforward reproducibility, we simply set $X_{\mathrm{ref}}$ to the first $k$, second $k$ and third $k$ train samples for the three different runs, respectively.
+Baselines
+We compare AnomalyDINO with a range of modern zero- and few-shot AD models, e.g., SPADE [8], PaDiM [10], PatchCore [33], WinCLIP+ [18], and APRIL-GAN [7].
+It is important to note that ACR [21] and MuSc [23] consider the batched zero-shot setting (see Table 1), thus, covering a different setting than AnomalyDINO.
+We adapt AnomalyDINO to this setting in Appendix D.
+Moreover, APRIL-GAN and AnomalyCLIP, require training on a related dataset, which is in contrast to our proposed training-free method.
+In Tables 2 and 3, results reported by [18] are indicated by , the result of the WinCLIP re-implementation by [20] (where J.
+Jeong is also coauthor) by .
+All other results are taken from the original publication.
+For ADP [20] we report the (usually slightly better) variant ADP (with access to the class label).
+For GraphCore no standard deviations are reported [41].
+4.1 Few-Shot Anomaly Detection and Anomaly Segmentation
+Results
+We summarize the results for few-shot anomaly detection on MVTec-AD and VisA in Table 2 and Table 3, respectively.
 […]
-Setting Method Detection Segmentation
+Method Classification Segmentation
 AUROC F1-max AP AUROC F1-max PRO
-1-shot SPADE 82.9$\pm$2.6 91.1$\pm$1.0 91.7$\pm$1.2 92.0$\pm$0.3 44.5$\pm$1.0 85.7$\pm$0.7
+1-shot
+SPADE 82.9$\pm$2.6 91.1$\pm$1.0 91.7$\pm$1.2 92.0$\pm$0.3 44.5$\pm$1.0 85.7$\pm$0.7
 PatchCore 86.3$\pm$3.3 92.0$\pm$1.5 93.8$\pm$1.7 93.3$\pm$0.6 53.0$\pm$1.7 82.3$\pm$1.3
@@
 […]
-AnomalyDINO 96.6$\pm$0.5 96.0$\pm$0.2 98.2$\pm$0.3 96.3$\pm$0.1 58.2$\pm$0.8 91.9$\pm$0.1
-2-shot SPADE 81.0$\pm$2.0 90.3$\pm$0.8 90.6$\pm$0.8 91.2$\pm$0.4 42.4$\pm$1.0 83.9$\pm$0.7
+AnomalyDINO-S (448) 96.5$\pm$0.4 96.0$\pm$0.2 98.1$\pm$0.3 96.3$\pm$0.1 57.9$\pm$0.8 91.7$\pm$0.1
+AnomalyDINO-S (672) 96.6$\pm$0.4 95.8$\pm$0.5 98.2$\pm$0.2 96.8$\pm$0.1 60.2$\pm$1.1 92.7$\pm$0.1
+2-shot
+SPADE 81.0$\pm$2.0 90.3$\pm$0.8 90.6$\pm$0.8 91.2$\pm$0.4 42.4$\pm$1.0 83.9$\pm$0.7
 PatchCore 83.4$\pm$3.0 90.5$\pm$1.5 92.2$\pm$1.5 92.0$\pm$1.0 50.4$\pm$2.1 79.7$\pm$2.0
-GraphCore 91.9$\pm$/ / / 96.9$\pm$/ / /
 […]
@@
 […]
-4-shot SPADE 84.8$\pm$2.5 91.5$\pm$0.9 92.5$\pm$1.2 92.7$\pm$0.3 46.2$\pm$1.3 87.0$\pm$0.5
+AnomalyDINO-S (448) 96.7$\pm$0.8 96.5$\pm$0.4 98.1$\pm$0.7 96.5$\pm$0.2 58.5$\pm$0.5 92.0$\pm$0.2
+AnomalyDINO-S (672) 96.9$\pm$0.7 96.1$\pm$0.3 98.2$\pm$0.5 97.0$\pm$0.2 61.0$\pm$0.5 93.1$\pm$0.2
+4-shot
+SPADE 84.8$\pm$2.5 91.5$\pm$0.9 92.5$\pm$1.2 92.7$\pm$0.3 46.2$\pm$1.3 87.0$\pm$0.5
 PatchCore 88.8$\pm$2.6 92.6$\pm$1.6 94.5$\pm$1.5 94.3$\pm$0.5 55.0$\pm$1.9 84.3$\pm$1.6
@@
 […]
+AnomalyDINO-S (448) 97.6$\pm$0.1 97.0$\pm$0.3 98.4$\pm$0.3 96.7$\pm$0.1 59.2$\pm$0.4 92.4$\pm$0.1
 […]
-8-shot GraphCore 95.9$\pm$/ / / 97.8$\pm$/ / /
+8-shot
+GraphCore 95.9$\pm$/ / / 97.8$\pm$/ / /
 WinCLIP+ 94.6$\pm$0.1 / / / / /
@@
 […]
+AnomalyDINO-S (448) 98.0$\pm$0.1 97.4$\pm$0.1 99.0$\pm$0.2 97.0$\pm$0.1 59.6$\pm$0.2 92.7$\pm$0.0
 […]
-16-shot WinCLIP+ 94.8$\pm$0.1 / / / / /
+16-shot
+WinCLIP+ 94.8$\pm$0.1 / / / / /
 […]
 […]
-AnomalyDINO 98.4$\pm$0.1 97.8$\pm$0.2 99.3$\pm$0.0 97.1$\pm$0.1 60.2$\pm$0.1 93.0$\pm$0.1
-Baselines
-We compare AnomalyDINO with a range of modern zero- and few-shot AD models, e.g., SPADE (Cohen and Hoshen, 2020), PaDiM (Defard et al., 2020), PatchCode (Roth et al., 2022), WinCLIP+ (Jeong et al., 2023), and APRIL-GAN (Chen et al., 2023).
-It is important to note that ACR (Li et al., 2024a) and MuSc (Li et al., 2024b) consider the batched zero-shot setting (see Table 1), thus, covering a different setting than AnomalyDINO.
-We adapt AnomalyDINO to this setting in Appendix D.
-Moreover, APRIL-GAN and AnomalyCLIP, require training on a related dataset, which is in contrast to our proposed training-free method.
-In Tables 2 and 3, results reported by Jeong et al.
-(2023) are indicated by , the result of the WinCLIP re-implementation by Kwak et al.
-(2024) (where J.
-Jeong is also coauthor) by .
-All other results are taken from the original publication.
-For ADP (Kwak et al., 2024) we report the (usually slightly better) variant ADP (with access to the class label).
-For GraphCore no standard deviations are reported (Xie et al., 2023).
-4.1 Few-Shot Anomaly Detection and Anomaly Segmentation
-Results
-We summarize the results for few-shot anomaly detection on MVTec-AD and VisA in Table 2 and Table 3, respectively.
-Regarding MVTec-AD, our method achieves state-of-the-art $k$-shot detection performance across all $k\in\{1,2,4,8,16\}$ in every reported metric, outperforming even approaches that require training on an additional data set (e.g., ADP and APRIL-GAN).
+AnomalyDINO-S (448) 98.3$\pm$0.1 97.7$\pm$0.2 99.3$\pm$0.0 97.1$\pm$0.1 60.0$\pm$0.1 92.9$\pm$0.1
+AnomalyDINO-S (672) 98.4$\pm$0.1 97.6$\pm$0.1 99.3$\pm$0.0 97.5$\pm$0.0 62.7$\pm$0.1 94.0$\pm$0.1
+Regarding MVTec-AD, our method achieves state-of-the-art $k$-shot detection performance across all $k\in{1,2,4,8,16}$ for every reported metric, outperforming approaches that require additional training data sets (such as ADP and APRIL-GAN).
+The method also demonstrates superior anomaly localization, with results showing that while detection performance remains comparable across different resolutions (448 vs.
+672), a higher resolution enhances localization performance.
+Furthermore, we observe clear improvements as the number of samples increases.
 […]
-Setting Method Detection Segmentation
+Method Classification Segmentation
 AUROC F1-max AP AUROC F1-max PRO
-1-shot SPADE 79.5$\pm$4.0 80.7$\pm$1.9 82.0$\pm$3.3 95.6$\pm$0.4 35.5$\pm$2.2 84.1$\pm$1.6
+1-shot
+SPADE 79.5$\pm$4.0 80.7$\pm$1.9 82.0$\pm$3.3 95.6$\pm$0.4 35.5$\pm$2.2 84.1$\pm$1.6
 PaDiM 62.8$\pm$5.4 75.3$\pm$1.2 68.3$\pm$4.0 89.9$\pm$0.8 17.4$\pm$1.7 64.3$\pm$2.4
 PatchCore 79.9$\pm$2.9 81.7$\pm$1.6 82.8$\pm$2.3 95.4$\pm$0.6 38.0$\pm$1.9 80.5$\pm$2.5
-WinCLIP+ 83.8$\pm$4.0 83.1$\pm$1.7 85.1$\pm$4.0 96.4$\pm$0.4 41.3$\pm$2.3 85.1$\pm$2.1
-APRIL-GAN 91.2$\pm$0.8 86.9$\pm$0.6 93.3$\pm$0.8 96.0$\pm$0.0 38.5$\pm$0.3 90.0$\pm$0.1
 […]
-2-shot SPADE 80.7$\pm$5.0 81.7$\pm$2.5 82.3$\pm$4.3 96.2$\pm$0.4 40.5$\pm$3.7 85.7$\pm$1.1
+[…]
+AnomalyDINO-S (448) 85.6$\pm$1.5 83.1$\pm$1.1 86.6$\pm$1.3 97.5$\pm$0.1 41.9$\pm$0.5 90.7$\pm$0.5
+[…]
+2-shot
+SPADE 80.7$\pm$5.0 81.7$\pm$2.5 82.3$\pm$4.3 96.2$\pm$0.4 40.5$\pm$3.7 85.7$\pm$1.1
 PaDiM 67.4$\pm$5.1 75.7$\pm$1.8 71.6$\pm$3.8 92.0$\pm$0.7 21.1$\pm$2.4 70.1$\pm$2.6
@@
 ADP 86.9$\pm$0.9 / / / / /
-APRIL-GAN 92.2$\pm$0.3 87.7$\pm$0.3 94.2$\pm$0.3 96.2$\pm$0.0 39.3$\pm$0.2 90.1$\pm$0.1
 […]
-4-shot SPADE 81.7$\pm$3.4 82.1$\pm$2.1 83.4$\pm$2.7 96.6$\pm$0.3 43.6$\pm$3.6 87.3$\pm$0.8
+AnomalyDINO-S (448) 88.3$\pm$1.8 84.8$\pm$1.2 89.2$\pm$1.3 97.8$\pm$0.1 44.2$\pm$0.3 91.7$\pm$0.5
+[…]
+4-shot
+SPADE 81.7$\pm$3.4 82.1$\pm$2.1 83.4$\pm$2.7 96.6$\pm$0.3 43.6$\pm$3.6 87.3$\pm$0.8
 PaDiM 72.8$\pm$2.9 78.0$\pm$1.2 75.6$\pm$2.2 93.2$\pm$0.5 24.6$\pm$1.8 72.6$\pm$1.9
-[…]
+PatchCore 85.3$\pm$2.1 84.3$\pm$1.3 87.5$\pm$2.1 96.8$\pm$0.3 43.9$\pm$3.1 84.9$\pm$1.4
 […]
 ADP 88.4$\pm$0.4 / / / / /
-APRIL-GAN 92.6$\pm$0.4 88.4$\pm$0.5 94.5$\pm$0.3 96.2$\pm$0.0 40.0$\pm$0.1 90.2$\pm$0.1
 […]
-8-shot WinCLIP+ 85.0$\pm$0.0 / / / / /
+AnomalyDINO-S (448) 91.3$\pm$0.8 87.5$\pm$1.0 91.8$\pm$0.7 98.0$\pm$0.0 46.1$\pm$0.3 92.5$\pm$0.2
+[…]
+8-shot
+WinCLIP+ 85.0$\pm$0.0 / / / / /
 ADP 89.2$\pm$0.1 / / / / /
 […]
+AnomalyDINO-S (448) 92.6$\pm$0.1 88.6$\pm$0.2 92.9$\pm$0.2 98.2$\pm$0.0 47.6$\pm$0.5 93.3$\pm$0.2
 […]
-16-shot WinCLIP+ 85.0$\pm$0.1 / / / / /
+16-shot
+WinCLIP+ 85.0$\pm$0.1 / / / / /
 ADP 90.1$\pm$0.5 / / / / /
 […]
-[…]
-AnomalyDINO reduces the gap between few-shot and full-shot methods in anomaly detection significantly, and it demonstrates the strongest performance in anomaly localization.
-Furthermore, we see a clear improvement in the method’s results as the quantity of samples increases.
-In terms of anomaly detection in the VisA benchmark (Table 3), APRIL-GAN demonstrates the strongest performance for $k\in\{1,2,4\}$.
-Nonetheless, AnomalyDINO consistently achieves second-best results, comparable results with APRIL-GAN in the 8-shot setting, and sets new state-of-the-art for $k=16$.
+AnomalyDINO-S (448) 93.8$\pm$0.1 89.9$\pm$0.3 94.2$\pm$0.3 98.3$\pm$0.0 48.6$\pm$0.3 93.8$\pm$0.2
+AnomalyDINO-S (672) 94.8$\pm$0.2 90.9$\pm$0.2 95.3$\pm$0.3 98.5$\pm$0.0 52.5$\pm$0.5 95.3$\pm$0.2
+In terms of anomaly detection in the VisA benchmark (Table 3), APRIL-GAN demonstrates the strongest performance for $k\in\{1,2\}$.
+Nonetheless, AnomalyDINO consistently achieves second-best results for $k\in\{1,2\}$, comparable results in the 4-shot setting, and sets new state-of-the-art for $k\in\{8,16\}$.
 This can be attributed to AnomalyDINO’s ability to benefit more from a richer memory bank $\mathcal{M}$ than APRIL-GAN.
-We hypothesize that meta-learning exerts a greater influence on APRIL-GAN (i.e., training on MVTec-AD, when testing on VisA, and vice-versa) compared to learning from the nominal features of the given reference samples.Note also that, AnomalyDINO outperforms all other training-free approaches.
-Table 4: Detection results for other settings. All results are AUROC values in (%).
+We hypothesize that meta-learning exerts a greater influence on APRIL-GAN (i.e., training on MVTec-AD, when testing on VisA, and vice-versa) compared to learning from the nominal features of the given reference samples.
+Note also that AnomalyDINO outperforms all other training-free approaches.
+Comparing the segmentation performance on VisA, Table 3 reveals a clear picture: AnomalyDINO consistently shows the strongest localization performance in all metrics considered.
+While AnomalyDINO-S (448) already demonstrates strong performance, the advantages of using a higher resolution (672) become more evident on the VisA dataset.
+We attribute this fact to smaller anomalous regions (for which smaller effective patch sizes are beneficial) and more complex scenes (compared to MVTec-AD).
+We have further adapted AnomalyDINO to the batched zero-shot setting (see Appendix D).
+This adaptation is straightforward and relatively simple, especially compared to MuSc.
+Notably, we did not employ additional techniques such as ‘Re-scoring with Constrained Image-level Neighborhood’ or ‘Local Neighborhood Aggregation with Multiple Degrees’ [23].
+The results obtained, as detailed in Table 4, are therefore quite satisfactory.
+We conclude that AnomalyDINO—despite its simplicity—rivals other methods in all settings and even comes out ahead in most of the settings (e.g., significantly reducing the gap between few-shot and full-shot methods for MVTec-AD).
+Within the class of training-free models, it is the clear winner essentially across the board.
+These results do not only demonstrate the virtues of AnomalyDINO itself but, more generally, highlight the merits of strong visual features as compared to highly engineered architectures.
+We provide further qualitative results in Appendix A, and discuss the limitations (such as detecting semantic anomalies) and specific failure cases of our approach in Appendix B.
+Table 4: Detection results for other settings. All results are AUROC values (in %).
 Setting Method MVTec-AD VisA
@@
 MuSc 97.8 92.8
-AnomalyDINO 93.3 89.6
-Comparing the segmentation performance on VisA, Table 3 reveals a mixed picture: WinCLIP+ shows strong localization performance in terms of pixel-level AUROC, and APRIL-GAN achieves the highest scores in PRO.
-AnomalyDINO consistently achieves results in per-region-overlap that are second-best compared to all other, and the best among the training-free methods.
-In addition, AnomalyDINO shows the best segmentation performance when measured in F1-max, a metric that accounts for the strong imbalance of nominal and anomalous pixels.
-We have further adapted AnomalyDINO to the batched zero-shot setting (see Appendix D).
-This adaptation is straightforward and relatively simple, especially compared to MuSc.
-Notably, we did not employ additional techniques such as ‘Re-scoring with Constrained Image-level Neighborhood’ or ‘Local Neighborhood Aggregation with Multiple Degrees’ (Li et al., 2024b).
-The results obtained, as detailed in Table 4, are therefore quite satisfactory.
-We conclude that AnomalyDINO—despite its simplicity—rivals other methods in all settings and even comes out ahead in most of the settings.
-Within the class of training-free models, it is the clear winner essentially across the board.
-These results do not only demonstrate the virtues of AnomalyDINO itself but, more generally, highlight the merits of strong visual features as compared to highly engineered architectures.
-We provide further qualitative results in Figures 3 and 4, and discuss the limitations (such as detecting semantic anomalies) and specific failure cases of our approach in Appendix B.
+AnomalyDINO-S (448) 93.0 89.7
+AnomalyDINO-S (672) 94.2 90.7
 4.2 Ablation Study
-We conduct additional experiments to assess the effect of specific design choices in our pipeline. The full results are provided in Appendix C. Here, we briefly describe the main insights.
+We conduct additional experiments to assess the effect of specific design choices in our pipeline.
+The full ablation study is provided in Appendix C, here we briefly summarize the main insights.
+Figure 3 supports the ablation study by highlighting the two key aspects: performance and runtime.
+Figure 3: Detection AUROC vs.
+inference time per sample on MVTec-AD in the 1-shot setting.
+The input resolution is given in parentheses after the method name.
+All runtimes are measured on a single NVIDIA A40 if not stated otherwise.
+(Note that for ADP and WinCLIP+ no official code is available.)
+Inference time
+Comparing the inference times of AnomalyDINO with the other approaches in the 1-shot setting (see Figure 3), we observe that AnomalyDINO achieves significantly faster inference times than SOTA few-shot competitors (note the logarithmic scale).
+For example, AnomalyDINO-S takes roughly 60ms to process an image at a resolution of 448.
+The only approaches with lower inference time are PatchCore and AnomalyDINO with ViT-backbones trained on ImageNet, which however sacrifice performance.
+Notably, while augmentations increase the memory bank, and thus, the potential runtime of the nearest neighborhood search, we find that the effect is negligible in practice.
+[^5] Our implementation leverages GPU-accelerated neighborhood search [19], leading to only slightly increased inference time.
+With increasing size of $\mathcal{M}$ reduction techniques such as coreset subsampling [33] are advisable.
+A more detailed runtime analysis is included in Appendix C, Table 9.
 Preprocessing
-Comparing the two different preprocessing settings, ‘agnostic’ (our default) and ‘informed’ we observe that the agnostic preprocessing slightly outperforms the informed counterpart, see Figures 8 and 9.
-Depending on the product category suitable preprocessing steps lead to substantial improvements.
-In some cases augmenting the reference sample with rotations also negatively impacts the detection performance.
+To study the impact of preprocessing, we compare the resulting detection AUROCs of AnomalyDINO with and without preprocessing in Figure 3.
+Across four different configurations of AnomalyDINO, we observe that by incorporating the proposed preprocessing, the resulting AUROCs increase by approximately $2\%$ without significantly affecting the inference time.
+The effect of proper preprocessing seems to increase with higher resolutions.
+Additionally, we compare the two different preprocessing settings, ‘agnostic’ (our default) and ‘informed’, and find that the agnostic preprocessing slightly outperforms the informed counterpart, see Figures 9 and 10.
+Depending on the product category, suitable preprocessing steps lead to substantial improvements.
+And while in principle augmenting the reference sample with rotations may negatively impact the detection performance, we observe this in only very few examples.
 For the full discussion, see Section C.1.
-Inference time
-Data augmentation increases the computational complexity of the proposed pipeline, as well as the memory bank size, and thus, the runtime of our nearest neighborhood approach.
-[^5] In practice, our implementation leverages GPU-accelerated neighborhood search (Johnson et al., 2019), leading to only slightly increased inference time.
-With increasing size of $\mathcal{M}$ reduction techniques such as coreset subsampling (Roth et al., 2022) are advisable.
-The effects of masking and augmentations on the inference time, presented in Figure 11, are not negligible but can be reduced by approximately 30% in the ‘informed’ scenario (on MVTec-AD).
-Aggregation statistic
+Aggregation choice
 We compare the default scoring method, the empirical tail value at risk, to other suitable aggregations in Section C.2.
 The mean of the 1% highest distances from test patches to $\mathcal{M}$ improves over the standard choice (maximum of the upsampled and smoothed patch distances).
-Architecture size
+Architecture size/choice
+We also evaluated the proposed framework using Vision Transformers pre-trained on ImageNet (see Figure 3).
+While these backbones slightly outperform PatchCore, they give weaker features compared to DINOv2 and are incompatible with the proposed masking procedure (zero-shot masking does not function effectively).
+This underlines that DINOv2 is indeed excellently suited for visual AD.
+In addition, ViT-based architectures like DINOv2 easily allow handling images with varying resolutions.
+In this context, we find that operating at a resolution of 448 gives the best trade-off between performance and inference time, but even higher detection AUROCs are possible at higher resolutions.
 In addition, we evaluated our pipeline with DINOv2 at different scales (ViT-S, ViT-B, ViT-L).
 The full results are given in Section C.3, Figure 14.
-We observe no considerable differences, in particular, also larger backbones give state-of-the-art results on MVTec-AD (all $k$) and VisA ($k=16$).
+We observe no considerable differences, in particular, also larger backbones give state-of-the-art results on MVTec-AD (all $k$) and VisA ($k\geq 8$).
 Interestingly, larger architectures do not necessarily translate to higher performance.
-As the dimensionality of the feature space $\mathcal{F}$ increases with larger architectures, the distances between nominal patches also increases.
 The smallest model demonstrates the best performance on MVTec-AD, which primarily consists of single objects and less complex scenes.
-In contrast, the larger architecture size L performs better on VisA, which often involves multiple and more complex objects.
+In contrast, the larger architecture sizes perform better on VisA, which often involves multiple and more complex objects.
 This suggests the importance of achieving the ‘just right’ level of abstraction for the specific task, particularly in the few-shot regime.
@@
 This paper proposes a vision-only approach for one- and few-shot anomaly detection.
-Our method is based on similarities between patch representations extracted by DINOv2 (Oquab et al., 2024).
-We carefully design means to populate the nominal memory bank with diverse and relevant features utilizing zero-shot segmentation and simple data augmentation.
+Our method is based on similarities between patch representations extracted by DINOv2 [29].
+We carefully design means to populate the nominal memory bank with diverse and relevant features while minimizing false positives utilizing zero-shot segmentation and simple data augmentation.
 Industrial settings require fast deployment, easy debugging and error correction, and rapid adaptation for covariance shifts in the normal data distribution.
@@
 The proposed method, AnomalyDINO, achieves state-of-the-art results on MVTec-AD and rivals all competing methods on VisA while outperforming all other training-free approaches.
+Thus, our approach achieves the best of both worlds: improved performance and reduced inference time, especially compared to more complex vision-language methods.
 Its simplicity and strong performance render AnomalyDINO an excellent candidate for practitioners for industrial anomaly detection and an effective baseline for assessing few-shot and even full-shot anomaly detection in ongoing research.
@@
 However, the individual parts of our pipeline can easily be exchanged for more sophisticated alternatives.
-For example, our simple masking approach could be replaced by more specialized and adaptive masking techniques (which may also be relevant to other methods building on features extracted by DINOv2), and the simple upsampling and smoothing approach could be substituted by more sophisticated methods (e.g., Fu et al., 2024) to obtain the anomaly map from patch features.
+For example, our simple masking approach could be replaced by more specialized and adaptive masking techniques (which may also be relevant to other methods building on features extracted by DINOv2), and the simple upsampling and smoothing approach could be substituted by more sophisticated methods like [14] to obtain the anomaly map from patch features.
 It would be interesting to see if that leads to further improvements in anomaly detection and localization, thereby further reducing the gap between few- and full-shot anomaly detectors.
@@
 We believe that simpler methods can be adapted more quickly, monitored more effectively, and are therefore more reliable.
-In this context, awareness of the risks of overreliance must be heightened (see Appendix B for identified failure cases for the proposed method).
+In this context, awareness of the risks of overreliance must be heightened (see Appendix B for identified failure cases of the proposed method).
 In addition, strong visual anomaly detectors could also lead to potentially malicious or unintended uses.
@@
 Appendix A Detailed experimental results
-In this section, we present the results of the default setting (AnomalyDINO-S and ‘agnostic’ preprocessing) in detail.
-The full results per category for MVTec-AD and VisA are given in Tables 5 and 6.
-Additional visualizations of anomaly maps are presented in Figures 3 and 4 for MVTec-AD and VisA, respectively.
+In this section, we present additional visualization and results for the proposed method.
 query GT 1-shot 8-shot query GT 1-shot 8-shot
 Figure 4: Examples – MVTec-AD.
-Depicted are, from left to right, a test sample per category (query), the ground truth anomaly annotation (GT), and the predicted anomaly map from AnomalyDINO in the 1- and 8-shot settings.
-The color coding is normalized by the maximum score on the ‘good’ test samples (per category).
+Depicted are, from left to right, a test sample per category (query), the ground truth anomaly annotation (GT), and the predicted anomaly map from AnomalyDINO-S (448) in the 1- and 8-shot settings.
+The color coding is normalized by the max.
+score over ‘good’ test samples.
+Further anomaly maps, predicted by AnomalyDINO, are presented in Figures 4 and 5. The full results per category for MVTec-AD and VisA are given in Tables 5 and 6, respectively.
 Table 5: Detailed results on MVTec-AD.
-Reported are results on anomaly detection (image-level AUROC) and segmentation (PRO) of AnomalyDINO with default preprocessing (mean and standard deviation over three independent runs for each metric, all results in %).
+Reported are results on anomaly detection (image-level AUROC) and segmentation (PRO) of AnomalyDINO-S (672) with default preprocessing (mean and standard deviation over three independent runs, all results in %).
 Shots 1-shot 2-shot 4-shot 8-shot 16-shot
 AUROC PRO AUROC PRO AUROC PRO AUROC PRO AUROC PRO
-Bottle 99.5$\pm$0.4 96.1$\pm$0.3 99.7$\pm$0.1 96.4$\pm$0.1 99.9$\pm$0.1 96.6$\pm$0.2 100.0$\pm$0.0 96.6$\pm$0.1 100.0$\pm$0.0 96.5$\pm$0.2
-Cable 91.6$\pm$1.2 88.2$\pm$0.7 92.1$\pm$0.7 88.5$\pm$0.7 93.8$\pm$1.4 89.5$\pm$0.5 94.7$\pm$0.3 89.6$\pm$0.3 94.4$\pm$1.1 89.5$\pm$0.7
-Capsule 89.8$\pm$5.9 96.4$\pm$0.4 88.9$\pm$7.1 96.7$\pm$0.5 95.4$\pm$0.9 97.3$\pm$0.1 94.7$\pm$1.3 97.4$\pm$0.1 95.5$\pm$0.3 97.5$\pm$0.1
-Carpet 100.0$\pm$0.0 97.9$\pm$0.1 100.0$\pm$0.0 97.9$\pm$0.0 100.0$\pm$0.0 97.7$\pm$0.1 100.0$\pm$0.0 97.7$\pm$0.1 100.0$\pm$0.0 97.6$\pm$0.2
-Grid 99.2$\pm$0.1 96.1$\pm$0.0 99.2$\pm$0.1 96.3$\pm$0.1 99.2$\pm$0.1 96.2$\pm$0.2 99.4$\pm$0.2 96.2$\pm$0.1 99.8$\pm$0.3 95.9$\pm$0.3
-Hazelnut 98.1$\pm$1.1 97.5$\pm$0.4 99.7$\pm$0.4 97.7$\pm$0.4 99.8$\pm$0.1 97.8$\pm$0.1 100.0$\pm$0.0 97.8$\pm$0.2 100.0$\pm$0.0 97.7$\pm$0.1
-Leather 100.0$\pm$0.0 97.9$\pm$0.2 100.0$\pm$0.0 97.5$\pm$0.3 100.0$\pm$0.0 97.1$\pm$0.3 100.0$\pm$0.0 97.0$\pm$0.1 100.0$\pm$0.0 96.7$\pm$0.1
-Metal nut 99.9$\pm$0.2 93.8$\pm$0.1 100.0$\pm$0.1 94.4$\pm$0.3 100.0$\pm$0.0 95.1$\pm$0.1 100.0$\pm$0.0 95.2$\pm$0.3 100.0$\pm$0.0 95.5$\pm$0.1
-Pill 93.1$\pm$1.5 96.9$\pm$0.2 94.9$\pm$2.2 97.1$\pm$0.2 96.3$\pm$0.1 97.3$\pm$0.0 97.1$\pm$0.4 97.4$\pm$0.1 97.6$\pm$0.3 97.5$\pm$0.1
-Screw 90.1$\pm$1.0 92.5$\pm$0.2 92.3$\pm$0.2 93.5$\pm$0.0 91.3$\pm$2.4 93.6$\pm$0.7 92.5$\pm$0.8 94.6$\pm$0.4 93.5$\pm$1.4 95.3$\pm$0.2
-Tile 100.0$\pm$0.0 88.6$\pm$0.2 100.0$\pm$0.0 88.2$\pm$0.2 100.0$\pm$0.0 87.9$\pm$0.2 100.0$\pm$0.0 87.5$\pm$0.2 100.0$\pm$0.0 86.9$\pm$0.6
-Toothbrush 99.4$\pm$0.5 94.5$\pm$0.1 99.9$\pm$0.1 94.9$\pm$0.3 99.9$\pm$0.1 95.0$\pm$0.3 99.9$\pm$0.1 95.5$\pm$0.5 99.6$\pm$0.3 95.9$\pm$0.6
-Transistor 90.9$\pm$2.1 64.4$\pm$0.5 89.2$\pm$5.9 64.8$\pm$3.1 91.9$\pm$3.5 66.5$\pm$1.4 96.4$\pm$1.8 70.3$\pm$1.6 98.3$\pm$0.5 72.7$\pm$1.0
-Wood 98.8$\pm$0.4 93.9$\pm$0.3 98.7$\pm$0.6 93.8$\pm$0.2 98.6$\pm$0.6 93.8$\pm$0.1 98.6$\pm$0.2 93.8$\pm$0.2 98.2$\pm$0.4 93.5$\pm$0.5
-Zipper 98.9$\pm$0.1 83.2$\pm$0.4 99.0$\pm$0.1 84.4$\pm$0.7 99.3$\pm$0.2 85.6$\pm$0.2 99.2$\pm$0.2 85.5$\pm$0.6 99.3$\pm$0.3 85.9$\pm$0.7
-Mean 96.6$\pm$0.5 91.9$\pm$0.1 96.9$\pm$0.8 92.1$\pm$0.2 97.7$\pm$0.1 92.5$\pm$0.1 98.2$\pm$0.1 92.8$\pm$0.0 98.4$\pm$0.1 93.0$\pm$0.1
+Bottle 99.7$\pm$0.2 95.9$\pm$0.3 99.9$\pm$0.1 96.3$\pm$0.0 99.9$\pm$0.1 96.7$\pm$0.1 100.0$\pm$0.0 96.5$\pm$0.4 99.9$\pm$0.1 96.4$\pm$0.4
+Cable 92.7$\pm$0.8 89.4$\pm$0.4 92.4$\pm$1.1 89.5$\pm$0.4 93.8$\pm$0.9 90.4$\pm$0.3 95.2$\pm$0.3 90.5$\pm$0.2 95.1$\pm$0.6 90.5$\pm$0.6
+Capsule 90.2$\pm$5.5 97.1$\pm$0.3 89.2$\pm$7.9 97.3$\pm$0.6 95.8$\pm$0.5 97.9$\pm$0.1 95.6$\pm$0.5 97.9$\pm$0.1 95.5$\pm$0.6 98.1$\pm$0.1
+Carpet 100.0$\pm$0.0 97.8$\pm$0.0 100.0$\pm$0.0 97.9$\pm$0.0 100.0$\pm$0.0 97.8$\pm$0.0 100.0$\pm$0.0 97.8$\pm$0.0 100.0$\pm$0.0 97.8$\pm$0.0
+Grid 99.1$\pm$0.2 97.2$\pm$0.1 99.2$\pm$0.4 97.2$\pm$0.1 99.5$\pm$0.3 97.2$\pm$0.1 99.5$\pm$0.1 97.2$\pm$0.0 99.7$\pm$0.3 97.0$\pm$0.2
+Hazelnut 97.5$\pm$2.6 97.4$\pm$0.4 99.6$\pm$0.5 98.0$\pm$0.3 99.8$\pm$0.1 98.0$\pm$0.1 100.0$\pm$0.0 98.1$\pm$0.1 100.0$\pm$0.0 98.1$\pm$0.1
+Leather 100.0$\pm$0.0 97.9$\pm$0.1 100.0$\pm$0.0 97.8$\pm$0.0 100.0$\pm$0.0 97.6$\pm$0.1 100.0$\pm$0.0 97.6$\pm$0.1 100.0$\pm$0.0 97.3$\pm$0.2
+Metal nut 99.9$\pm$0.1 94.2$\pm$0.0 100.0$\pm$0.0 94.6$\pm$0.2 100.0$\pm$0.0 95.3$\pm$0.1 100.0$\pm$0.0 95.4$\pm$0.3 100.0$\pm$0.0 95.7$\pm$0.1
+Pill 93.7$\pm$0.9 97.3$\pm$0.1 95.4$\pm$0.7 97.5$\pm$0.1 96.0$\pm$0.2 97.6$\pm$0.1 97.2$\pm$0.2 97.7$\pm$0.1 97.9$\pm$0.1 97.8$\pm$0.1
+Screw 93.2$\pm$0.3 93.4$\pm$0.4 93.5$\pm$0.8 94.3$\pm$0.4 92.7$\pm$2.3 94.5$\pm$0.9 93.5$\pm$1.1 95.2$\pm$0.5 94.7$\pm$0.9 95.9$\pm$0.3
+Tile 100.0$\pm$0.0 88.0$\pm$0.2 100.0$\pm$0.0 87.6$\pm$0.4 100.0$\pm$0.0 87.1$\pm$0.4 100.0$\pm$0.0 86.7$\pm$0.2 100.0$\pm$0.0 86.0$\pm$0.5
+Toothbrush 97.4$\pm$0.5 94.0$\pm$0.8 98.1$\pm$1.0 94.7$\pm$0.3 97.5$\pm$0.6 94.7$\pm$0.3 97.7$\pm$1.6 95.1$\pm$0.8 98.1$\pm$1.8 95.8$\pm$1.0
+Transistor 90.9$\pm$1.2 67.3$\pm$2.1 89.4$\pm$4.6 68.4$\pm$3.1 93.2$\pm$2.2 70.6$\pm$1.4 96.2$\pm$1.3 75.3$\pm$1.9 97.6$\pm$0.3 78.2$\pm$1.4
+Wood 98.0$\pm$0.2 94.7$\pm$0.1 98.0$\pm$0.1 94.6$\pm$0.0 97.9$\pm$0.2 94.6$\pm$0.1 98.3$\pm$0.4 94.4$\pm$0.3 98.3$\pm$0.6 94.2$\pm$0.4
+Zipper 97.4$\pm$0.9 89.2$\pm$1.2 98.9$\pm$0.4 90.2$\pm$0.4 99.0$\pm$0.4 91.2$\pm$0.3 99.6$\pm$0.1 91.1$\pm$0.4 99.6$\pm$0.3 91.7$\pm$0.5
+Mean 96.6$\pm$0.4 92.7$\pm$0.1 96.9$\pm$0.7 93.1$\pm$0.2 97.7$\pm$0.2 93.4$\pm$0.1 98.2$\pm$0.2 93.8$\pm$0.1 98.4$\pm$0.1 94.0$\pm$0.1
+Table 6: Detailed results on VisA.
+Reported are results on anomaly detection (image-level AUROC) and segmentation (PRO) of AnomalyDINO-S (672) with default preprocessing (mean and standard deviation over three independent runs, all results in %).
+Shots 1-shot 2-shot 4-shot 8-shot 16-shot
+AUROC PRO AUROC PRO AUROC PRO AUROC PRO AUROC PRO
+Candle 87.9$\pm$0.3 96.8$\pm$0.4 89.4$\pm$3.0 97.0$\pm$0.2 91.3$\pm$2.9 97.2$\pm$0.1 93.5$\pm$1.2 97.3$\pm$0.2 94.5$\pm$0.5 97.6$\pm$0.2
+Capsules 98.4$\pm$0.5 95.1$\pm$0.7 98.9$\pm$0.1 95.5$\pm$0.2 99.2$\pm$0.1 96.3$\pm$0.4 99.2$\pm$0.1 96.7$\pm$0.2 99.2$\pm$0.2 97.2$\pm$0.3
+Cashew 86.1$\pm$3.6 96.1$\pm$0.9 89.4$\pm$3.8 96.7$\pm$0.7 94.5$\pm$0.7 97.4$\pm$0.5 95.3$\pm$0.6 97.3$\pm$0.2 96.0$\pm$0.3 97.3$\pm$0.1
+Chewinggum 98.0$\pm$0.4 92.0$\pm$1.0 98.6$\pm$0.4 92.9$\pm$0.3 98.8$\pm$0.2 93.0$\pm$0.1 98.8$\pm$0.2 93.1$\pm$0.3 98.8$\pm$0.2 93.0$\pm$0.3
+Fryum 94.8$\pm$0.5 93.2$\pm$0.2 96.5$\pm$0.2 93.9$\pm$0.3 97.0$\pm$0.1 94.5$\pm$0.4 97.6$\pm$0.4 94.9$\pm$0.3 97.9$\pm$0.2 95.1$\pm$0.0
+Macaroni1 87.5$\pm$1.1 97.5$\pm$0.3 87.5$\pm$0.9 97.9$\pm$0.3 89.5$\pm$1.4 98.3$\pm$0.2 90.1$\pm$1.7 98.6$\pm$0.2 90.4$\pm$1.1 98.7$\pm$0.1
+Macaroni2 62.2$\pm$4.3 92.0$\pm$0.7 66.9$\pm$1.9 93.0$\pm$0.4 70.0$\pm$1.7 93.9$\pm$0.8 74.9$\pm$0.4 95.0$\pm$0.6 77.6$\pm$0.8 95.7$\pm$0.3
+PCB1 91.5$\pm$2.0 92.6$\pm$0.2 91.2$\pm$2.7 92.5$\pm$0.5 94.0$\pm$2.1 93.3$\pm$0.5 95.5$\pm$0.5 93.9$\pm$0.2 96.8$\pm$0.7 94.2$\pm$0.2
+PCB2 84.8$\pm$1.2 89.9$\pm$0.2 88.1$\pm$2.5 90.7$\pm$0.3 91.1$\pm$1.7 91.4$\pm$0.2 92.6$\pm$0.3 92.0$\pm$0.1 93.2$\pm$0.1 92.5$\pm$0.2
+PCB3 84.9$\pm$3.3 88.5$\pm$1.3 89.4$\pm$3.8 90.8$\pm$0.5 94.3$\pm$0.4 91.7$\pm$0.4 95.6$\pm$0.2 93.1$\pm$0.3 96.5$\pm$0.3 93.9$\pm$0.3
+PCB4 79.9$\pm$13.7 78.5$\pm$6.8 87.4$\pm$11.3 82.0$\pm$6.1 96.2$\pm$2.6 84.1$\pm$1.5 98.0$\pm$0.3 87.9$\pm$2.3 99.0$\pm$0.4 90.4$\pm$1.8
+Pipe fryum 92.7$\pm$2.7 98.0$\pm$0.0 93.3$\pm$1.2 97.9$\pm$0.2 94.6$\pm$1.9 97.8$\pm$0.1 95.0$\pm$1.6 97.6$\pm$0.1 97.2$\pm$0.9 97.7$\pm$0.1
+Mean 87.4$\pm$1.2 92.5$\pm$0.5 89.7$\pm$1.3 93.4$\pm$0.6 92.6$\pm$0.9 94.1$\pm$0.1 93.8$\pm$0.3 94.8$\pm$0.2 94.8$\pm$0.2 95.3$\pm$0.2
 query GT 1-shot 8-shot query GT 1-shot 8-shot
-Figure 4: Examples – VisA.
-Depicted are, from left to right, a test sample per category (query), the ground truth anomaly annotation (GT), and the predicted anomaly map from AnomalyDINO in the 1- and 8-shot settings.
+Figure 5: Examples – VisA (1/2).
+Depicted are, from left to right, a test sample per category (query), the ground truth anomaly annotation (GT), and the predicted anomaly map from AnomalyDINO-S (448) in the 1- and 8-shot settings.
 The color bar is normalized by the maximum score on the ‘good’ test samples (per category).
@@
 Best viewed at a higher zoom level as some anomalies are quite small.
-Table 6: Detailed results on VisA.
-Reported are results on anomaly detection (image-level AUROC) and segmentation (PRO) of AnomalyDINO with default preprocessing (mean and standard deviation over three independent runs for each metric, all results in %).
-Shots 1-shot 2-shot 4-shot 8-shot 16-shot
-AUROC PRO AUROC PRO AUROC PRO AUROC PRO AUROC PRO
-Candle 83.8$\pm$0.7 77.6$\pm$0.6 86.2$\pm$3.8 78.4$\pm$0.1 88.0$\pm$3.0 79.3$\pm$0.3 89.9$\pm$1.0 79.7$\pm$0.2 92.0$\pm$0.8 80.7$\pm$0.3
-Capsules 98.4$\pm$0.3 80.4$\pm$0.5 99.1$\pm$0.2 81.0$\pm$0.4 99.1$\pm$0.2 81.5$\pm$0.2 99.1$\pm$0.5 81.9$\pm$0.1 98.8$\pm$0.4 82.4$\pm$0.2
-Cashew 85.3$\pm$5.6 95.0$\pm$1.5 89.3$\pm$5.6 95.7$\pm$1.1 93.8$\pm$0.5 96.2$\pm$0.8 95.1$\pm$0.7 95.9$\pm$0.3 95.5$\pm$0.1 96.2$\pm$0.3
-Chewinggum 98.2$\pm$0.3 90.6$\pm$0.8 98.7$\pm$0.3 91.8$\pm$0.8 98.8$\pm$0.2 92.2$\pm$0.5 98.9$\pm$0.1 92.5$\pm$0.3 99.0$\pm$0.0 92.6$\pm$0.0
-Fryum 95.1$\pm$0.3 92.6$\pm$0.3 96.6$\pm$0.5 93.2$\pm$0.1 97.2$\pm$0.5 93.6$\pm$0.1 97.9$\pm$0.6 93.8$\pm$0.1 97.8$\pm$0.6 94.0$\pm$0.1
-Macaroni1 83.4$\pm$3.0 86.7$\pm$1.0 83.3$\pm$1.7 87.6$\pm$0.2 85.8$\pm$0.9 88.7$\pm$0.2 86.9$\pm$0.8 89.6$\pm$0.2 88.2$\pm$0.7 90.0$\pm$0.2
-Macaroni2 61.4$\pm$3.8 66.3$\pm$0.7 67.1$\pm$5.3 67.5$\pm$0.3 70.0$\pm$1.8 68.5$\pm$0.4 74.0$\pm$1.2 70.3$\pm$0.6 76.5$\pm$1.5 71.9$\pm$0.5
-PCB1 89.0$\pm$1.2 90.6$\pm$0.3 88.9$\pm$1.5 90.6$\pm$0.6 92.3$\pm$2.3 91.2$\pm$0.4 92.9$\pm$0.6 91.8$\pm$0.2 95.2$\pm$0.8 92.5$\pm$0.3
-PCB2 87.8$\pm$2.0 86.5$\pm$1.1 89.4$\pm$2.5 87.5$\pm$0.2 91.8$\pm$0.7 88.4$\pm$0.1 92.8$\pm$0.6 89.1$\pm$0.2 93.0$\pm$0.4 89.5$\pm$0.3
-PCB3 84.1$\pm$2.4 85.6$\pm$0.8 88.5$\pm$1.3 87.1$\pm$0.6 91.8$\pm$0.3 87.8$\pm$0.4 93.8$\pm$0.7 89.4$\pm$0.6 94.5$\pm$0.2 90.0$\pm$0.2
-PCB4 76.7$\pm$11.5 79.8$\pm$5.0 86.5$\pm$7.2 83.3$\pm$3.7 94.9$\pm$3.0 86.1$\pm$0.8 97.3$\pm$0.6 88.8$\pm$1.1 99.0$\pm$0.3 90.6$\pm$0.9
-Pipe fryum 91.8$\pm$2.6 97.4$\pm$0.3 93.5$\pm$3.0 97.5$\pm$0.3 94.8$\pm$2.7 97.2$\pm$0.1 96.5$\pm$1.6 97.3$\pm$0.1 98.0$\pm$0.3 97.3$\pm$0.1
-Mean 86.2$\pm$1.4 85.8$\pm$0.6 88.9$\pm$1.9 86.8$\pm$0.4 91.5$\pm$0.7 87.6$\pm$0.2 92.9$\pm$0.1 88.3$\pm$0.2 94.0$\pm$0.0 89.0$\pm$0.2
 The results presented in Tables 5 and 6 show that a) anomalies in some categories are more difficult to than others, b) detection and segmentation performance correlate strongly (measured in image-level AUROC and PRO), and c), that the performance of AnomalyDINO increases across the board with more available reference samples.
@@
 In addition, we see that also the variance in the reported metrics decreases with the number of nominal samples.
-In this context, we observe a peculiarity of the few-shot regime (which does not occur for sufficiently populated $\mathcal{M}$).
-The results crucially depend on the chosen reference image(s).
+In this context, we observe a peculiarity of the few-shot regime (which does not occur for sufficiently populated $\mathcal{M}$). The results crucially depend on the chosen reference image(s).
+Table 7: Full-shot results on MVTec-AD and VisA with AnomalyDINO-S in the default setting (no std reported as results are deterministic when all samples are considered).
+Dataset Resolution Detection Segmentation
+AUROC F1-max AP AUROC F1-max PRO
+MVTec-AD 448 99.3 98.8 99.7 97.9 61.8 93.9
+672 99.5 99.0 99.8 98.2 64.3 95.0
+VisA 448 97.2 93.7 97.6 98.7 50.5 95.0
+672 97.6 94.5 98.0 98.8 53.8 96.1
 The high variances, predominantly in the 1- and 2-shot setting, for category ‘Capsule’ for MVTec-AD, or ‘Cashew’ or ‘PCB4’ for VisA demonstrate this.
 We discuss the potential failure cases of choosing a sub-optimal reference sample in the following section (Section B.2).
+Finally, we also report the full-shot performance, see Table 7.
+As the diversity within $\mathcal{M}$ is sufficiently high given the large number of reference samples, we only apply masing here (no augmentations).
+The results demonstrate that the performance nicely improves for all considered metrics.
 Appendix B Limitations and failure cases
@@
 In contrast, AnomalyDINO is developed with low-level sensory anomalies in mind.
+Consider, for instance, the anomalies shown in Figure 6.
 (a) Reference sample
@@
 Depicted are a nominal reference sample, an example of a sensory anomaly (Figure 6(b)) and a semantic anomaly (Figure 6(c)) with anomaly maps predicted by AnomalyDINO (1-shot).
-Consider, for instance, the anomalies shown in Figure 5.
 While the cable with anomaly type ‘Bent Wire’ (Figure 6(b)) contains patches that cannot be well matched to patches of the reference image, this does not apply to the anomaly ‘Cable Swap’ (Figure 6(c)).
@@
 B.2 The importance of informative reference samples
+(a) Reference sample (left) and estimated anomaly map of ‘good’ test sample (right).
+(b) Reference sample (left) and estimated anomaly map of ‘good’ test sample (right).
+Figure 7: Example of an uninformative reference sample in category ‘Capsule’ (MVTec-AD).
+Some reference samples do not resemble the full concept of normality (here, the sample in Figure 7(b) does not show the text on the capsule, i.e., any nominal sample with text be considered anomalous).
+Anomaly maps predicted by AnomalyDINO based on the depicted reference sample.
 The second failure case occurs if the reference sample(s) does not resemble all concepts of normality, therefore $\mathcal{M}$ does not capture all variations of the nominal distribution $p_{\mathrm{norm}}$.
@@
 We like to remark, that this is only relevant to the few-shot setting, and with an increasing number of reference samples (and higher diversity of nominal patches in $\mathcal{M}$), the variance decreases notably.
-(a) Reference sample (left) and estimated anomaly map of ‘good’ test sample (right).
-(b) Reference sample (left) and estimated anomaly map of ‘good’ test sample (right).
-Figure 6: Example of an uninformative reference sample in category ‘Capsule’ (MVTec-AD).
-Some reference samples do not resemble the full concept of normality (here, the sample in Figure 6(b) does not show the text on the capsule such that any nominal sample with text will get a high anomaly score).
-Anomaly maps predicted by AnomalyDINO based on the depicted reference sample.
 (a) Test sample (left), the same test sample with ground-truth anomaly annotation (center), and a magnified view of a region with strong background artifacts (right).
-(b) Anomaly map predicted by AnomalyDINO (1-shot) without (left) and with masking (right).
-Figure 7: Visualization of the effect of masking in the presence of high background noise for the category ‘Capsules’ in VisA.
+(b) Anomaly map predicted by AnomalyDINO (1-shot) without masking (left) and with masking (right).
+Figure 8: Visualization of the effect of masking in the presence of high background noise for the category ‘Capsules’ in VisA (1-shot).
 As in Figure 5 we depict the anomaly map for ‘Capsules’ using a different colormap (red instead of yellow) for better visibility.
@@
 Discarding background patches helps to mitigate the problem of potential background noise, thereby reducing the number of false positives.
-We find that some samples of anomalous samples also show artifacts in the background.
-An example of the category ‘Capsules’ from VisA is depicted in Figure 7.
-Without masking, the method would correctly predict the depicted sample to show an anomaly, but for the wrong reason (high anomaly scores for the background noise).
-We conclude that masking is essential to faithfully detect anomalies whenever higher variations in the background are expected.
-A quantitative analysis is provided in the next paragraphs.
-In addition to masking (which decreases the size of $\mathcal{M}$), we consider augmenting the reference sample with rotations (which increases the size of $\mathcal{M}$).
+A representative example, here from the category ‘Capsules’ from the VisA dataset, is depicted in Figure 8.
+Without masking, the method would correctly predict the depicted sample to show an anomaly, but for the wrong reason (high anomaly scores caused by background noise/contamination).
+In contrast, the irrelevant background areas are discarded by our masking approach, such that the anomalous region is correctly identified.
+We conclude that masking is essential to faithfully detect anomalies whenever higher variations in the background are expected—and as the issue of low variations is particularly pressing in the few-shot regime, caused by the minimal amount of nominal reference samples, suitable masking can significantly boost the performance in these cases.
+This is showcased by the superior anomaly localization performance A quantitative analysis is provided in the next paragraphs.
+In addition to masking (which decreases the size of $\mathcal{M}$), we consider augmenting the reference sample with rotations (which increases the size and diversity of $\mathcal{M}$).
 Here we distinguish the ‘agnostic’ scenario, where we do not know a priori about potential rotations, and augment by default.
-Whenever we know about potential rotations of reference or test samples, we can deactivate the augmentation to reduce the test time.
-The preprocessing decision per category for MVTec-AD and VisA, inferred (solely) from the first nominal reference sample in the test set (to comply with the one-shot setting), are given in Table 7.
+Whenever we know about potential rotations of reference or test samples, we can deactivate the augmentation to reduce the size of the memory bank $\mathcal{M}$, the time to construct $\mathcal{M}$, as well as the test time itself.
 Table 8: Default preprocessing steps for MVTec-AD and VisA.
-We do not mask textures, indicated by (T).
-In addition, we do not apply masking when the masking test failed, indicated by (MT).
-See Section 3 and Figure 2 for further discussion and a visualization.
+We do not mask textures, as indicated by (T).
+In addition, we do not apply masking when the masking test on the first train sample failed, as indicated by (MT).
+See Section 3 and Figure 2 for further discussion and visualization.
 MVTec-AD Mask? Rotation?
@@
 Pipe fryum
+Figure 9: Effect of preprocessing on MVTec-AD.
+Anomaly detection of AnomalyDINO-S (448) in the 1-shot setting for different choices of the preprocessing pipeline (detection AUROC on image-level in %, mean and standard deviation over three independent runs).
+The preprocessing decision per category for MVTec-AD and VisA, inferred (solely) from the first nominal reference sample in $X_{\mathrm{ref}}$ (to comply with the one-shot setting), are given in Table 8.
+Figure 10: Effect of preprocessing on VisA.
+Anomaly detection with AnomalyDINO-S (448) in the 1-shot setting for different choices of the preprocessing pipeline (detection AUROC on image-level in %, mean and standard deviation over three independent runs).
+For VisA, the ‘informed’ scenario is equivalent to only applying masking (all categories), while ‘agnostic’ is equivalent to masking and augmentations (all categories), see Table 8.
 Effect of preprocessing on detection performance
-As discussed in Section 3 (and in the previous paragraph), suitable means to fill $\mathcal{M}$ influence the detection performance.
-The right preprocessing is in some cases more important than others.
-The full results are given in Figures 8 and 9.
-Figure 8: Effect of preprocessing on MVTec-AD.
-Anomaly detection of AnomalyDINO in the 1-shot setting for different choices of the preprocessing pipeline (detection AUROC on image-level in %, mean and standard deviation over three independent runs).
-Figure 9: Effect of preprocessing on VisA.
-Anomaly detection with AnomalyDINO in the 1-shot setting for different choices of the preprocessing pipeline (detection AUROC on image-level in %, mean and standard deviation over three independent runs).
+As discussed in Section 3 (and in the previous paragraph), suitable means to fill $\mathcal{M}$ and preprocess test samples, influence the detection performance.
+The results per object for MVTec-AD and VisA are given in Figures 9 and 10, respectively.
 Rotation
+Figure 11: Rotations as anomalies (‘misplaced’ transistor from MVTec-AD).
+The left anomaly map is estimated from a reference sample with rotations (‘agnostic’), and the anomaly is not detected.
+In contrast, the right anomaly map is based on a reference sample without rotations (‘informed’), and the anomaly is successfully detected.
+This example highlights the importance of a carefully designed preprocessing pipeline for each object.
 Increasing the diversity of nominal patch representations in $\mathcal{M}$ by rotating the reference sample can significantly improve the detection performance.
-Consider, e.g., the category ‘Screw’ in MVTec-AD, where the detection AUROC can be boosted from 65.6% to 90.1% in the 1-shot setting.
+Consider, e.g., the category ‘Screw’ in MVTec-AD, where the detection AUROC can be boosted from 65.6% to 89.2% in the 1-shot setting.
 This is intuitive, as the test samples of ‘Screw’ are taken from various angles.
 With sufficiently many reference samples, such data augmentation is not necessary anymore, but for the few-shot setting, we see major improvements.
-Figure 10: Rotations as anomalies (‘misplaced’ transistor from MVTec-AD).
-The left anomaly map is estimated from a reference sample with rotations (‘agnostic’), while the right anomaly map is based on a reference sample without rotations (‘informed’).
-The same holds—although to a lesser extent—for the categories ‘Hazelnut’, and ‘Metal nut’ (both MVTec-AD), as well as for ‘Macaroni2’, ‘Cashew’, ‘PCB2’ and ‘PCB4’ categories (all VisA).
-However, we also observe that rotations of the reference sample can actually decrease the detection performance in some categories, namely ‘PCB1’ and ‘PCB3’ (both VisA), and to a small extent also ‘Transistor’ (MVTec-AD).
+The same holds—although to a lesser extent—for the categories ‘Hazelnut’, ’Cable’ and ‘Wood’ (all MVTec-AD).
+However, we also observe that rotations of the reference sample can also decrease the detection performance in some categories, namely ‘PCB1/2/3’ and ‘Macaroni1’ (all VisA), and to a small extent also ‘Transistor’ (MVTec-AD).
 We attribute this to the fact the specific anomalies for these printed circuit boards contain rotated or bent connectors (see the PCB examples in Figure 5, right column).
@@
 The anomaly refers to a rotated transistor (anomaly type ‘misplaced’), and when (falsely) rotating the reference sample, such anomalies will not be detected—in contrast to the ‘informed’ preprocessing (right side of Figure 11).
+This highlights the importance of carefully designing the pre- and postprocessing pipeline for each object/product considered.
 Masking
-Across the board, we see that masking improves the detection performance.
-The only exception is ‘Candle’ from VisA with a slight decrease in performance.
-On further inspection, we find that the estimated foreground masks of DINOv2 occasionally fail to capture all relevant patches (the decision on whether to mask or not is inferred from the first nominal sample only, 1-shot setting).
-We leave the development of more robust and adaptive masking techniques for future work.
+Figure 12: Masking examples from MVTec-AD for all categories that passed the masking test (on a single train sample, see Table 8). Examples have not been cherry-picked.
+We utilize the zero-shot masking capabilities of DINOv2 to keep the overhead for this preprocessing step minimal.
+By applying a threshold to the first principal component [29], we can typically distinguish between the background and the foreground.
+We also employ a straightforward rule-based strategy to enhance the robustness and generalizability of the PAC-based masking in industrial settings: ensuring that the center crop of the image is predominantly occupied by the object of interest.
+This minor adjustment is necessary because industrial images often differ significantly from the data on which DINOv2 was originally trained.
+In addition, we use dilation and morphological closing to improve the quality of the mask.
+In our experiments, we applied masking only to the test samples as the size of $\mathcal{M}$ did not matter in the few-shot regime.
+Across the board, we see that the proposed masking technique improves the detection performance—in many categories even significantly, e.g., for ‘Capsules’, ‘Macaroni1/2’, or ‘PCB1/2/3/4’ (all VisA).
+We leave further improvements and the exploration of more advanced masking techniques for future work.
+Figure 13: Masking examples from VisA (not cherry-picked).
+Table 9: Runtime analysis on MVTec-AD (mean and std of inference time over all 1725 test samples, and mean and std of time to populate the memory bank for each object) for different shots (Tab.
+9(a)), preprocessing choices (Tab.
+9(b)), sample resolutions (Tab.
+9(c)) and model sizes (Tab.
+9(d)).
+All times are reported in seconds, measured on a single NVIDIA A40 with GPU warmup and CUDA kernel synchronization.
+The default setting is 1-shot, agnostic preprocessing, model size S, and resolution 448 (underlined for reference).
+Shots Inference Memory Bank
+1 0.060$\pm$0.012 00.52$\pm$0.040
+2 0.059$\pm$0.012 00.85$\pm$0.070
+4 0.059$\pm$0.012 01.58$\pm$0.080
+8 0.060$\pm$0.011 03.05$\pm$0.170
+16 0.063$\pm$0.012 06.02$\pm$0.390
+full (masking only) 0.067$\pm$0.010 16.62$\pm$4.430
+full (agnostic) 0.130$\pm$0.044 93.72$\pm$20.56
+(a) Runtime in dependence of shots.
+Mask? Rotate? Inference Memory Bank
+no no 0.055$\pm$0.011 0.17$\pm$0.03
+no yes 0.055$\pm$0.012 0.51$\pm$0.05
+yes no 0.068$\pm$0.012 0.18$\pm$0.03
+yes yes 0.067$\pm$0.012 0.51$\pm$0.04
+informed 0.059$\pm$0.016 0.22$\pm$0.11
+agnostic 0.060$\pm$0.012 0.52$\pm$0.04
+(b) Runtime in dependence of preprocessing choices.
+Resolution Inference Memory Bank
+224 0.043$\pm$0.010 0.31$\pm$0.04
+448 0.060$\pm$0.012 0.52$\pm$0.04
+672 0.086$\pm$0.014 0.75$\pm$0.06
+896 0.141$\pm$0.021 1.26$\pm$0.05
+(c) Runtime in dependence of image resolution.
+Model Size Inference Memory Bank
+S (21 M) 0.060$\pm$0.012 0.52$\pm$0.04
+B (86 M) 0.084$\pm$0.021 0.77$\pm$0.05
+L (300 M) 0.141$\pm$0.029 1.24$\pm$0.06
+G (1,100 M) 0.306$\pm$0.034 2.67$\pm$0.14
+(d) Runtime in dependence of model size.
 Runtime analysis
-The design choices in our pipeline influence the test time.
-To see the potential effects we measure the test time per sample on MVTec for the different combinations of masking (yes/no) and rotation (yes/no) for the full pipeline.
-This includes preparation of the memory bank $\mathcal{M}$ (i.e., feature extraction and the optional augmentation and masking operations), as well as the inference time for all test samples (i.e., feature extraction and optional masking).
-The results are given in Figure 11.
+The design choices in our pipeline influence the run time of the proposed method.
+To see the potential effects we measure the inference time per sample on MVTec-AD in various scenarios together with the time to set up the memory bank $\mathcal{M}$.
 The runtime was assessed utilizing a single NVIDIA A40 GPU, consistently employed throughout all experiments detailed in this paper (each experiment can be executed on a single NVIDIA A40 GPU).
-The resulting run time increases by a factor of 4, when comparing the baseline (no masking, no rotation) to the scenario where masking and rotations are applied to each product.
-Not rotating a reference sample in cases where it is unnecessary (the difference between ‘informed’ and ‘agnostic’) lowers the test time by approximately 30%.
-Figure 11: Effect on masking and augmentation on test time (per sample) in the 1-shot setting with AnomalyDINO.
-We measure the test time on a single NVIDIA A40 GPU and report mean and standard deviation over three independent runs on MVTec-AD.
+The runtime is measured with GPU warmup and CUDA kernel synchronization for a fair comparison.
+The results are given in Table 9.
+The average inference time per sample with AnomalyDINO-S (448) amounts to approximately 60ms in the few-shot regime ($\approx$ 16.7 samples/s), and only moderately increases with a larger memory bank (to 67ms ($+11\%$) for the full-shot scenario without augmentations).
+Compared to SOTA competitors in the one-shot regime, this is at least one order of magnitude faster (see Figure 3).
+Without any preprocessing steps, the 1-shot inference time amounts to approximately 55ms per sample ($\approx$ 18 samples/s).
+When applying both masking and rotations, the runtime increases from 55ms to 67ms, an moderate increase of roughly 23% (compared to the scenario without any preprocessing steps).
+The informed scenario can reduce the time to build the memory bank, but inference time is only affected for larger sample sizes.
+As expected, higher resolutions and larger architectures lead to increased runtimes.
 C.2 Scoring
+Table 10: Effect of aggregation statistics $q$ on the detection performance on MVTec-AD and VisA, evaluated for AnomalyDINO-S (448).
+$\mathcal{D}$ denotes the set of patch distances to the nominal memory bank, $H_{0.01}(\mathcal{D})$ the 1% highest entries thereof, and $\mathcal{A}$ denotes the anomaly map derived from $\mathcal{D}$ (anomaly scores for each pixel, after upsampling and smoothing, see Section 3).
+All results in %.
+Scoring $q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ $q=\mathrm{max}(\mathcal{\mathcal{D}})$ $q=\mathrm{max}(\mathcal{A})$
+Shots AUROC F1-max AP AUROC F1-max AP AUROC F1-max AP
+MVTec-AD 1 96.5$\pm$0.4 96.0$\pm$0.2 98.1$\pm$0.3 95.0$\pm$0.5 94.7$\pm$0.2 97.4$\pm$0.4 94.9$\pm$0.7 94.6$\pm$0.6 97.5$\pm$0.4
+2 96.7$\pm$0.8 96.5$\pm$0.4 98.1$\pm$0.7 95.5$\pm$1.0 95.3$\pm$0.5 97.4$\pm$0.6 95.0$\pm$1.3 94.8$\pm$0.8 97.4$\pm$0.9
+4 97.6$\pm$0.1 97.0$\pm$0.3 98.4$\pm$0.3 96.6$\pm$0.2 96.0$\pm$0.1 97.8$\pm$0.4 96.3$\pm$0.4 95.7$\pm$0.2 98.1$\pm$0.3
+8 98.0$\pm$0.1 97.4$\pm$0.1 99.0$\pm$0.2 97.2$\pm$0.1 96.5$\pm$0.4 98.6$\pm$0.1 97.0$\pm$0.0 96.3$\pm$0.1 98.6$\pm$0.1
+16 98.3$\pm$0.1 97.7$\pm$0.2 99.3$\pm$0.0 97.6$\pm$0.2 96.9$\pm$0.3 98.9$\pm$0.1 97.4$\pm$0.1 96.6$\pm$0.2 98.8$\pm$0.1
+VisA 1 85.6$\pm$1.5 83.1$\pm$1.1 86.6$\pm$1.3 82.4$\pm$1.9 81.4$\pm$1.0 84.0$\pm$1.7 80.5$\pm$1.3 80.6$\pm$0.8 82.1$\pm$0.9
+2 88.3$\pm$1.8 84.8$\pm$1.2 89.2$\pm$1.3 85.1$\pm$1.8 82.5$\pm$1.3 86.4$\pm$1.1 83.3$\pm$1.8 82.2$\pm$1.1 84.5$\pm$1.5
+4 91.3$\pm$0.8 87.5$\pm$1.0 91.8$\pm$0.7 88.4$\pm$0.3 84.9$\pm$0.6 89.0$\pm$0.4 86.8$\pm$1.4 84.3$\pm$0.9 87.6$\pm$1.3
+8 92.6$\pm$0.1 88.6$\pm$0.2 92.9$\pm$0.2 90.0$\pm$0.3 86.3$\pm$0.2 90.6$\pm$0.5 88.8$\pm$0.5 85.3$\pm$0.3 89.7$\pm$0.3
+16 93.8$\pm$0.1 89.9$\pm$0.3 94.2$\pm$0.3 91.6$\pm$0.5 87.2$\pm$0.7 92.2$\pm$0.4 90.5$\pm$0.3 86.8$\pm$0.5 91.5$\pm$0.4
 In Section 3 we investigate different ways of aggregating the anomaly scores $\mathcal{D}$ on patch-level (in our case, distances to the nominal memory bank $\mathcal{M}$) to an image-level anomaly score via a statistic $q$.
 Note that the segmentation results are therefore not affected by the choice of $q$.
-A standard choice is upsampling the patch distances of lower resolution to the full resolution of the test image using bilinear interpolation, then applying Gaussian smoothing operation (we follow Roth et al., 2022, and set $\sigma=4.0$), to obtain an anomaly map $\mathcal{A}$, and set $q=\max{(\mathcal{A})}$.
+A standard choice is upsampling the patch distances of lower resolution to the full resolution of the test image using bilinear interpolation, then applying Gaussian smoothing operation (we follow [33] and set $\sigma=4.0$), to obtain an anomaly map $\mathcal{A}$, and set $q=\max{(\mathcal{A})}$.
 We also evaluate two potential alternatives of $q$, our default choice $q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ (mean of the 1% highest entries in $\mathcal{D}$) and the maximum of $\mathcal{D}$.
-Table 8: Effect of aggregation statistics $q$ on the detection performance on MVTec-AD and VisA.
-$\mathcal{D}$ denotes the set of patch distances to the nominal memory bank, $H_{0.01}(\mathcal{D})$ the 1% highest entries thereof, and $\mathcal{A}$ denotes the anomaly map derived from $\mathcal{D}$ (anomaly scores for each pixel, after upsampling and smoothing, see Section 3).
-All results in %.
-Scoring $q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ $q=\mathrm{max}(\mathcal{\mathcal{A}})$ $q=\mathrm{max}(\mathcal{D})$
-Shots AUROC F1-max AP AUROC F1-max AP AUROC F1-max AP
-MVTec-AD 1 96.6$\pm$0.5 96.0$\pm$0.2 98.2$\pm$0.3 95.8$\pm$0.6 95.4$\pm$0.3 97.7$\pm$0.4 95.1$\pm$0.5 94.8$\pm$0.1 97.4$\pm$0.4
-2 96.9$\pm$0.8 96.6$\pm$0.4 98.2$\pm$0.7 96.2$\pm$1.1 96.0$\pm$0.4 97.7$\pm$0.7 95.6$\pm$1.0 95.4$\pm$0.5 97.4$\pm$0.6
-4 97.7$\pm$0.1 97.1$\pm$0.3 98.5$\pm$0.3 97.2$\pm$0.2 96.7$\pm$0.3 98.1$\pm$0.4 96.7$\pm$0.2 96.1$\pm$0.1 97.9$\pm$0.4
-8 98.2$\pm$0.1 97.5$\pm$0.1 99.1$\pm$0.2 97.7$\pm$0.1 97.1$\pm$0.3 98.8$\pm$0.1 97.3$\pm$0.2 96.5$\pm$0.4 98.6$\pm$0.1
-16 98.4$\pm$0.1 97.8$\pm$0.2 99.3$\pm$0.0 98.0$\pm$0.1 97.4$\pm$0.2 99.1$\pm$0.1 97.7$\pm$0.2 96.9$\pm$0.3 98.9$\pm$0.1
-VisA 1 86.2$\pm$1.4 83.5$\pm$1.0 87.7$\pm$1.4 85.1$\pm$1.7 82.7$\pm$1.0 87.1$\pm$1.3 83.8$\pm$1.8 82.0$\pm$1.0 85.8$\pm$1.6
-2 88.9$\pm$1.9 85.3$\pm$1.4 90.4$\pm$1.4 87.7$\pm$1.9 84.3$\pm$1.4 89.7$\pm$1.2 86.6$\pm$1.9 83.4$\pm$1.2 88.5$\pm$1.2
-4 91.5$\pm$0.7 87.6$\pm$0.9 92.6$\pm$0.6 90.7$\pm$0.3 86.4$\pm$0.6 92.1$\pm$0.3 89.8$\pm$0.3 85.7$\pm$0.5 91.2$\pm$0.5
-8 92.9$\pm$0.1 89.0$\pm$0.3 93.8$\pm$0.1 92.0$\pm$0.1 87.8$\pm$0.3 93.2$\pm$0.3 91.2$\pm$0.1 87.0$\pm$0.2 92.4$\pm$0.4
-16 94.0$\pm$0.0 90.0$\pm$0.1 94.8$\pm$0.1 93.3$\pm$0.2 88.9$\pm$0.4 94.4$\pm$0.2 92.6$\pm$0.3 88.2$\pm$0.4 93.7$\pm$0.3
-The results in Table 8 show that just the maximum patch distance leads to already good results, but upsampling and smoothing the patch distances improves upon that (maximum of $\mathcal{A}$).
+The results in Table 10 show that just the maximum patch distance leads to already good results while upsampling and smoothing the patch distances giving slightly weaker results (maximum of $\mathcal{A}$).
 Taking the mean of all patch distances above the 99% quantile ($q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$) improves above the standard choice.
@@
 The comparison including the best competing methods is depicted in Figure 14.
-We see that different architecture sizes have some effect on the image-level AUROC.
-AnomalyDINO-L outperforms smaller architectures, S (our default) and B, on VisA in the few-shot setting with $k\in\{1,2,4\}$.
-For $k\in{8,16}$ AnomalyDINO-L is on par with AnomalyDINO-S (and APRIL-GAN).
-For $k=16$ all architecture sizes outperform the best competitor (APRIL-GAN).
-On MVTec-AD, we see that AnomalyDINO-S performs slightly better, but the gap between AnomalyDINO-S and AnomalyDINO-B reduces with increasing shot size.
+We see that different architecture sizes have indeed an effect on the image-level AUROC.
+On MVTec-AD, we see that AnomalyDINO-S performs best, followed by the next larger model AnomalyDINO-B (which might contrast the common belief that larger models always perform better).
 In particular, all architecture sizes outperform the closest competitor (ADP).
-Refer to captionRefer to caption
-Figure 12: Effect of model size on detection AUROC.
-We report mean and standard deviation for the 1-shot setting on MVTec-AD (left) and VisA (right) over three seeds.
-Note, that the results for WinCLIP+ for 8 and 16 shots are those of the WinCLIP re-implementation by Kwak et al.
-(2024).
+Regarding VisA, larger architectures slightly outperform our default setting, AnomalyDINO-S.
+All architecture sizes are on par with APRIL-GAN at $k=4$, but outperform all competitors for $k>4$.
+Figure 14 also showcases that the performance of the proposed method scales preferably with the number of reference samples.
+(a) Detection ARUOC for MVTec-AD.
+(b) Detection ARUOC for VisA.
+Figure 14: Effect of model size on detection AUROC for MVTec-AD and VisA.
+We report mean and standard deviation for the 1-shot setting over three seeds.
+The image resolution of AnomalyDINO is set to 672.
+Note, that the results for WinCLIP+ for 8 and 16 shots are those of the WinCLIP re-implementation by [20].
 Appendix D Extending AnomalyDINO to the Batched Zero-Shot Setting
@@
 Recall, that in this setting all test samples $X_{\mathrm{test}}$ are provided (or at least a sufficiently large batch), but no (labeled) training or reference samples.
-The underlying (and necessary) assumption to meaningfully predict anomalies solely based on test samples, is that the majority of samples (or in our case, patches) at test time are from the nominal data distribution (see e.g., Li et al., 2024a).
+The underlying (and necessary) assumption to meaningfully predict anomalies solely based on test samples, is that the majority of samples (or in our case, patches) at test time are from the nominal data distribution (see e.g., [21]).
 We need to alter our method, outlined in Section 3, only slightly to adapt it to the batched setting.
-We score a test sample $\mathbf{x}^{(j)}\in X_{\mathrm{test}}$ in comparison to all remaining test sample $X_{\operatorname{test}}\setminus\{\mathbf{x}^{(j)}\}$, following the idea of mutual scoring (Li et al., 2024b).
+We score a test sample $\mathbf{x}^{(j)}\in X_{\mathrm{test}}$ in comparison to all remaining test sample $X_{\operatorname{test}}\setminus\{\mathbf{x}^{(j)}\}$, following the idea of mutual scoring [23].
 For each test sample $\mathbf{x}^{(j)}\in X_{\mathrm{test}}$ we therefore collect all patch representations not belonging to $\mathbf{x}^{(j)}$ in a memory bank, again utilizing DINOv2 as patch-level feature extractor $f$,
-$\mathcal{M}_{j}:=\bigcup_{\mathbf{x}^{(i)}\in X_{\operatorname{test}}\setminus\{\mathbf{x}^{(j)}\}}\bigl{\{}\mathbf{p}_{m}\;|\;f(\mathbf{x}^{(i)})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n}),\,m\in\{1,\dots,n\}\bigr{\}}\enspace.$ (5)
+$\mathcal{M}_{j}:=\bigcup_{\mathbf{x}^{(i)}\in X_{\operatorname{test}}\setminus\{\mathbf{x}^{(j)}\}}\bigl{\{}\mathbf{p}_{m}\mid f(\mathbf{x}^{(i)})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n}),\,m\in[n]\bigr{\}}\enspace.$ (5)
 We need to infer anomaly scores for each patch representation $\mathbf{p}_{\mathrm{test}}$ of $\mathbf{x}^{(j)}$.
@@
 We set $\alpha=0.1\%$ as anomalous patches are rare by assumption, and $\mathcal{D}(\mathbf{p}_{\mathrm{test}},\mathcal{M}_{j})$ large enough to accurately estimate the tail of the distribution.
-[^6] For MVTec-AD the total number of test patches extracted by DINOv2 (ViT-S) range between 43.008 and 171.008 per category, and for VisA between 163.200 and 334.464.
+[^6] For MVTec-AD the total number of test patches extracted by DINOv2 (ViT-S) at a resolution of 448 range between 43.008 and 171.008 per category, and for VisA between 163.200 and 334.464.
 The image-level score $s(\mathbf{x}_{\mathrm{test}})$ is again given by aggregating the patch-level anomaly score following Equation 4.
-Computation of the cosine distances and the proposed aggregation statistics can be effectively implemented as matrix operations on GPU such that the batched zero-shot inference time for AnomalyDINO amounts to roughly 0.1 seconds/sample for MVTec-AD (again measured on an NVIDIA A40; feature extraction takes approximately the same amount of time).
+Computation of the cosine distances and the proposed aggregation statistics can be effectively implemented as matrix operations on GPU such that the batched zero-shot inference time for AnomalyDINO amounts to roughly 60 ms/sample for MVTec-AD at a resolution of 448 (again measured on an NVIDIA A40 GPU).
 Some resulting anomaly maps in the batched zero-shot setting are depicted in Figure 15.
-Figure 13: Anomaly maps for the batched zero-shot setting on MVTec-AD. The left-most sample in each category is a ‘good’ test sample for reference, followed by four samples with anomalies.
+Figure 15: Anomaly maps for the batched zero-shot setting on MVTec-AD.
+The left-most sample in each category is a ‘good’ test sample for reference, followed by four randomly picked samples with anomalies.

### v3

@@
 We show that this approach does not only rival existing techniques but can even outmatch them in many settings.
-Our proposed vision-only approach, AnomalyDINO, is based on patch similarities and enables both image-level anomaly prediction and pixel-level anomaly segmentation.
+Our proposed vision-only approach, AnomalyDINO, follows the well-established patch-level deep nearest neighbor paradigm, and enables both image-level anomaly prediction and pixel-level anomaly segmentation.
 The approach is methodologically simple and training-free and, thus, does not require any additional data for fine-tuning or meta-learning.
@@
 1 Introduction
+footnotetext: This paper was accepted at the Winter Conference on Applications of Computer Vision (WACV), 2025. The final version is available here (Open Access) and on IEEE Xplore.
 Anomaly detection (AD) in machine learning attempts to identify instances that deviate substantially from the nominal data distribution $p_{\mathrm{norm}}(x)$.
@@
 The distances of those to the nominal representations in $\mathcal{M}$ give rise to an anomaly map and the corresponding anomaly score $s(\mathbf{x}_{\mathrm{test}})$ using the aggregation statistic $q$.
-For both, masking and feature extraction, we utilize DINOv2.
-Further examples for other categories are depicted on the right (and in Figures 4 and 5 in Appendix A).
+For both, masking and feature extraction, we utilize DINOv2 ($f$).
+Further examples for other categories are depicted on the right (and in Figures 4, 5, 6 and 7 in Appendix A).
 AD for industrial images has gained tremendous interest over the last couple of years.
@@
 This perspective is motivated by the observation that few-shot anomaly detection is feasible for human annotators based on visual features only, and does not require additional textual description of the given object or the expected types of anomalies (which are typically not known a priori).
-Our approach, termed AnomalyDINO, is memory bank-based and leverages DINOv2 [29] as a backbone (but the approach could be adapted to other architectures with strong patch-level feature extraction).
-We carefully design a suitable preprocessing pipeline for the one-shot scenario, using the segmentation abilities of DINOv2 (which alleviates the additional overhead of another model).
-At test time, anomalous samples are detected based on the high distances between their patch representations and the closest counterparts in the nominal memory bank.
-Due to its simplicity, AnomalyDINO can be deployed in industrial contexts very easily—in strong contrast to more complex approaches such as those proposed by [7] or [20].
+Our approach, termed AnomalyDINO, follows the well-established AD framework of patch-level deep nearest neighbor [46, 34], and leverages DINOv2 [30] as a backbone.
+We carefully design a suitable preprocessing pipeline for the few-shot scenario, using the zero-shot segmentation abilities of DINOv2 (which alleviates the additional overhead of another segmentation model).
+At test time, anomalous samples are detected based on the high distances between their patch representations and the closest counterparts in the nominal memory bank $\mathcal{M}$.
+Due to its simplicity, AnomalyDINO can be deployed in industrial contexts very easily—in strong contrast to more complex approaches such as [7] or [21].
 Yet, the proposed method achieves new state-of-the-art performance on anomaly detection in the few-shot regime on MVTec-AD [2] and outperforms all but one competing method on VisA [51].
@@
 DINO [5, 30] leverages a self-supervised student-teacher framework based on vision transformers [12].
-It employs a multi-view strategy to train models to predict softened teacher outputs, thereby learning robust and high-quality features for downstream tasks.
+It employs a multi-view strategy to train Vision Transformers (ViT) [11] to predict softened teacher outputs, thereby learning robust and high-quality features for downstream tasks.
 DINOv2 [30] combines ideas from DINO with patch-level reconstruction techniques [49] and scales to larger architectures and datasets.
@@
 Several works tackled this task by either training an anomaly classifier [38] or a generative model, which allows for reconstruction-based or likelihood-based AD [42, 26, 48, 9, 28].
-However, this work focuses on few-shot AD, which tackles anomaly detection under strong restrictions on the number of available training samples.
-Various recent methods recognized the effectiveness of pre-trained vision models for few-shot AD.
-The underlying idea in these approaches is to leverage the feature representations originating from pre-trained ResNets [17], Wide ResNets [44], and Vision Transformers (ViT) [11].
-These models can extract representations on a patch and pixel level, which can be compared against a memory bank, which aggregates patch and pixel features provided by the training set.
-The comparison with the memory bank is typically conducted using a nearest neighbor approach [1, 8, 10, 33, 41, 23].
-Loosely speaking, patches and pixels corresponding to anomalies are expected to have a high distance to their nearest neighbor in the memory bank.
+A well-established approach is that of deep nearest neighbor AD, where test instances are scored according to the distance to their nearest neighbors (in feature space) from the memory bank $\mathcal{M}$, containing features of nominal instances.
+This approach dates back to (at least) 2002 [13], but got later popularized first on image-level [1] and then by the seminal works on patch-level: Patch support vector data description (SVDD) [46] trains a patch-based feature extractor (with a mixture of DeepSVDD [37] and a self-supervised loss), and PatchCore [34] utilizes pre-trained classification models on Imagenet for patch-level feature extraction.
+Various further approaches build on this simple, but effective AD approach [8, 43, 24].
+Besides kNN matching, other ‘classical’ AD methods like the Mahalanobis distance are employed [10].
+Evidently, the performance of these approaches crucially depends on the feature extractor $f$.
+Classical choices are ResNets [18], Wide ResNets [47], and Vision Transformers [11], mostly pre-trained on a supervised task.
 Another line of work builds upon the success of pre-trained language-vision models in zero-shot classification.
@@
 This section introduces AnomalyDINO, which leverages DINOv2 to extract meaningful patch-level features.
-The strong feature extraction capabilities of the utilized backbone allow us to fall back on a simple patch-based matching strategy to perform few-shot and training-free AD for image data.
-We first gather relevant patch representations of nominal samples in a memory bank $\mathcal{M}$.
+We build upon the well-established deep nearest neighbor approaches [13, 46, 8, 34, 43, 1, 24], i.e., we first gather relevant patch representations of nominal patches in a memory bank $\mathcal{M}$.
 Then, for each test patch, we compute its distance to the nearest nominal patch in $\mathcal{M}$.
-A suitable aggregation of the patch-based distances gives the anomaly score on image-level.
-Inspired by PatchCore [33], our method simplifies the memory bank framework by selecting DINOv2 as the backbone, thereby reducing the complexity of the feature engineering stage—specifically in deciding which representations or layers to use and how to aggregate them.
-We also tailor the memory bank concept to the few-shot learning regime by introducing an improved aggregation statistic, and a new preprocessing pipeline that incorporates zero-shot masking and augmentations.
+A suitable aggregation of the patch-based distances gives anomaly scores on image-level.
+Our work differs from previous deep nearest neighbor approaches [34, 46] by tailoring the memory bank concept to the few-shot regime, specifically utilizing the strong features from DINOv2: We design a pipeline that incorporates zero-shot masking and augmentations, and we propose a more robust aggregation statistic.
+Importantly, we identify DINOv2 as the ideal backbone for our scenario due to its strong patch-level features and masking ability.
+In addition, this simplifies the deep nearest neighbor framework by reducing the complexity of the feature engineering stage (e.g., by making it unnecessary to think of which representations/layers to use and how to aggregate them), and increasing the flexibility w.r.t.
+input resolution (that can be chosen to be any multiple of $14$).
 The proposed method is described in detail in the following subsections.
 3.1 Anomaly Detection via Patch (Dis-) Similarities
-Let us briefly review the idea of the memory bank approach.
+Let us briefly review the idea of patch-level deep nearest neighbor AD.
 We assume to have access to a suitable feature extractor $f:\mathcal{X}\rightarrow\mathcal{F}^{n}$ that maps each image $\mathbf{x}\in\mathcal{X}$ to a tuple of patch-features $f(\mathbf{x})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n})$, where $\mathcal{X}$ denotes some space of images and $\mathcal{F}$ the feature space.
-Note that $n$ depends on the image resolution and the patch size.
-Given $k\geq 1$ nominal reference sample(s) $X_{\text{ref}}:=\bigl{\{}\mathbf{x}^{(i)}\mid\,i\in[k]\bigr{\}}$ (with shorthand $[k]:=\{1,\dots,k\}$), we collect the nominal patch features and store them in a memory bank
+Note that $n$ depends on the image resolution and the patch size (in the case of DINOv2, a patch is $14\times 14$ pixels).
+Given $k\geq 1$ nominal reference sample(s) $X_{\text{ref}}:=\{\mathbf{x}^{(i)}\mid\,i\in[k]\}$ (with shorthand $[k]:=\{1,\dots,k\}$), we collect the nominal patch features and store them in a memory bank
 $\begin{split}\mathcal{M}:=\bigcup_{\mathbf{x}^{(i)}\in X_{\operatorname{ref}}}\bigl{\{}\mathbf{p}_{j}^{(i)}\mid f(\mathbf{x}^{(i)})=(\mathbf{p}_{1}^{(i)},\dots,\mathbf{p}_{n}^{(i)}),\,j\in[n]\bigr{\}}\enspace.\end{split}$ (1)
 To score a test sample $\mathbf{x}_{\mathrm{test}}$, we collect the extracted patch representations $f(\mathbf{x}_{\mathrm{test}})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n})$ and then check how well they comply with $\mathcal{M}$.
-To do so, we leverage a nearest-neighbor approach to find the distance to the closest reference patch for a given test patch $\mathbf{p}\in\mathcal{F}$
+To do so, we leverage a nearest neighbor approach to find the distance to the closest reference patch for a given test patch $\mathbf{p}\in\mathcal{F}$
 $d_{\operatorname{NN}}(\mathbf{p};\mathcal{M}):=\min_{\mathbf{p}_{\operatorname{ref}}\in\mathcal{M}}d(\mathbf{p},\,\mathbf{p}_{\operatorname{ref}})$ (2)
@@
 $s(\mathbf{x}_{\mathrm{test}}):=q\big{(}\{d_{\operatorname{NN}}(\mathbf{p}_{1};\mathcal{M}),\dots,d_{\operatorname{NN}}(\mathbf{p}_{n};\mathcal{M})\}\big{)}\enspace.$ (4)
-Throughout this paper, we define $q$ as the average distance of the 1% most anomalous patches, i.e, $q(\mathcal{D}):=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ with $H_{0.01}(\mathcal{D})$ containing the 1% highest values in $\mathcal{D}$.
-The statistic $q$ can be understood as an empirical estimate of the tail value at risk for the 99% quantile [26] and turns out suitable for a wide range of settings because it balances two desirable properties: First, we want $s(\mathbf{x}_{\mathrm{test}})$ to depend on the highest patch distances, as they may provide the strongest anomaly signal.
+Throughout this paper, we define $q$ as the average distance of the 1% most anomalous patches, i.e, $q(\mathcal{D}):=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ with $H_{0.01}(\mathcal{D})$ containing the 1% highest values in the set $\mathcal{D}$.
+The statistic $q$ can be understood as an empirical estimate of the tail value at risk for the 99% quantile [27] and turns out to be suitable for a wide range of settings because it balances two desirable properties: First, we want $s(\mathbf{x}_{\mathrm{test}})$ to depend on the highest patch distances, as they may provide the strongest anomaly signal.
 Similarly, we want a certain degree of robustness against a singular high patch distance (in particular in few-shot scenarios where $\mathcal{M}$ is sparsely populated).
@@
 Following [34], we utilize bilinear upsampling and Gaussian smoothing ($\sigma=4.0$) to turn the patch distances into pixel-level anomaly scores for localization of potential defects.
-Examples of the resulting anomaly maps are visualized in Figure 1 and Appendix A (Figures 4 and 5).
+Examples of the resulting anomaly maps are visualized in Figure 1 and Appendix A (Figures 4, 5, 6 and 7).
 We extend AnomalyDINO also to the batched zero-shot scenario, see Appendix D (Figure 18).
@@
 This is essential because the variability at test time is likely to be significantly greater than in the reference data, $X_{\mathrm{ref}}$.
+In contrast, full-shot methods aim to reduce the size of $\mathcal{M}$ to reduce inference time, e.g., [34].
 To avoid irrelevant areas of the test image from leading to falsely high anomaly scores, we propose masking the object of interest.
 This approach mitigates the risk of false positives, particularly in the few-shot regime where the limited reference samples may not adequately capture the natural variations in the background, as exemplified in Figure 10, Appendix C.
-It is important to note that the appropriate preprocessing technique—whether to apply masking and/or augmentations—should depend on the specific characteristics of the object of interest.
+It is important to note that the appropriate preprocessing technique—whether to apply masking and/or augmentations—should depend on the specific characteristics of the object(s) of interest.
 For a more detailed discussion on the challenges and considerations in designing an effective preprocessing pipeline, see Section C.1.
 Masking
-Masking, i.e., delineating the primary object(s) in an image from its background, can help to reduce false-positive predictions.
+Masking, i.e., delineating the primary object(s) in an image from its background, can help to reduce false-positive predictions, thereby improving the robustness in the low-data regime.
 To minimize the overhead of the proposed pipeline, we utilize DINOv2 also for masking.
@@
 In addition, we utilize dilation and morphological closing to eliminate small holes and gaps within the predicted masks.
-See Figures 12 and 13 for examples of this masking procedure, Table 8 for the outcomes of the masking test per object, and Figure 8 for a visualization of the benefits of masking in the presence of background noise (all in Appendix C).
+See Figures 14 and 15 for examples of this masking procedure, Table 8 for the outcomes of the masking test per object, and Figure 10 for a visualization of the benefits of masking in the presence of background noise (all in Appendix C).
 In general, we do not mask textures (e.g., ‘Wood’ or ‘Tile’ in MVTec-AD).
@@
 Moreover, APRIL-GAN and AnomalyCLIP, require training on a related dataset, which is in contrast to our proposed training-free method.
-In Tables 2 and 3, results reported by [18] are indicated by , the result of the WinCLIP re-implementation by [20] (where J.
+In Tables 2 and 3, results reported from [19] are indicated by , the result of the WinCLIP re-implementation [21] (where J.
 Jeong is also coauthor) by .
@@
 We summarize the results for few-shot anomaly detection on MVTec-AD and VisA in Table 2 and Table 3, respectively.
+Regarding MVTec-AD, our method achieves state-of-the-art $k$-shot detection performance across all $k\in\{1,2,4,8,16\}$ for every reported metric, outperforming approaches that require additional training data sets (such as ADP and APRIL-GAN).
+The method also demonstrates superior anomaly localization, with results showing that while detection performance remains comparable across different resolutions (448 vs.
+672), a higher resolution enhances localization performance.
+Furthermore, we observe clear improvements as the number of samples increases.
 […]
@@
 1-shot
-SPADE 82.9$\pm$2.6 91.1$\pm$1.0 91.7$\pm$1.2 92.0$\pm$0.3 44.5$\pm$1.0 85.7$\pm$0.7
-PatchCore 86.3$\pm$3.3 92.0$\pm$1.5 93.8$\pm$1.7 93.3$\pm$0.6 53.0$\pm$1.7 82.3$\pm$1.3
+SPADE 81.0$\pm$2.0 90.3$\pm$0.8 90.6$\pm$0.8 91.2$\pm$0.4 42.4$\pm$1.0 83.9$\pm$0.7
+PatchCore 83.4$\pm$3.0 90.5$\pm$1.5 92.2$\pm$1.5 92.0$\pm$1.0 50.4$\pm$2.1 79.7$\pm$2.0
 […]
@@
 2-shot
-SPADE 81.0$\pm$2.0 90.3$\pm$0.8 90.6$\pm$0.8 91.2$\pm$0.4 42.4$\pm$1.0 83.9$\pm$0.7
-PatchCore 83.4$\pm$3.0 90.5$\pm$1.5 92.2$\pm$1.5 92.0$\pm$1.0 50.4$\pm$2.1 79.7$\pm$2.0
+SPADE 82.9$\pm$2.6 91.1$\pm$1.0 91.7$\pm$1.2 92.0$\pm$0.3 44.5$\pm$1.0 85.7$\pm$0.7
+PatchCore 86.3$\pm$3.3 92.0$\pm$1.5 93.8$\pm$1.7 93.3$\pm$0.6 53.0$\pm$1.7 82.3$\pm$1.3
 […]
@@
 AnomalyDINO-S (672) 98.4$\pm$0.1 97.6$\pm$0.1 99.3$\pm$0.0 97.5$\pm$0.0 62.7$\pm$0.1 94.0$\pm$0.1
-Regarding MVTec-AD, our method achieves state-of-the-art $k$-shot detection performance across all $k\in{1,2,4,8,16}$ for every reported metric, outperforming approaches that require additional training data sets (such as ADP and APRIL-GAN).
-The method also demonstrates superior anomaly localization, with results showing that while detection performance remains comparable across different resolutions (448 vs.
-672), a higher resolution enhances localization performance.
-Furthermore, we observe clear improvements as the number of samples increases.
 […]
@@
 The effect of proper preprocessing seems to increase with higher resolutions.
-Additionally, we compare the two different preprocessing settings, ‘agnostic’ (our default) and ‘informed’, and find that the agnostic preprocessing slightly outperforms the informed counterpart, see Figures 9 and 10.
+Additionally, we compare the two different preprocessing settings, ‘agnostic’ (our default) and ‘informed’, and find that the agnostic preprocessing slightly outperforms the informed counterpart, see Figures 11 and 12.
 Depending on the product category, suitable preprocessing steps lead to substantial improvements.
@@
 For the full discussion, see Section C.1.
-Aggregation choice
+Aggregation statistic
 We compare the default scoring method, the empirical tail value at risk, to other suitable aggregations in Section C.2.
@@
 Architecture size/choice
-We also evaluated the proposed framework using Vision Transformers pre-trained on ImageNet (see Figure 3).
-While these backbones slightly outperform PatchCore, they give weaker features compared to DINOv2 and are incompatible with the proposed masking procedure (zero-shot masking does not function effectively).
+We also evaluated the proposed framework to Vision Transformers pre-trained on ImageNet, see Figure 3 and Sec.
+C.4.
+While these backbones slightly outperform PatchCore, they give weaker features compared to DINOv2 and are incompatible with the proposed masking procedure.
 This underlines that DINOv2 is indeed excellently suited for visual AD.
@@
 In this context, we find that operating at a resolution of 448 gives the best trade-off between performance and inference time, but even higher detection AUROCs are possible at higher resolutions.
-In addition, we evaluated our pipeline with DINOv2 at different scales (ViT-S, ViT-B, ViT-L).
+We also evaluated our pipeline with DINOv2 at different distillation sizes (ViT-S, ViT-B, ViT-L).
 The full results are given in Section C.3, Figure 16.
@@
 Our method is based on similarities between patch representations extracted by DINOv2 [30].
-We carefully design means to populate the nominal memory bank with diverse and relevant features while minimizing false positives utilizing zero-shot segmentation and simple data augmentation.
+We carefully design means to populate the nominal memory bank with diverse and relevant features while minimizing false positives by utilizing zero-shot segmentation and simple data augmentation.
 Industrial settings require fast deployment, easy debugging and error correction, and rapid adaptation for covariance shifts in the normal data distribution.
@@
 However, the individual parts of our pipeline can easily be exchanged for more sophisticated alternatives.
-For example, our simple masking approach could be replaced by more specialized and adaptive masking techniques (which may also be relevant to other methods building on features extracted by DINOv2), and the simple upsampling and smoothing approach could be substituted by more sophisticated methods like [14] to obtain the anomaly map from patch features.
+For example, our simple masking approach could be replaced by more specialized and adaptive masking techniques (which may also be relevant to other methods building on DINOv2), and the simple upsampling and smoothing approach could be substituted by more sophisticated methods like [15].
 It would be interesting to see if that leads to further improvements in anomaly detection and localization, thereby further reducing the gap between few- and full-shot anomaly detectors.
 We also plan to improve the batched zero-shot performance of AnomalyDINO.
-Broader Impacts
-Advancing few-shot visual anomaly detection methods can offer various benefits by enhancing manufacturing quality control through the rapid identification of defects with minimal nominal examples, which improves efficiency, reduces waste, and improves overall safety in the product lifecycle.
-Similar positive benefits can be expected outside the industrial domain, e.g., for healthcare diagnostics or environmental monitoring.
-It is, however, essential to recognize the shortcomings of automated anomaly detection systems.
-We believe that simpler methods can be adapted more quickly, monitored more effectively, and are therefore more reliable.
-In this context, awareness of the risks of overreliance must be heightened (see Appendix B for identified failure cases of the proposed method).
-In addition, strong visual anomaly detectors could also lead to potentially malicious or unintended uses.
-To address these concerns, including potential privacy infringements and possible socioeconomic impacts of automation, strategies such as establishing robust data governance, and implementing strict privacy protections are essential.
-Additionally, investing in workforce development can help manage the socioeconomic effects of automation and leverage the full potential of strong visual anomaly detection.
+Acknowledgement
+The authors acknowledge funding from TRR 391 Spatio-temporal Statistics for the Transition of Energy and Transport by the German Research Foundation (DFG).
+Supplementary Material
 Appendix A Detailed experimental results
-In this section, we present additional visualization and results for the proposed method.
-query GT 1-shot 8-shot query GT 1-shot 8-shot
-Figure 4: Examples – MVTec-AD.
+Query GT 1-shot 8-shot
+Figure 4: Examples – MVTec-AD (1/2).
 Depicted are, from left to right, a test sample per category (Query), the ground truth anomaly annotation (GT), and the predicted anomaly map from AnomalyDINO-S (448) in the 1- and 8-shot settings.
@@
 score over ‘good’ test samples.
-Further anomaly maps, predicted by AnomalyDINO, are presented in Figures 4 and 5. The full results per category for MVTec-AD and VisA are given in Tables 5 and 6, respectively.
+Further anomaly maps, predicted by AnomalyDINO, are presented in Figures 4, 5, 6 and 7. The full results per category for MVTec-AD and VisA are given in Tables 6 and 7, respectively.
+The results presented in Tables 6 and 7 show that a) anomalies in some categories are more difficult to than others, b), that the performance of AnomalyDINO increases across the board with more available reference samples.
+In particular, more complex objects like ‘PCB3’ and ‘PCB4’ in VisA or ‘Transistor’ in MVTec-AD seem to benefit the most from more available reference samples.
+In addition, we see that also the variance in the reported metrics decreases with the number of nominal samples.
+In this context, we observe a peculiarity of the few-shot regime (which does not occur for sufficiently populated $\mathcal{M}$).
+The results crucially depend on the chosen reference image(s).
+[^6] Note that this holds for all one- and few-shot methods, not only for AnomalyDINO.
+The high variances, predominantly in the 1- and 2-shot setting, for category ‘Capsule’ for MVTec-AD, or ‘Cashew’ or ‘PCB4’ for VisA demonstrate this.
+We discuss the potential failure cases of choosing a sub-optimal reference sample in the following section (Section B.2).
+Finally, we also report the full-shot performance, see Table 5.
+As the diversity within $\mathcal{M}$ is sufficiently high given the large number of reference samples, we only apply masing here (no augmentations).
+The results demonstrate that the performance further improves for all considered metrics.
+Notably, AnomalyDINO-S (672) achieves new state-of-the-art segmentation performance measured in (AU)PRO in the full-shot setting (see here, accessed 11/26/2024).
+Table 5: Full-shot results on MVTec-AD and VisA with AnomalyDINO-S in the default setting (no std reported as results are deterministic when all samples are considered).
+Dataset Resolution Detection Segmentation
+AUROC F1-max AP AUROC F1-max PRO
+MVTec-AD 448 99.3 98.8 99.7 97.9 61.8 93.9
+672 99.5 99.0 99.8 98.2 64.3 95.0
+VisA 448 97.2 93.7 97.6 98.7 50.5 95.0
+672 97.6 94.5 98.0 98.8 53.8 96.1
+Query GT 1-shot 8-shot
+Figure 5: Examples – MVTec-AD (2/2). See Fig. 4 for a description.
 Table 6: Detailed results on MVTec-AD.
@@
 Mean 87.4$\pm$1.2 92.5$\pm$0.5 89.7$\pm$1.3 93.4$\pm$0.6 92.6$\pm$0.9 94.1$\pm$0.1 93.8$\pm$0.3 94.8$\pm$0.2 94.8$\pm$0.2 95.3$\pm$0.2
-query GT 1-shot 8-shot query GT 1-shot 8-shot
+Query GT 1-shot 8-shot
 Figure 6: Examples – VisA (1/2).
@@
 Best viewed at a higher zoom level as some anomalies are quite small.
-The results presented in Tables 5 and 6 show that a) anomalies in some categories are more difficult to than others, b) detection and segmentation performance correlate strongly (measured in image-level AUROC and PRO), and c), that the performance of AnomalyDINO increases across the board with more available reference samples.
-In particular, more complex objects like ‘PCB3’ and ‘PCB4’ in VisA or ‘Transistor’ in MVTec-AD seem to benefit the most from more available reference samples.
-In addition, we see that also the variance in the reported metrics decreases with the number of nominal samples.
-In this context, we observe a peculiarity of the few-shot regime (which does not occur for sufficiently populated $\mathcal{M}$). The results crucially depend on the chosen reference image(s).
-Table 7: Full-shot results on MVTec-AD and VisA with AnomalyDINO-S in the default setting (no std reported as results are deterministic when all samples are considered).
-Dataset Resolution Detection Segmentation
-AUROC F1-max AP AUROC F1-max PRO
-MVTec-AD 448 99.3 98.8 99.7 97.9 61.8 93.9
-672 99.5 99.0 99.8 98.2 64.3 95.0
-VisA 448 97.2 93.7 97.6 98.7 50.5 95.0
-672 97.6 94.5 98.0 98.8 53.8 96.1
-The high variances, predominantly in the 1- and 2-shot setting, for category ‘Capsule’ for MVTec-AD, or ‘Cashew’ or ‘PCB4’ for VisA demonstrate this.
-We discuss the potential failure cases of choosing a sub-optimal reference sample in the following section (Section B.2).
-Finally, we also report the full-shot performance, see Table 7.
-As the diversity within $\mathcal{M}$ is sufficiently high given the large number of reference samples, we only apply masing here (no augmentations).
-The results demonstrate that the performance nicely improves for all considered metrics.
+Query GT 1-shot 8-shot
+Figure 7: Examples – VisA (2/2). See Fig. 6 for the description.
 Appendix B Limitations and failure cases
@@
 Consider, for instance, the anomalies shown in Figure 8.
-(a) Reference sample
+(a) Reference sample (No anomalies)
 (b) Anomaly type ‘Bent Wire’
@@
 Figure 8: Example of a semantic anomaly in category ‘Cable’ (MVTec-AD).
-Depicted are a nominal reference sample, an example of a sensory anomaly (Figure 6(b)) and a semantic anomaly (Figure 6(c)) with anomaly maps predicted by AnomalyDINO (1-shot).
+Depicted are a nominal reference sample, a sensory anomaly (Figure 8(b)), and a semantic anomaly (Figure 8(c)) with anomaly maps predicted by AnomalyDINO (1-shot).
 While the cable with anomaly type ‘Bent Wire’ (Figure 8(b)) contains patches that cannot be well matched to patches of the reference image, this does not apply to the anomaly ‘Cable Swap’ (Figure 8(c)).
@@
 Therefore, parts of the text of a nominal test sample may be falsely recognized as anomalies.
-Due to the strong dependency on a suitable reference sample, we observe higher AD variances for some products in the one-shot setting, as shown in Tables 5 and 6.
+Due to the strong dependency on a suitable reference sample, we observe higher AD variances for some products in the one-shot setting, as shown in Tables 6 and 7.
 We like to remark, that this is only relevant to the few-shot setting, and with an increasing number of reference samples (and higher diversity of nominal patches in $\mathcal{M}$), the variance decreases notably.
@@
 Whenever we know about potential rotations of reference or test samples, we can deactivate the augmentation to reduce the size of the memory bank $\mathcal{M}$, the time to construct $\mathcal{M}$, as well as the test time itself.
+The preprocessing decision per category for MVTec-AD and VisA, inferred (solely) from the first nominal reference sample in $X_{\mathrm{ref}}$ (to comply with the one-shot setting), are given in Table 8.
 Table 8: Default preprocessing steps for MVTec-AD and VisA.
@@
 Anomaly detection of AnomalyDINO-S (448) in the 1-shot setting for different choices of the preprocessing pipeline (detection AUROC on image-level in %, mean and standard deviation over three independent runs).
-The preprocessing decision per category for MVTec-AD and VisA, inferred (solely) from the first nominal reference sample in $X_{\mathrm{ref}}$ (to comply with the one-shot setting), are given in Table 8.
 Figure 12: Effect of preprocessing on VisA.
@@
 As discussed in Section 3 (and in the previous paragraph), suitable means to fill $\mathcal{M}$ and preprocess test samples, influence the detection performance.
-The results per object for MVTec-AD and VisA are given in Figures 9 and 10, respectively.
+The results per object for MVTec-AD and VisA are given in Figures 11 and 12, respectively.
 Rotation
+Increasing the diversity of nominal patch representations in $\mathcal{M}$ by rotating the reference sample can significantly improve the detection performance.
+Consider, e.g., the category ‘Screw’ in MVTec-AD, where the detection AUROC can be boosted from 65.6% to 89.2% in the 1-shot setting.
+This is intuitive, as the test samples of ‘Screw’ are taken from various angles.
+With sufficiently many reference samples, such data augmentation is not necessary anymore, but for the few-shot setting, we see major improvements.
 Figure 13: Rotations as anomalies (‘misplaced’ transistor from MVTec-AD).
@@
 This example highlights the importance of a carefully designed preprocessing pipeline for each object.
-Increasing the diversity of nominal patch representations in $\mathcal{M}$ by rotating the reference sample can significantly improve the detection performance.
-Consider, e.g., the category ‘Screw’ in MVTec-AD, where the detection AUROC can be boosted from 65.6% to 89.2% in the 1-shot setting.
-This is intuitive, as the test samples of ‘Screw’ are taken from various angles.
-With sufficiently many reference samples, such data augmentation is not necessary anymore, but for the few-shot setting, we see major improvements.
-The same holds—although to a lesser extent—for the categories ‘Hazelnut’, ’Cable’ and ‘Wood’ (all MVTec-AD).
+The same holds—although to a lesser extent—for the categories ‘Hazelnut’, ’Cable’, and ‘Wood’ (all MVTec-AD).
 However, we also observe that rotations of the reference sample can also decrease the detection performance in some categories, namely ‘PCB1/2/3’ and ‘Macaroni1’ (all VisA), and to a small extent also ‘Transistor’ (MVTec-AD).
+Figure 14: Masking examples from MVTec-AD for all categories that passed the masking test (on a single train sample, see Table 8).
 We attribute this to the fact the specific anomalies for these printed circuit boards contain rotated or bent connectors (see the PCB examples in Figure 6, right column).
@@
 This highlights the importance of carefully designing the pre- and postprocessing pipeline for each object/product considered.
+Figure 15: Masking examples from VisA.
 Masking
-Figure 12: Masking examples from MVTec-AD for all categories that passed the masking test (on a single train sample, see Table 8). Examples have not been cherry-picked.
 We utilize the zero-shot masking capabilities of DINOv2 to keep the overhead for this preprocessing step minimal.
@@
 We leave further improvements and the exploration of more advanced masking techniques for future work.
-Figure 13: Masking examples from VisA (not cherry-picked).
 Table 9: Runtime analysis on MVTec-AD (mean and std of inference time over all 1725 test samples, and mean and std of time to populate the memory bank for each object) for different shots (Tab.
@@
 Without any preprocessing steps, the 1-shot inference time amounts to approximately 55ms per sample ($\approx$ 18 samples/s).
-When applying both masking and rotations, the runtime increases from 55ms to 67ms, an moderate increase of roughly 23% (compared to the scenario without any preprocessing steps).
+When applying both masking and rotations, the runtime increases from 55ms to 67ms, a moderate increase of roughly 23% (compared to the scenario without any preprocessing steps).
 The informed scenario can reduce the time to build the memory bank, but inference time is only affected for larger sample sizes.
@@
 A standard choice is upsampling the patch distances of lower resolution to the full resolution of the test image using bilinear interpolation, then applying Gaussian smoothing operation (we follow [34] and set $\sigma=4.0$), to obtain an anomaly map $\mathcal{A}$, and set $q=\max{(\mathcal{A})}$.
-We also evaluate two potential alternatives of $q$, our default choice $q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ (mean of the 1% highest entries in $\mathcal{D}$) and the maximum of $\mathcal{D}$.
+We also evaluate two potential alternatives of $q$, our default choice $q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ (mean of the 1% highest entries in $\mathcal{D}$) and $q=\max(\mathcal{D})$.
 The results in Table 10 show that just the maximum patch distance leads to already good results while upsampling and smoothing the patch distances giving slightly weaker results (maximum of $\mathcal{A}$).
@@
 C.3 Architecture Size
+(a) Detection AUROC for MVTec-AD.
+(b) Detection AUROC for VisA.
+Figure 16: Effect of model size on detection AUROC for MVTec-AD and VisA (mean and std over three seeds).
+The image resolution of AnomalyDINO is set to 672.
+Note, that the results for WinCLIP+ for 8 and 16 shots are those of the WinCLIP re-implementation [21].
 DINOv2 is available in different distillation sizes (S, B, L, and G).
@@
 Figure 16 also showcases that the performance of the proposed method scales preferably with the number of reference samples.
-(a) Detection ARUOC for MVTec-AD.
-(b) Detection ARUOC for VisA.
-Figure 14: Effect of model size on detection AUROC for MVTec-AD and VisA.
-We report mean and standard deviation for the 1-shot setting over three seeds.
-The image resolution of AnomalyDINO is set to 672.
-Note, that the results for WinCLIP+ for 8 and 16 shots are those of the WinCLIP re-implementation by [20].
+(a) Detection AUROC for MVTec-AD.
+(b) Detection AUROC for VisA.
+Figure 17: Effect of backbone choice DINOv2 (trained in a self-supervised fashion) compared to ViT (supervised training on ImageNet) on detection AUROC for MVTec-AD and VisA (mean and std over three seeds).
+For better comparability, image resolution of AnomalyDINO is 224 for all models.
+C.4 Backbone Choice
+In our experiments, we find that DINOv2 provides excellently suited features for few-shot AD.
+This is already evident in Fig.
+3, here we investigate the effect on the detection performance on MVTec-AD and VisA in more detail.
+We compare the performance in few-shot AD of ViT-B and ViT-L pre-trained on ImageNet to that of AnomalyDINO-B and AnomalyDINO-L (utilizing DINOv2 in distillation sizes B and L).
+To ensure comparability, we deactivate any pre- and postprocessing for AnomalyDINO (no augmentations, no zero-shot masking) and set the image resolution to 224 for all models.
+The results are depicted in Figure 17.
+We observe that DINOv2 significantly improves the detection performance compared to those of the ViT pre-trained on the image classification tasks, giving a performance gain of at least $+4\%$ AUROC on MVTec-AD and VisA.
+This demonstrates that the features extracted by DINOv2 are better suited compared to those of ImageNet-pretrained ViT-{B/L}.
+In Sections 3 and 4 we demonstrate that further (substantial) improvements are possible with suitable pre- and postprocessing like augmentations and zero-shot masking (which is not possible based on the features from supervised features) and that the image resolution (or the effective patch size) can also greatly boost performance.
+See also Fig.
+16 for the performance in dependence of the model size.
 Appendix D Extending AnomalyDINO to the Batched Zero-Shot Setting
@@
 Note, however, that $\mathcal{M}_{j}$ may now contain nominal and anomalous patches.
-Thus, the nearest-neighbor approach is not suitable anymore: the nearest neighbor in $\mathcal{M}_{j}$ for (the representation of) an anomalous patch might also be abnormal, and the resulting distance therefore not informative.
+Thus, the nearest neighbor approach is not suitable anymore: the nearest neighbor in $\mathcal{M}_{j}$ for (the representation of) an anomalous patch might also be abnormal, and the resulting distance therefore not informative.
 A simple solution, based on the assumption that the majority of patches are nominal, is to replace the nearest neighbor with a suitable aggregation statistic over the distribution of patches distances in $\mathcal{M}_{j}$
@@
 The left-most sample in each category is a ‘good’ test sample for reference, followed by four randomly picked samples with anomalies.
+Appendix E Broader Impacts
+Advancing few-shot visual anomaly detection methods can offer various benefits by enhancing manufacturing quality control through the rapid identification of defects with minimal nominal examples, which improves efficiency, reduces waste, and improves overall safety in the product lifecycle.
+Similar positive benefits can be expected outside the industrial domain, e.g., for healthcare diagnostics or environmental monitoring.
+It is, however, essential to recognize the shortcomings of automated anomaly detection systems.
+We believe that simpler methods can be adapted more quickly, monitored more effectively, and are therefore more reliable.
+In this context, awareness of the risks of overreliance must be heightened (see Appendix B for identified failure cases of the proposed method).
+In addition, strong visual anomaly detectors could also lead to potentially malicious or unintended uses.
+To address these concerns, including potential privacy infringements and possible socioeconomic impacts of automation, strategies such as establishing robust data governance, and implementing strict privacy protections are essential.
+Additionally, investing in workforce development can help manage the socioeconomic effects of automation and leverage the full potential of strong visual anomaly detection.